A Complete Bibliography of Publications in Communications in Statistics: Simulation

and Computation: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

15 June 2022 Version 1.05

Title word cross-reference

(k, l) [RSG21]. 2 [NN20, OP20]. 3 [OP20]. ^{1} [GZZ22]. _{2.5} [GZ21b]. _{RS} [CKCG21]. A [MPD20]. X̄ [MMMG21, CKT^{+}20, HCTZ21, NKSY22]. C_{pk} [PDO^{+}20]. D [LFMMRH22, GBHA20, GBHA21]. F [SX22]. G [AAHR20]. I [CLJK21]. I^{2} [HBM^{+}20]. K [LW21, FY21, GPP21, MAMA20]. L [Hut21]. L_{2} [Zhe21]. L_{p} [SD22]. M [Pol20, RYA20, SCK21, GBHA20]. p [CL22b, DL20]. ρ [Bar20a]. S_{B} [MM20]. sin^{k}(x) [MS22]. t [BA22, NIN20, Pan20c, RBN^{+}22]. T^{2} [AMK^{+}21a, AF22b, MLR^{+}22, SSBA22]. U [GK20]. φ [LY20]. W [NCC^{+}22, SMM22]. X [PM20].

-class [FY21]. -coefficient [LW22b]. -control [PM20]. -copulas [GBHA21]. -estimation [DA21]. -estimator [RYA20, SCK21]. -estimators [Hut21]. -Inflated [RSG21]. -Loss [Zhe21]. -means [LW21]. -mixing [LY20]. -modulated [DL20]. -of- [SMM22]. -optimal [LFMMRH22, MPD20].

-regression [Pol20]. -sample [GBHA20]. -spline [CLJK21]. -stage [NN20]. -statistics [GK20]. -transformed [Pan20c]. -value [CL22b].

1 [BC21].

AB/BA [ZL20]. abrupt [MMMG21]. absence [EH22]. absolutely [RK20]. accelerated [ABPK21, Bar20a, CSL20, GAS20, ZLSC20]. Acceptance [Ger20, DR22]. Accounting [ZL22]. accuracy [AT21, CL22b, PY20]. across [De 20]. acyclic [HPZ^{+}21]. adaLASSO [EGVdR21]. adapted [AF22b, dSdSTC21]. Adaptive [ESES20, FY21, AH20, Ali22, ASM22, GI20, KB21, Lev21, LYY21, NN20, STdL21, Yad22, YDHW21]. additive [AS21b, LMP20, PHN22, PBLM22, SKV20a, ZG22a, ZFG22, ZGL22]. Addressing [MAAA21, Yad22]. Adjusted [TNN20, TAA22]. adjustment [CSE^{+}21, Hep21b]. Advances [FP20]. against [AG21, OP22]. age [NG22]. agent [YH21a]. aggregating [MV22]. aggregation [TK22b]. agreement [TDD20]. AIDS [TWTT21]. air [KKO21]. aircraft [LRF20]. algorithm [ASM21, CCH20, CL22c, DKKA22, GW20, HYSC21, KTZ20, LW21, MS22, MW20, NN22, RM20, SMO21, Tel22, TWTT20, VPT21]. algorithmic [MPD20]. Algorithms [JPH21, BV22, ES22, NIN20]. allocation [AM21, CMMV21]. almost [VW21]. alternating [GW20]. alternative [AF22b, Bap22, SKV20a]. alternatives [AG21]. American [MFS21]. Analysis [RSG21, ALB22, ANP20, Avc21, BS21, CSL20, CKK21, CDB^{+}21, CSG22, CKCG21, DA22, DF22, FN20, GGMM22, GW20, HAT^{+}22, Har21, KOT20, MFPP21, OM21, Pan20a, PRH21, Pol22, RM20, RY21, SMB22, SBZG22, WP22, WMS22, Yas22, YWT22, ZA21, ZL22, ZNKN22, ZL20, dOMD^{+}21]. analysis-based [FN20]. Analytic [LPP21]. analytics [ML22]. analyzing [LYZ^{+}21]. animal [PHN22]. Anticipated [LX22]. Application [ASM21, ULN21, YH21a, Abd18, AW22, AM22, CL22a, CSX20, CS22, CDB^{+}21, DA22, GZ21b, HAT^{+}22, HORA22, HYSC21, IMR21, LLZ22, PASB21, RBRAa21, SPG^{+}22, SS21a, fSLC21, SKJ20, TWTT21, XZJ^{+}21]. Applications [Alt20b, AAHR20, BAEk21, BSNT21, BE20, GNE21, JKH20, KS20c, KR22, KE22, LX22, LRF20, Yas22]. Applied [MSL21, AA20d, BFAR20, NJ22, RCSS22, Ver21]. Applying [GAS20]. approach [AES21, Alg20, Alt20a, Ami21, Avc21, AA20d, BDD^{+}22, CYS21, CLL20, De 20, GI20, Gai21, GAS20, HI22, Hut21, MPD20, Nad21, PHS21, PASB21, RM20, SFG21, ULN21, YC22]. Approaches [Wes20, PV20b]. Approximate [HK20, YW20]. approximation [AY21b, CL22b, Lav21, Lee20]. Approximations [NZ21, GBHA21, Hom20]. arbitrarily [Hor22]. ARCH} [OB21, BGM20]. area [IC21, MW20, MSS20]. Areal [BJ20]. arising [BAEk21]. array [Ger20]. Artificial [KE22]. asian} [CHL22]. ask [LWC22]. aspects [FP20]. Assessing [Lee20, CP22]. assessment [ANP20]. asset [AM21, MFS21]. asset-allocation [AM21]. associated [DYC22, GABG22, GBHA20, HZAF22]. assumption [Par20]. assumptions [DYC22]. assurance [BJU20]. asymmetric [Hep21a, Zhe21]. asymmetry [GZ22b, ZG22a]. Asymptotic [BPN22, Esk21, Mil20, BS21, MVLMA22, TKF^{+}20, WLC^{+}22]. asymptotically [DRB20, DYC22]. asymptotics [GW22, dXLY20a, dXGLY20, dXLY20b]. atmospheric [KKO21]. attribute [ABPK21, WZ21a]. attribute-variable [WZ21a]. attributed [MASA22]. attributes [MS20]. augmented [AGA20]. autocorrelated [SMM22]. autocorrelation [RAG21]. automated [EGVdR21]. Automatic [AT21, VPT21, Tel22]. Autometrics [EGVdR21]. Autoregressive [GKM20, AE21, GNE21, LLZ22, LW22b, NIN20, WWYX21, WHW21]. auxiliary [AS21a, BPP20, HB22, JuAH20, NKSY22, uARS22, SVM20, SVK22]. availability [SS21a]. average [AK20, CLS22, LLZ22, uARS22, OP20]. averaging [HD22, HCZ21, dSdSTC21]. axially [VWZ21].

B [YZ21]. B-spline [YZ21]. BA [ZL20]. background [CS22]. backward [LX22]. balanced [AATD21, MPD20]. balancing [HAT^{+}22]. band [DLRZ21]. banded [JM20]. bandwidth [ESES20, Sol21, ZZ22]. Banerjee [PM20]. barrier [LG21]. Bartlett [Pic21]. base [YWT22]. based [ASZN21, ABHF22, AMK^{+}21a, AY21a, AS22, AJ22, AA22b, ABPK21, ABJ21b, AA20d, ASP22, BC21, BM21, CKK21, CLS22, CS22, CMMV21, CCM20, CNL21, DKKA22, De 20, DA21, Ere20, FN20, GK22a, GK22b, GWZC22, HI22, HR22, Hep21b, HCTZ21, HE21, IM20, IS20, KR21, KS22, LES21, LLL22, MG21, MNL21, MSL21, MQ21, MLR^{+}22, Mil20, MB21, MBBS22, MW20, MK22, NdSA20, NN21, NKSY22, PHY20, PD21, PHS21, RYA20, RAJ21, RCSS22, SI20, STdL21, SBTP20, Sha22, SL21, SMWJ20, SS21b, SB21b, SM20, SZJG21, SKS21, SCK22, SCK20, SCK21, SLE20, TWTT20, VDN22, WS20, WWYX21, WJZ^{+}22, WMJ21, Wu21, XXH22, XWL22, XH21, YÖT20, YZ21, ZZFG20, ZGC21, ZG22a, ZNKN22, ZB20, ZLSC20, ZBZS20, vO21]. baseline [DSL20, PBLM22]. basketball [SS21b, SZS20]. Basu [dOMA21]. Basu-Dhar [dOMA21]. batch [PK22]. Bayes [AMK21b, Esk21, YW20, ZGC21]. Bayesian [ASZN21, AA20b, AA20c, Ami21, AJ22, ALVQ22, Avc21, AA20d, BJ20, BC21, CSL20, CKK21, CLY21, DA22, FLPB20, GGB20, GJK21, Han20, Han21, HvZ21, dAIIS20, IT21, KK22, KP21, Lev21, MNN21, MdN22, MZZF21, OLOK22, PCWL22, RdAIS^{+}20, SLE20, VDN22, WP22, Yad22, YBC22, YWT22, ZB20, ZLSC20]. behavior [WJB21]. behavioral [ML22]. behaviors [EFD21]. Ben [PM20]. Berk [HM21]. Berkson [WCF21]. Bernstein [DY22]. Berry [DYC22]. best [Ami21, CZ22, Ver21]. beta [CNL21, KP21]. Better [IR21]. between [AK20, AE21, CST21, Lui22, MNN21, MPS21, OP20, PK22, PAT22, UYMK22, YC22, ZZ21]. between-batch [PK22]. Bi [SS21a]. Bi-dimensional [SS21a]. Bias [KAVR20, BZ22a, BR21, GCF^{+}20, HZAF22, SCHS22, ZPW21]. biased [CZS21]. biasing [ÖT20, SCK20]. bid [LWC22]. bidimensional [GTW22]. big [XH21]. BiMM [SWC^{+}20]. bimodal [ROL^{+}20]. binary [DKKA22, FS22, LFMMRH22, MASA22, OBN22, SI20, SWC^{+}20]. BINMA [SKJ20]. binomial [AF22a, GPP21, KS20c, SK22a, UYMK22, WWYX21, ZWY20]. biochemical [HAT^{+}22]. biosimilar [PK22]. Birnbaum [BA22]. birth [RdAIS^{+}20]. birthdates [PRH21]. Bivariate [Ere20, MPS21, ASP22, Bap20, Bap22, GAS22, MLR^{+}22, NAKK20, NM21, RK20, SSBA22, SZ21, WLT22, dXLY20a, dXGLY20, dOMA21]. blank [SKV20b]. blinded [TF22]. block [Bey21, BJVV21, MPD20, MBDK21, SD22]. blocked [ZK20b]. boosting [LYP22]. Bootstrap [BCP21, FH22, ZZ22, Bey21, BRA20, Hut21, MBDK21, MSS20, PDO^{+}20, PV20b, Sol21]. Bootstrapping [CS20, XW20]. bottleneck [ML22]. bound [DYC22]. boundary [NZ21]. bounded [dALFB21]. branching [RY21]. breaks [De 20]. Burr [AS22]. Bursa [NN22]. business [ML22].

calculations [CFM21]. calibrated [Mag20]. Calibration [SSS20]. cancer [HAT^{+}22, RBRAa21, ZA21]. canonical [GW20, Pol22]. capability [ABJ21b, PG22, PDO^{+}20, RAJ21]. capture [Kal21]. Capturing [BCR^{+}21]. cards [SKV20b]. care [SCE^{+}20]. Carlo [CL22b, FPAA21, KKZ21, KOT20, uIKKH22, Kur21c, LPNP20, Uya22, WMS22, ZL20]. case [CSE^{+}21, GKZ^{+}22, KOT20, MNL21, MAMA20, NN22, ONE22, Pan20a, SB20, SPG^{+}22, SCE^{+}20, SMB22, SRO20, dOMA21]. case- [MAMA20]. case-cohort [Pan20a]. case-mix [CSE^{+}21, SCE^{+}20]. CAT [AW22]. categorical [AMK^{+}21a]. category [Avc21]. Cauchy [KAVR20]. causes [CSL20]. CDF [Ras20]. censored [ASZN21, ARAB21, AJ22, AMK21b, AHG20, ASP22, CLY21, Lee20, MG21, MK21, MAMA20, MB21, MBBS22, PHY20, PCWL22, PV20a, RBRAa21, SBTP20, Sha22, Wu21, ZA21]. censoring [CSL20, CKK21, GAS20, KB21, KP21, LTR21, MJ20, Mis20, PMK20, WLT22, WL22, YBC22]. censorship [FZZ20]. CEV [ARAB21]. CEV-Hybrid [ARAB21]. chain [HE21, SS21b, SLE20]. change [BSNT21, MMMG21, QYC22, ZT20]. change-point [BSNT21]. change-points [ZT20]. changes [BVJV22]. chaos [BCDA21]. character [SKV20b]. characteristics [CSE^{+}21, MK21, YBC22]. characterization [Mil20]. chart [AMK^{+}21a, AK20, AK22, ASM22, ARJ20, ABPK21, CLS22, CKT^{+}20, GK22a, HU20, HORA22, HCTZ21, JuAH20, KL21, LK21a, dALFB21, MMMG21, MLR^{+}22, uARS22, SAF21, SAA21, STdL21, SB21b, TAA22, WZ21a]. charts [AH20, ARAB21, EAJ20, FN20, GK22b, HB22, KBR21, KS20a, NKSY22, PM20, SSBA22, SN22]. checkerboard [GBHA21]. checking [ZG22a]. children [IM20]. choice [ALVQ22, LW22a]. Cholesky [vO21]. Cholesky-based [vO21]. Choosing [Ver21]. CIR [MFS21, TMZL21]. circular [AATD21, BZ22a]. claim [TK22b]. class [AF22a, ATYK20, BZ22b, CCM20, FY21, Ruz22, SK22b]. classes [NG22]. classification [DKKA22, IAG21, MS20, YÖT20, ZC22]. classifications [dOVQP21]. classifiers [DKKA22]. classifying [PK21]. Closed [Lev21, HK20, LWC22, PHN22]. Closed-form [Lev21, LWC22]. closeness [WWTM22]. cluster [BM21, CDB^{+}21, KS20c, dOMD^{+}21]. cluster-based [BM21]. clustered [SWC^{+}20]. clustering [AY21a, CGS21, CCH20, HYSC21, LH20, Tel22, VPT21]. clusters [BDD^{+}22, FBNRG21, Tel22, ZNKN22]. codifference [GTW22]. coefficient [AS21a, CKL22, ESS22, GWZC22, GZZ22, GP20, HCZ21, IR21, KJ20, LZ22, LW22b, Mos22, SGG22, SVM20, SBZG22, TNN20, ZZFG20, ZG22a, Zha22, ZZWH20, ZZ22, ZXF22]. coefficient-based [ZG22a]. cognitive [ML22]. cohort [Pan20a]. Coincidence [Pic21]. cointegrating [AÖ20b]. cointegration [Hep21a, uIKKH22]. COM [AK22, SKJ20]. COM-Poisson [AK22, SKJ20]. combination [REN22, SU21]. combinations [De 20, YH21a, Yad22]. combine [CL21]. Combined [Hua20, WZ21a]. Combining [CGS21, ÖA22, IC21]. Comments [CL22a]. Common [Kal21, KJ20, TNN20]. commonly [Lui22, Uya22]. commonly-used [Lui22]. communication [SMWJ20]. compact [WJB21]. company [BS21]. comparative [AE21, BA22, Bay21, KGR21, MFPP21]. compare [CYS21]. Comparing [AG21, CSG22, dAIIS20, SS20b, Dog22, FO22, SG21, SS20a]. Comparison [CLL20, CCBN22, ONE22, PBLM22, UR22, AKS^{+}21, CMMV21, GLT20, IM20, IRA21, JM20, uIKKH22, Lui22, MSL21, MPS21, SCHS22, SS20a, Uya22, WL20]. Comparisons [AGA20, HM21, Wu21]. competing [AHG20, CSL20, WLT22, WL22]. complete [ASZN21, ASP22]. complex [BCR^{+}21, PK21, PHS21]. component [AE21, FN20, Kur21b, ZNKN22]. components [LMP20, VLvdW22, YW20]. Composite [MMR22, DA22, HHY21, LMP20, MW20, Pan20b, SK22b, TWTT21]. compositional [GGMM22]. compound [ZK20b]. Computation [GLT20, Bar20a, SMO21]. Computational [BFAR20, FP20, YÖT20]. computationally [LW22a, SY22]. Computing [Nag21a, SVK22]. concepts [ZNKN22]. conceptual [KÖ21a]. concomitant [AA22b]. concomitant-based [AA22b]. condition [FPAA21]. conditional [BCP21, DKL22, ESES20, HCTZ21, KBR21, dXLY20a, XWL22]. conditionally [Mil21, YH21b]. conditions [LW22b]. conduct [dOVQP21]. Confidence [CST21, DP20, UYMK22, BCP21, DLRZ21, IRS20, IR21, LLL22, MNN21, PDO^{+}20, TNN20, WWTM22]. confounders [DNWG22]. conic [YÖT20]. considering [MASA22]. consistent [MNL21]. constant [LG21]. constrained [Pan20a, ZBZS20]. constraints [DY22, MQ21, RVV22]. constructed [JHR21, SSBA22]. Construction [Bey21, MPD20]. contagion [CDB^{+}21]. containing [GAS22]. contaminated [Dog22, SFG21]. contamination [AM22]. context [DKL22]. contingency [KJK20]. continual [UKK20, ZWY21]. continuous [AMK^{+}21a, GGB20, RSG21, RK20, ZG22a]. Control [EAJ20, KBR21, AMK^{+}21a, AK22, ASM22, ARJ20, ABPK21, Ava21, DC20, FN20, GK22a, GK22b, HORA22, HE21, JuAH20, KL21, dALFB21, MMMG21, MLR^{+}22, uARS22, PM20, RYA20, SAA21, STdL21, SSBA22, SMO21, SS20b, SN22, TAA22, WZ21a, Wu21, Yas22]. Controlling [LAZ21]. convergence [GBHA20]. converting [Mag20]. coordinates [GGMM22, KS20b]. Copula [ADK21, QL21, GAS20, GBHA20, HD20, HE21, KBR21, LES21, SLE20, dXLY20a, dXGLY20, YC22]. copula-based [HE21, LES21, SLE20]. copulas [GBHA21, SSBA22, JHR21]. correct [CNL21]. Corrected [GCF^{+}20, Özk21]. Corrected-Hill [GCF^{+}20]. Correcting [BR21]. correction [BZ22a, HZAF22, Hep21a, ZPW21]. correlated [AA22a, AS21b, LES21, Yad22]. Correlation [CZP21, ZG22a, Bap20, Bap22, GWZC22, GZZ22, GW20, MS22, Pol22, UR22, Wes20, vO21]. cost [MYVD21, PM20, SA22]. costs [ZC22]. count [HD20, Hom20, MV22, OLOK22, TEO20]. counting [GBHA20, LZ22, PAT22]. counts [MPS21, TK22b, dSdSTC21]. course [ML22]. covariance [ALB22, AM21, BZ22b, GZ21a, JM20, RAG21, ZK20b]. covariate [HAT^{+}22, MAMA20]. covariates [DSL20, MZZF21, NCC^{+}22, SBZG22, WHW21, WCF21]. Coverage [IRS20]. Cox [AJ22, CMA21, Pan20a, SI20, ULN21, ZGC21]. Cramér [ZZFG20]. credit [CHL22, GHW20]. cricket [BT20]. crisis [NN22]. Criteria [Hom20, CCM20, Kim21, Wes20]. criterion [BSNT21, DRB20, HK22, HCZ21, KÖ21a, Mis20]. Cross [GTW22, Alg20, Pol22]. Cross-codifference [GTW22]. cross-language [Pol22]. cross-validation [Alg20]. crossing [NZ21]. crossover [KS22, ZL20]. cum [KGF20]. cumulative [CHC20]. cure [RdAIS^{+}20, SL21]. current [SPS22]. curtailed [CLY21]. curve [ANP20, LK21b, PY20]. curves [CHC20]. customer [YWT22]. customer-base [YWT22]. customized [KS20a]. CUSUM [AH20, HB22, KL21, RYA20].

D [Ava21, OP20]. D-Trace [Ava21]. daily [SMB22]. Danish [MNL21]. Data [VWZ21, Abd18, AMK^{+}21a, AY21a, AA20b, ARAB21, AM21, AJ22, ALVQ22, AMK21b, AÖ20b, AM22, AHG20, ASP22, BJ20, BDD^{+}22, CL22a, CLL20, CCH20, CZP21, CZS21, DP20, DA22, FZM21, FZZ20, GI20, GAS22, GGB20, GGMM22, GP20, GZ21b, HD20, HCZ21, HYSC21, IAG21, IC21, KS22, LH20, LYY21, dALFB21, LES21, LYZ^{+}21, LRF20, MT21, MAMA20, MB21, MBBS22, MV22, Mos22, Nad21, Nak21, NAG21b, OLOK22, PHY20, PCWL22, Pap20, PK21, PHS21, PV20b, RBRAa21, SAA21, SBTP20, SCHS22, SMB22, SFG21, Sha22, SKK21, TY21, TK22b, TWTT20, TEO20, UM22, UR22, WHW21, WL22, WWTM22, WJB21, XH21, ZA21, ZGL22, ZB20, ZZWH20, ZL20]. database [Ver21]. datasets [CL22c]. Daya [PM20]. deal [SKK21, SK22b]. debit [GHW20]. deciles [SB21b]. decision [MM22, NG22, PMK20, SWC^{+}20]. Decrease [EL21]. decreasing [PM20]. deeper [ML22]. default [CHL22, fSLC21]. degradation [CSDT20, SBZG22]. degree [IMR21]. degrees [TEO20]. delay [SMWJ20, SMO21]. delay-based [SMWJ20]. delayed [YH21a]. delta [MNN21]. delta-lognormal [MNN21]. dengue [HYSC21]. denoising [Lev21, TP22]. densities [QL21]. density [Bar20a, BCP21, BZ22a, CZS21, HK22, dAIIS20, LJC^{+}20, QL21, Sha20, SD22, WP22, WN21b]. dependence [ADK21, CDB^{+}21]. dependency [YC22]. dependent [AÖ20b, ARJ20, ABPK21, ABJ21a, ABJ21b, BCR^{+}21, CSL20, JHR21, LZ22, MAMA20, Med22, NG22, PV20b, RAJ21, Sha20, SM20, TKF^{+}20, dOVQP21, WLT22, ZD22]. derivatives [YDHW21]. derived [UYMK22]. Design [MNL21, PHS21, SPG^{+}22, ABJ21a, CP22, CKT^{+}20, KGF20, KL21, KK22, PM20, RGA21, REN22, SY22, Yad22, ZWY21, ZD22]. Design-based [PHS21]. Design-consistent [MNL21]. designed [WZ21a]. Designing [ABPK21, BJU20, MS20, HORA22, SAF21]. designs [AATD21, AGA20, AJ22, AAA20, BJVV21, BVJV22, FTC21, GJK21, KS22, LFMMRH22, Lui22, MPD20, NN20, NKSY22]. desirability [SCHS22]. destructive [RdAIS^{+}20]. detect [GK22a, GK22b]. Detecting [AE21, OP20, WZ21a, FBNRG21]. Detection [GZ22b, GZ22a, WS21a, ZGL21, AMK^{+}21a, BDD^{+}22, SCK22, TK22a, UEA20]. detections [QYC22]. determinant [BZ22b]. Determination [ABJ21a, SGG22]. Determining [PJL21, IM20]. deterministic [JGGDMRLD21]. detrending [GG22]. Developing [NN21]. development [FS22, PK22]. Deviance [Kim21]. device [SKV20b, SSK22]. Dewma [ARAB21]. Dhar [dOMA21]. diagnosed [IM20]. diagnostic [ANP20, HBM^{+}20, HE21, KL21]. diagnostics [AÖ20a, KUA21]. dichotomous [ZL20]. difference [CST21, MNN21, UYMK22]. differences [OP20]. different [AM21, CYS21, DSL20, GLT20, Han20, Han21, HHY21, Pol20, ROL^{+}20, TK22b, UKK20]. differential [LX22, MZZF21]. diffusion [GHW20, LLLL22, MFS21, SBZG22, WMJ21]. Diggle [CLL20]. dimension [CZP21]. dimensional [CKCG21, DNWG22, FY21, IAG21, JM20, Kur21a, Kur21b, LYY21, MT21, Nak21, SS21a, VLvdW22, WX21, WCF21]. dimensions [HHY21]. directed [HPZ^{+}21]. direction [GW20]. directional [ADK21, BJVV21, NAG21b]. Dirichlet [LH20, MVLMA22, WJB21]. discontinuous [AA20d]. discovery [BR21]. Discrete [HZAF22, ADK21, SS21b, VPT21]. discrete-time [SS21b]. discriminant [KS20b, WP22]. Discriminating [PAT22]. Discrimination [UM22, TP22]. disorder [IM20]. dispersion [AS21a, ESS22, SVM20, dOMD^{+}21]. Distance [NdSA20, AY21a]. Distance-based [NdSA20]. Distinguishing [GKZ^{+}22]. distortion [dXLY20b, ZG21, ZG22a, ZFG22, ZGL22]. distributed [DDK21, HORA22, IRS20, MAAA21, ÖT20]. Distribution [Alt20b, TF22, Abd18, ASZN21, ABHF22, AK22, ARAB21, Ami21, AMK21b, AA22b, ABPK21, ASP22, BA22, Bar20a, BAEk21, CL22a, CSDT20, CCBN22, Dog22, GPP21, GABG22, GW22, HR22, IMR21, JAP21, JKH20, KAVR20, KR21, Kal21, KLK21, KTZ20, KBR21, KB21, KS20c, KR22, KP21, LH20, LPP21, LTR21, LRF20, LY20, MK21, MSL21, MM20, MBBS22, NAKK20, NdSA20, NIN20, NM21, ON21, OP22, PHN22, PBLM22, PASB21, PMK20, RL21, RBRAa21, Ras20, Ric21, RK20, Sha22, UKK20, VG22, WLT22, YBC22, YWLD20, ZNLW20, ZBZS20, ZL20, dOMA21, Biç20].

Distributions [GBHA20, AS22, AAHR20, ARJ21, BJU20, BPN22, CKK21, CCBJ21, CLY21, FN20, Hom20, Kim21, KGR21, MJ20, MNN21, MB21, MFPP21, Pan20c, PAT22, RYA20, SKG22, TJ20, UEA20, WZY^{+}21]. divergence [LK21b, MK22, PHY20]. Diverse [DKKA22]. dividend [LG21]. dose [NN20]. dose-response [NN20]. Double [LK21a, AK20, dALFB21, MMMG21]. doubly [PIA22, Wu21]. dropout [CLL20]. drug [Yad22]. DSLRIG [SAD^{+}21]. Dual [fSLC21, HB22, LFZ21]. Dual-frailty [fSLC21]. Dynamic [dSdSTC21, DNWG22, MPS21, RRA20, SMWJ20, TEM20, TAA22]. dynamics [DF22].

E-Bayesian [Han20, Han21]. E-MSE [Han21]. E-posterior [Han20]. each [Ver21]. early [PM20, TK22a]. EBLUP [MSS20]. Economic [NKSY22, ABJ21a, PM20, RGA21, REN22]. economic-statistical [NKSY22, REN22]. economical [DR22]. economically [WZ21a]. Education [SN22]. educational [UM22]. effect [BJVV21, GZ21a, Har21, IM20, ML22, MV22, NSN21, SVK22]. Effects [PK22, TEM20, AW22, BCR^{+}21, EH22, GLT20, Hua20, Lar20, MASA22, PHN22, Pol20, SLW20, TWTT20, TWTT21, YW20, ZK20a, ZL20]. Efficiency [ÖT20, CKL22, Mil20, SSBA22, XWL22]. Efficient [BM21, MYVD21, NIN20, SB20, SU21, Esk21, MS22, SVK22, SY22]. eigenvalue [Ava21, WS21b]. elastic [LYY21, Pol22]. elastic-net [LYY21]. elements [VPT21]. elicitation [SK22a]. elongated [Tel22]. emerging [TK22a]. emphasis [HBM^{+}20]. Empirical [Che21, LK21b, LLL22, LY20, CZ22, Esk21, Gho20, MQ21, ÖA22, SKS21, TY21, WN21b, YWT22, ZZWH20]. energy [YC22]. enhanced [HU20]. Enhancing [Ver21]. Ensemble [IAG21, LYP22, DKKA22]. entrance [XXH22]. entropies [NYM20]. entropy [CZP21, Mis20]. envelope [YV20]. environmental [dALFB21]. epistemic [BCDA21]. equality [AG21, MJ20, PFM21]. equation [Dog22, MZZF21]. equations [LX22, LYZ^{+}21, REMO22, Wes20]. equilibrium [TEM20]. equivalence [DC20]. ergodicity [TMZL21]. error [CR21, CYS21, DC20, EH22, EL21, GZ22c, Gha22, Han21, Hep21a, Hor22, IRS20, LAZ21, MSS20, uARS22, PK22, SCE^{+}20, Sha20, WX21, WN21a, WCF21, YA21, YRAB22, ZNLW20, ZGC21]. error-in-variable [WCF21]. error-in-variables [WX21]. errors [GZ22b, GZ22a, GWZC22, GZZ22, KZG21, MAAA21, MS20, PZ20, RGA21, Sha20, SVK22, SBB22, WS20, WLY22, YDLX20, ZG21, ZFG22, ZGL22]. errors-in-variables [WS20]. Esseen [DYC22]. estimate [Gho20, SBB22]. estimated [CKT^{+}20, LK21a, SAF21, YHGA21]. estimates [Gho20, KAVR20]. Estimating [GBHA21, LTR21, TK22b, YDHW21, FTC21, KLM^{+}22, LYZ^{+}21, Ras20, REMO22, Wes20]. Estimation [ALB22, HPZ^{+}21, JHR21, KSAR21, KB21, LES21, MK21, MES21, MBBS22, NS20, PKA22, PZ20, SA22, Sha20, SKV20b, SKJ20, VLvdW22, WWYX21, ZNLW20, ZGC21, ZXF22, AG20, ASZN21, AS21a, Ava21, AM22, BA22, BPP20, BRA20, BZ22b, CCBJ21, CSX20, CS22, CLJK21, CZS21, DRB20, De 20, DDK21, DNWG22, ESES20, Esk21, FY21, FLPB20, GNE21, GCF^{+}20, HD20, HHY21, dAIIS20, IC21, KZG21, KN22, KP21, LPNP20, LJC^{+}20, LK21b, LW21, LYY21, LMP20, LRF20, MG21, MZ22, MZZF21, Mil21, Mos22, MFPP21, MYVD21, NAKK20, NIN20, Nis21, ÖT20, ÖA22, Pan20a, QL21, RBN^{+}22, RVV22, SB20, SZ20, SCE^{+}20, SD22, SKV20a, SK22b, SPS22, SSK22, Sol21, SCK20, TKF^{+}20, TF22, Tel22, TÇA22, WY21, WZ21b, XWL22, YBC22, YRAB22, YWLD20, ZG22b, DA21]. estimations [Han20, Han21, NYM20, fSLC21]. estimator [AÖ20a, AA22a, AA20a, AK21, BZ22a, Gha22, IS20, IMR21, KGF20, LW22a, LWC22, PN21, RYA20, SCK21, TEM20, VW21, VG22, YA21, YW20]. Estimators [LFMMRH22, AF22a, AKS^{+}21, AM21, ACP21, AMK21b, AA22b, BZ22b, CK20, CCM20, DYC22, ESS22, GLT20, HZAF22, Hua20, Hut21, JM20, Kal21, Kur21c, LFZ21, MM20, MW20, ÖK21, PKW22, Ruz22, SZ21, SVM20, SU21, SK22b, SVK22, SBB22, WLC^{+}22]. Evaluating [PG22, DC20, Hom20]. Evaluation [HBM^{+}20, UEA20, Han21, KR21, LPP21, OBN22, SMWJ20, ZL20]. Evaluations [MW20, CL21, SN22]. event [ONE22]. events [AK20, BDD^{+}22, PRH21]. evidence [BT20, Kur21c]. EWMA [ASM22, HCTZ21, NN21, NKSY22, SAA21, TAA22]. Exact [GHW20, ZBZS20, Nag21a, SAF21]. example [BGM20, MMR22]. exceedance [Ere20]. existence [YÖT20]. exit [GHW20]. exp [AAHR20]. exp- [AAHR20]. expansions [BCDA21]. expectation [dXLY20a]. expected [Han21]. experiment [ZBZS20]. experimental [LFMMRH22, MMR22]. experiments [LAZ21, RCSS22, SPG^{+}22, SCHS22]. experimentwise [LAZ21]. experts [YWLD20]. exploitation [SN22]. exploratory [FBNRG21]. Exponential [LRF20, ASZN21, AE21, CLY21, GJK21, KR21, MJ20, PMK20, RL21, RK20, SBTP20, WL22, Wu21, MBBS22]. Exponential-Poisson [LRF20]. Exponentially [uARS22, AK20]. exponentiated [RBRAa21, Ras20]. expression [HYSC21]. Extended [AAHR20, Abd18, ASP22, CL22a, RCSS22]. extensions [MdN22]. extreme [BGM20, BE20, KGR21]. extremes [MdN22]. Extropy [NJ22]. eye [OP20]. Eyraud [dXLY20a, dXGLY20].

facing [FO22]. factor [CSG22, MPS21]. factorial [BVJV22, LAZ21]. factors [ML22]. failure [CSL20, NG22]. false [BR21]. families [Bar20a, BAEk21]. family [AAHR20, BFAR20, BPN22, CKK21, CCBJ21, HPZ^{+}21, MG21, MB21]. familywise [DC20, Hor22]. Farlie [dXLY20a, dXGLY20]. fast [LYP22, LW22a]. feature [CGS21, SX22]. features [TWTT20]. fiber [HORA22]. Fiducial [WZY^{+}21, CP22, YDLX20]. fields [HI22, Lev21]. filter [Lev21]. Filtering [DL20]. financial [AM21, CDB^{+}21, NN22, WS21a, ZT20]. finding [Tel22]. findings [ÖT20]. Finite [TJ20, BRA20, DKL22, KK22, LY20, PKW22, QL21, REMO22, SS21b]. finite-sample [PKW22]. finite-state [SS21b]. fires [TBH22]. First [LZ22, Sol21, TBH22]. First-order [LZ22, Sol21]. Fisher [HR22]. Fit [LK21b, ABHF22, BH20, CMA21, Dog22, FS22, Gai21, KGR21, MG21, MQ21, Mil20, NM21, OP22, PHY20, Sul22]. fitting [LK21b, MW20]. fixed [CHC20, Hua20]. flexible [BCR^{+}21, ROL^{+}20, WHW21]. Focused [HCZ21]. following [TF22]. forecast [AT21, Bey21, CL21, RM20]. Forecasting [De 20, EGVdR21, IRA21, WZ21b]. forecasts [XZJ^{+}21]. Forest [SLW20, MNL21]. forests [CGS21]. form [Lev21, LWC22, dALNPdOF21]. formula [Bar20a, SZJG21]. formulas [DL20]. Formulation [SBB22, RBN^{+}22]. four [WL20]. Fourier [dAIIS20]. fractional [LPP21, XW20]. frailty [PBLM22, fSLC21, ULN21, ZGC21]. frame [LFZ21]. framework [WHW21]. free [TK22a]. frequency [AM21, CL21]. frontier [RRA20]. Function [NCC^{+}22, AA22b, BCP21, CHC20, CZP21, Gho20, KSAR21, KE22, Lav21, LY20, MYVD21, Nad21, PKA22, PZ20, SS21a, SA22, SD22, ZNLW20, dOMD^{+}21]. Functional [ZHW21, BV22, CCH20, GP20, KS20b, Mos22, NSN21, Sha20, YZ21]. functions [AMK21b, AA20d, Bar20a, DSL20, Han20, Han21, KEE20, Pic21, Pol20, WN21b, YDHW21]. further [ÖK20]. fuzzy [KEE20, Tür20].

gains [AM21]. games [LX22, SZS20]. Gamma [NAG21b, AA20a, BJU20, KLK21, Kur21c, OP22, PKA22, Ras20, SAA21, VG22, WZY^{+}21, YH21a, ZK20a]. gap [SL21]. GARCH [Alt20a, De 20, DDK21, IT21, TKF^{+}20]. GARCH-TSLx [Alt20a]. Gaussian [CSDT20, DE22, HD20, Lev21, Nag21a, WCF21]. GEE [XWL22]. gene [HPZ^{+}21, HYSC21]. General [WN21a, AÖ20a, Baz20, MG21, TEM20, ZG21]. generalised [BJ20]. Generalization [Alt20b, GP20, JKH20, dOMA21]. Generalized [BE20, GNE21, LYZ^{+}21, Pan20a, SPS22, YDLX20, ARJ21, ARJ20, ABJ21a, BAEk21, BFAR20, CKK21, DA21, FZM21, GKM20, HK20, KLK21, LH20, MK21, Özk21, OP22, PM20, RAJ21, REMO22, RCSS22, RK20, SBTP20, TY21, TNN20, VLvdW22, WLT22, WL22, Wes20, ZWY20, ZD22]. generally [CLS22]. Generating [NCC^{+}22, Gho20, MS22]. Generation [KTZ20, BVJV22, VWZ21]. Genetic [ES22, PHN22, VPT21]. Geometric [Alt20b, Biç20, NAKK20, PAT22, PIA22, RL21, dOMA21]. geostatistical [DRB20, HD20]. Gibbs [AA20c, OKO21]. Gini [SKS21]. glass [HORA22]. glaucoma [OP20]. global [ES22]. GLS [GG22]. GM [DA21, CSX20, CS22]. GMDH [DKKA22]. GMDH-type [DKKA22]. GMWB [OO21]. goal [AES21]. Gompertz [CL22a, Abd18, GABG22]. good [EFD21]. Goodness [ABHF22, CMA21, KGR21, NM21, OP22, PHY20, FS22, Gai21, LK21b, MG21, MQ21, Mil20, Sul22]. Goodness-of-fit [CMA21, KGR21, NM21, LK21b, FS22, Gai21, Mil20, Sul22]. Gradient [Yas22, DKL22, LYP22]. graph [HvZ21]. graphical [Kur21c]. graphs [HPZ^{+}21]. gray [CS22]. grey [CSX20]. gross [WLY22]. Group [SX22, CKT^{+}20, GK22a, GK22b, ULN21]. grouped [DP20]. growth [MD22]. GSTAR [IRA21]. GSTAR-SUR [IRA21]. GSTAR-SUR/neural [IRA21]. Guaranteed [HCTZ21]. Gumbel [dXLY20a, dXGLY20]. GWMA [HU20].

Half [SMB22, RBRAa21]. Half-spectral [SMB22]. Handling [PRH21]. having [DSL20, SKG22]. hazard [AMK21b, PM20, SKG22]. hazards [CMA21]. head [OP20]. health [SCE^{+}20]. heavy [GI20, QYC22, dXLY20a, dXGLY20]. heavy-tailed [GI20]. hedonic [DE22]. Heligman [SRO20]. heritability [VLvdW22]. Heston [MFS21]. heterogeneity [Har21]. heterogeneous [Nad21]. heteroscedastic [MAAA21, YH21b]. heteroscedasticity [Uya22, Wu21, WCF21, ZGL21]. heteroskedastic [AS21b]. HEWMA [JuAH20]. Hierarchical [AY21a, CFM21]. High [MT21, AM21, CZP21, CL21, CKCG21, FY21, FS22, IAG21, JM20, Kur21a, Kur21b, LYY21, Nak21, VLvdW22, WX21, WCF21]. high-dimension [CZP21]. High-dimensional [MT21, CKCG21, FY21, IAG21, JM20, LYY21, Nak21, VLvdW22]. high-frequency [AM21, CL21]. Higher [MVLMA22, SN22, Bap20, BPP20]. Higher-order [MVLMA22]. highly [KS20a]. Hill [GCF^{+}20]. histograms [DP20]. HIV [SKG22]. Hjorth [YBC22]. homogeneity [ATYK20, GZY22, PD21]. homoscedasticity [FO22]. Horvitz [IMR21]. Hotelling [MLR^{+}22, SSBA22]. house [DE22]. HPD [LPNP20]. humidity [KKO21]. hybrid [CSL20, CLY21, IRA21, KB21, MBBS22, NG22, RBRAa21, SBTP20, WLT22, WL22, ARAB21]. hyper [KR22]. hyper-Poisson [KR22]. hyperbolic [ARJ21, GKM20]. hyperkinetic [IM20]. hypotheses [ZK20b]. Hypothesis [ZD22, NdSA20, PD21, YV20].

I/II [Yad22]. IBNR [TK22b]. ICBayes [PCWL22]. identifiability [RBN^{+}22]. Identification [AZH^{+}22, SI20, LMP20, TEM20]. identifying [PY20, ZT20]. ignorable [RSG21]. II [AMK21b, CKK21, GAS20, KB21, KP21, LTR21, MG21, MK21, MJ20, PV20a, REN22, SBTP20, Sha22, Yad22, YBC22, YHGA21]. image [Lev21]. images [OP20]. imbalanced [Nak21]. Impact [SCE^{+}20, UKK20, CSE^{+}21, GZY22, SBZG22]. implementation [CKCG21, JPH21]. implied [BH20]. importance [NT20]. improve [SPG^{+}22]. Improved [CS22, MM20, PDO^{+}20, BZ22b]. Improvement [LPNP20, CSX20, ZK20a]. Improving [CL22b, LW21, PY20, SN22, Lev21, XWL22]. imputation [BPP20, FH22, SKK21]. in-play [SS21b]. In}-type [CK20]. inadequate [CSE^{+}21]. INAR [WZ21b, Lar20, LZ22, SB20, SM20]. incentives [WMJ21]. incidence [CHC20]. incomplete [KJK20, MPD20]. Inconsistency [KM20]. incorrect [CNL21]. increase [GAS22]. independence [KuIK20]. independent [ACP21, Gho20]. index [AA20b, ABJ21b, ARKA22, BT20, CS20, GZ22a, GZ22c, Nad21, NN21, PDO^{+}20, RAJ21, Sha20, ZNLW20, ZGL21, ZHW21, ZFG22, ZXF22]. indicator [Özk21]. indicators [MMR22]. indices [ANP20, OP20, Sol21, SKS21]. individual [LAZ21]. individuals [JAP21]. industry [YC22]. infant [SKG22]. infected [JAP21]. Inference [ASP22, HR22, Hut21, MJ20, RBRAa21, SBTP20, Sha22, WX21, WLT22, WL22, Alt20a, AHG20, BC21, Biç20, CHC20, FH22, FP20, GABG22, HK20, IT21, KJ20, LW22b, MdN22, MB21, PIA22, SB21a, SCE^{+}20, SKS21, SLE20, WZY^{+}21, YDLX20, Zha22, ZB20, ZBZS20]. Inferences [AS22, SRO20]. inferiority [CP22]. infinite [Che21, GTW22]. Inflated [RSG21, ASM22, CFM21, GZY22, KR22, SZ20]. Influential [KUA21, MAMA20, AZH^{+}22, SI20]. information [AS21a, BSNT21, HB22, HR22, HCZ21, HHY21, IMR21, JuAH20, Kim21, NKSY22, Pol22, SVM20, SVK22, XH21]. information-weighted [HHY21]. initial [FPAA21]. INLA [Ver21]. innovation [SM20, WMJ21, YH21b]. innovations [GKM20, SKJ20, TKF^{+}20]. inputs [SLW20]. inspection [RGA21, WZ21a]. instrumental [ZZWH20]. insurance [BS21, DA22]. integer [AES21, WWYX21, OB21]. integer-valued [WWYX21, OB21]. integrated [PM20, SNA20]. integration [Bap20, EL21, XW20]. intensity [NG22, fSLC21]. inter [TDD20]. inter-rater [TDD20]. interarrival [PIA22]. intercept [AJ22]. interest [GHW20, LG21, XZJ^{+}21]. interesting [EFD21]. interpolation [HI22]. interquantile [ACP21]. interquartile [SA22]. Intersection [AÖ20b]. Interval [ACP21, AHG20, CST21, LPP21, MZ22, MAMA20, MB21, PD21, SZ20, SB21b]. intervals [BCP21, CNL21, DP20, IRS20, IR21, LPNP20, MNN21, PDO^{+}20, TNN20, UYMK22, WJB21]. intervel [PCWL22]. intervel-censored [PCWL22]. intervened [KS20c]. intra [ATYK20]. intra-class [ATYK20]. Intuitionistic [KEE20]. inventory [MNL21]. inverse [AA20d, CSDT20, WL20]. inverted [CKK21]. Investigation [GZ21a, PKW22, Kur21c]. Iranian [SMB22]. irregularly [Mos22]. issue [SK22b]. Item [Brz20, Mag20, UM22]. iterative [MW20]. Iteratively [Zhe21].

Jackknife [SKS21, WY21, WLY22]. Jeffreys [MK22]. Jensen [LK21b]. Jensen-Shannon [LK21b]. Johnson [MM20, PY20]. joint [GGB20, HPZ^{+}21, KKO21]. Jones [HM21]. jump [LLLL22]. jumps [MFS21, WS21a, ZZ22].

kaphom [ATYK20]. kappa [ATYK20]. Kenward [CLL20]. kernel [Ali22, BZ22a, HZAF22, PK21, Pol22]. known [AZH^{+}22]. Kolmogorov [ZZFG20]. kriging [ASM21]. Kumaraswamy [KB21, dALFB21, WLT22]. kurtosis [EHHS20, PFM21].

label [HvZ21]. lag [FPAA21, MAAA21, ÖT20, SKG22]. Lambert [NCC^{+}22]. landmark [SCK22]. language [Pol22]. Laplace [ZBZS20]. Laplacian [HvZ21, LH20]. large [CL22c, HvZ21, PV20b, ZC22]. large-sample [PV20b]. lasso [AA20c, HPZ^{+}21, EGVdR21, ULN21, WJZ^{+}22]. LASSO/adaLASSO [EGVdR21]. late [Yad22]. late-onset [Yad22]. latency [RdAIS^{+}20]. latent [HK20]. lattice [EFD21]. learning [GZ21b, LYP22, RCSS22, SPG^{+}22]. least [DRB20, KLM^{+}22, WY21, WLY22, ZG22b, Zhe21]. left [MAMA20, MB21]. left-truncated [MAMA20]. Legendre [LJC^{+}20]. length [BC21, FPAA21, HCTZ21]. Leone [AMK21b]. level [AY21b, BVJV22, CL22b, FS22, LAZ21, NS20]. levels [TK22b]. Leveraging [SAD^{+}21]. Lévy [JHR21, OO21]. life [ABPK21, BJU20, CSL20, DR22, GAS20, PZ20, SBTP20, SZJG21, ZLSC20]. lifetime [Abd18, Ami21, CL22a, PAT22]. lifetimes [Wu21]. Likelihood [TWTT20, BA22, CLL20, Che21, DDK21, DKL22, GNE21, HD20, HK20, KAVR20, LK21b, LWC22, LLL22, LY20, MQ21, MM20, OKO21, ÖA22, SZ21, SKS21, TKF^{+}20, TY21, YH21b, ZGC21, ZZWH20, ZBZS20]. Likelihood-based [TWTT20, LLL22, SKS21]. Lilliefors [Sul22]. limit [XXH22]. limitations [AW22]. limits [TAA22]. Lin [Mil20, PHY20]. Lin-Wong [PHY20]. Lindley [Biç20, WZ21b, GABG22, GW22, PBLM22, Sha22]. Linear [GAS22, YW20, ZG21, AG20, AÖ20a, AA22a, AZH^{+}22, AES21, AK21, ALVQ22, BJ20, BSNT21, CCM20, DLRZ21, EGVdR21, FY21, FP20, GI20, GZ22b, GZ22a, GLT20, Gha22, HK20, HCZ21, IRS20, KSAR21, KN22, KJK20, KÖ21a, KÖ21b, Lee20, LMP20, MYVD21, Nag21a, Nis21, ÖA22, ÖK20, Özk21, RAG21, RBN^{+}22, Ric21, RCSS22, RVV22, REN22, SA22, SFG21, SCK21, TY21, TÇA22, Tür20, VLvdW22, WX21, WCF21, YA21, YRAB22, YZ21, ZNLW20, ZGL21, ZD22, ZZWH20, ZXF22]. link [Nad21]. liquid [GHW20]. Liu [AÖ20a, AA20a, Gha22, KÖ21b, ÖK20, VW21]. Liu-type [AA20a]. local [Nis21]. location [AG21, Hut21, Kal21, MMMG21, MB21, PKW22, RYA20, STdL21, WLY22]. Log [MSL21, GGMM22, KJK20, KP21, Kur21c, LJC^{+}20, MAMA20, MB21]. log-density [LJC^{+}20]. log-gamma [Kur21c]. log-location-scale [MB21]. Log-normal [MSL21, MAMA20]. log-ratio [GGMM22]. logarithmic [SBB22]. logistic [IAG21, IS20, Mil21, NN20, ONE22, OBN22, PJL21, RBRAa21, SAD^{+}21, UYMK22, dOVQP21, VW21, ZZ21]. logit [DSL20]. lognormal [MNN21, PY20]. long [AS21b, FZZ20, ROL^{+}20]. long-term [FZZ20, ROL^{+}20]. Longitudinal [NSN21, ALVQ22, GI20, GZ21a, GGB20, HCZ21, RSG21, SWC^{+}20, TY21, TWTT20, TWTT21, XWL22, ZZWH20]. Loss [Zhe21, ARKA22, CZP21, Han20, Han21, dXLY20a, dXGLY20, dXLY20b]. low [CZP21]. low-sample-size [CZP21]. luck [GKZ^{+}22]. LVQ [LW21].

M [BC21, BC21, DA21]. M/M/1 [BC21]. Machine [RCSS22, Nak21, SPG^{+}22]. machines [FN20, MT21]. maintenance [NG22, PM20, SNA20, SPG^{+}22]. Makeham [CL22a, Abd18]. making [MM22]. Malaysia [NN22]. Mansson [ÖT20]. MAR [CLL20]. Mardia [EHHS20]. margin [ZC22]. Marginal [KÖ21a, Mil21, SL21, NSN21, PRH21, ZGL21]. marginalized [OKO21]. marginals [ADK21, ZWY20]. markers [ANP20]. market [CL21, YC22, ZT20]. Markov [BCP21, HE21, KBR21, Lev21, LES21, OKO21, SS21b, SLE20, TMZL21, Yas22]. Markov-switching [OKO21]. Markov-type [Yas22]. martingale [Lee20, LPP21]. masked [CSL20]. Masking [SCHS22]. matched [CP22, WWTM22]. matched-pairs [CP22]. matches [BT20, SS21b]. matching [KLK21, OBN22, ZL22]. maternal [SKG22]. matrices [MS22, vO21]. matrix [Ava21, AM22, RAG21, WS21b]. matters [Lav21]. Max [JuAH20]. Max-HEWMA [JuAH20]. maxima [MBDK21, RGA21]. Maximum [BA22, HD20, CLL20, DDK21, DKL22, GNE21, KAVR20, LK21b, LWC22, MM20, SZ21, TKF^{+}20, ZNKN22]. Maxwell [SAF21]. MCEM [TWTT20]. MCMC [OKO21]. MCMC4Extremes [MdN22]. MD [TEM20]. Mean [KZG21, AH20, AK22, BJU20, Baz20, BPP20, DC20, DLRZ21, EL21, GK22a, Han21, HU20, HB22, KGF20, KBR21, MZ22, MNN21, MSS20, MYVD21, Pap20, PFM21, PKA22, PZ20, RAG21, SZ20, SMM22, SMO21, SKV20b, SKV20a, SVK22, SBB22, SPS22, SSK22, Wu21]. means [Bar20a, FO22, LW21, MJ20, MNN21, MES21, PD21]. measure [Bap22, ESS22, LK21b, MK22, SSK22, dXLY20b, ZG22a]. measurement [Brz20, CR21, Gha22, KZG21, uARS22, PZ20, SCE^{+}20, SVK22, SBB22, YDLX20, YA21, YRAB22, ZGC21, ZG21, ZFG22, ZGL22]. measurements [AATD21]. measures [Dog22, MAMA20, TDD20, Tür20, UR22]. Measuring [GWZC22, SLW20]. mechanics [CHL22]. mechanism [RSG21]. media [KKZ21]. median [CST21, GK22b, HCTZ21]. mediator [PHS21]. memory [AS21b]. MEPDF [WN21b]. mesh [Ver21]. meta [Avc21, Har21]. meta-analysis [Avc21, Har21]. method [Ali22, DA21, GW20, LW21, Med22, MBDK21, MK22, ÖT20, ÖK20, Pan20c, PK21, SAF21, SD22, SZJG21, SWC^{+}20, SY22, UKK20, UEA20, WP22, WS20, WY21, WLY22, YWT22, ZT20, ZWY21]. methodology [XWL22]. methods [AÖ20a, BFAR20, BPP20, CHC20, CSE^{+}21, CS22, CCBN22, CMMV21, CL22b, EGVdR21, HZAF22, IM20, KLM^{+}22, Kur21a, Kur21b, MES21, Nis21, OLOK22, OBN22, ÖA22, Ras20, SKK21, WL20, YÖT20, ZGC21]. metric [CSG22]. MEWMA [SSBA22]. Mexican [SRO20]. MGR [CKT^{+}20]. MIDAS [XZJ^{+}21]. Minimal [AATD21, Ami21]. minimization [JPH21]. minimizing [CCM20]. minimum [AM22, BVJV22, BZ22b]. Mis [SBZG22]. Mis-specification [SBZG22]. misclassification [dOVQP21]. Mises [BZ22a, ZZFG20]. missing [GAS22, PRH21, RSG21, SKK21]. misspecification [GZY22]. misspecifying [ZL20]. mix [AMK^{+}21a, CSE^{+}21, SCE^{+}20, XWL22]. mix-GEE [XWL22]. Mixed [AES21, AY21a, ALB22, AW22, ALVQ22, Gha22, GGB20, HK20, KÖ21a, KÖ21b, NSN21, ÖK20, RSG21, RBN^{+}22, TWTT20, TWTT21, WHW21, YA21, YRAB22]. mixed-effects [AW22]. mixing [LY20]. mixture [BAEk21, CLL20, DA21, KK22, LH20, LES21, Nad21, OB21, QL21, REMO22, RCSS22, TJ20, WJB21, YWLD20]. mixtures [CKCG21, Kim21, LFMMRH22, TJ20]. MNAR [CLL20]. modal [KSAR21]. mode [Ruz22]. Model [ALVQ22, Bar20b, HD22, HE21, TKF^{+}20, AG20, AÖ20a, AA22a, AZH^{+}22, ALB22, AA20a, Alg20, AK21, Alt20a, AJ22, AHG20, BC21, BSNT21, CSL20, CLL20, CSDT20, CSX20, CS22, CCM20, CMA21, CNL21, DE22, DA22, DSL20, DYC22, DY22, Dog22, DNWG22, EH22, FZM21, GZ22b, GZ22a, GWZC22, GZZ22, GHW20, GZY22, GJK21, HCZ21, IRA21, IRS20, JAP21, JHR21, JM20, KSAR21, KL21, KK22, KÖ21a, Kur21c, Lee20, LW22a, LLL22, LG21, MNL21, MAMA20, MFS21, MVLMA22, MD22, MES21, NN20, dALNPdOF21, NG22, NAG21b, OBN22, ÖT20, Pan20a, RRA20, ROL^{+}20, RBN^{+}22, RCSS22, SB20, SNA20, SI20, Sha20, SL21, SMWJ20, SS21b, SM20, SRO20, SLW20, SCK22, fSLC21, SCK20, SCK21, SKJ20, TMZL21, UKK20, ULN21, dOVQP21, WS20, WLC^{+}22, WMJ21, XXH22, dXLY20b, XZJ^{+}21, XH21, YDLX20, YV20, YA21]. model [ZHW21, Zha22, ZXF22, dSdSTC21]. model* [YW20]. model-based [MNL21, SCK22]. modeled [OO21]. Modeling [BS21, CL21, PHN22, SKG22, YC22, AT21, CDB^{+}21, Dog22, DA21, DKL22, GI20, GGB20, KJK20, SWC^{+}20, TJ20, Tür20, Ver21]. Modelling [SZS20, AA20d, GZ21a, WHW21]. models [AW22, ALVQ22, AE21, BCDA21, BJ20, BH20, BCR^{+}21, BE20, Brz20, CR21, CST21, CFM21, De 20, DLRZ21, EGVdR21, FY21, FZZ20, FS22, FP20, GZ22b, GZ22a, GWZC22, GZ22c, GZZ22, GNE21, GLT20, Gha22, HK22, HD20, Hep21a, HK20, HCZ21, HE21, HHY21, IAG21, IT21, IRA21, JGGDMRLD21, KKO21, KZG21, KÖ21a, KÖ21b, LLZ22, LES21, LMP20, LW22b, Mag20, MAAA21, MZZF21, MPS21, Mos22, Nad21, Nag21a, NSN21, NIN20, ONE22, OKO21, OB21, ÖA22, ÖK20, Özk21, PHN22, PBLM22, PV20a, PM20, RVV22, SB21a, SKV20a, SLE20, TY21, TEM20, TF22, TWTT20, TWTT21, TEO20, UM22, VLvdW22, WP22, WWYX21, WX21, WJB21, WCF21, XWL22, YRAB22, YWLD20, ZNLW20, ZZFG20, ZGL21, ZGC21, ZG21, ZD22, ZG22b, ZFG22, ZZWH20, ZZ22, ZK20b, dSdSTC21]. modes [GPP21]. Modification [Ali22, Med22]. Modified [BSNT21, Kur21c, SAA21, Sul22, VW21, Alg20, ABPK21, CKT^{+}20, DSL20, HYSC21, KUA21, NG22, RYA20, SZ21, SSK22, TK22a]. modulated [DL20]. moment [BPP20, Che21, Gho20, MQ21]. moments [LPP21, WN21a]. monitor [HORA22, SMM22]. Monitoring [BNK21, PV20a, ASM22, EAJ20, KS20a, dALFB21, MT21, MMMG21, MASA22, PASB21, RAG21, RYA20, SMM22, TAA22, YHGA21]. monotone [CLJK21]. monotonicity [RVV22]. Monte [CL22b, FPAA21, KKZ21, KOT20, uIKKH22, Kur21c, LPNP20, Uya22, WMS22, ZL20]. Morgenstern [dXLY20a, dXGLY20]. mortality [SRO20]. Most [KuIK20, uIKKH22, Uya22]. moving [AK20, CLS22, LLZ22, NZ21, uARS22]. moving-average [LLZ22]. MSE [Han21]. Multi [BDD^{+}22, KS22, REN22, Avc21, Bey21, EAJ20, Gai21, MFS21]. multi-asset [MFS21]. multi-category [Avc21]. Multi-objective [BDD^{+}22, REN22].

Multi-response [KS22]. multi-sample [Gai21]. multi-stage [EAJ20]. multi-step [Bey21]. Multicategory [ZC22]. multicollinearity [SCK21, ZZ21]. multidimensional [UM22]. multilevel [FS22]. multinomial [DSL20, Esk21, LW22a]. Multiple [DF22, ARJ20, ABPK21, ABJ21a, ABJ21b, BDD^{+}22, CSG22, GK20, KS22, RAJ21, SS20a, TWTT20, WHW21, Wes20, Wu21, YDLX20, ZT20, ZHW21]. multiple-index [ZHW21]. multiplicative [BZ22a, HZAF22]. multipliers [GW20]. multistage [BNK21, KN22]. Multivariate [EHHS20, WN21b, ARJ21, BNK21, Baz20, CKL22, Dog22, FN20, HZAF22, IT21, JM20, Kal21, KS20b, LYZ^{+}21, MVLMA22, MV22, MSS20, NAG21b, Pan20c, Par20, RAG21, UEA20, dXLY20b, YHGA21]. Multiwavelet [CZS21].

Nadaraya [Ali22, ESES20]. national [MNL21]. NBA [SS21b, SZS20]. nearest [FH22]. Necessary [WWTM22]. need [dOVQP21]. negative [GPP21, KS20c, WWYX21, ZWY20]. negatively [DYC22]. neighbors [FH22]. neighbour [BJVV21]. nerve [OP20]. nested [EH22]. nested-error [EH22]. net [LYY21, Pol22]. network [DKKA22, IRA21, IMR21]. networks [GZ21b, HPZ^{+}21, KE22, MASA22, SMWJ20]. neural [DKKA22, IRA21, KE22]. neutrosophic [ARKA22]. newly [TK22a]. noise [AS21b, DDK21, ZK20a]. noisy [MZZF21]. nominal [UR22]. nomination [RGA21]. Non [Bar20b, Sol21, AATD21, ASZN21, ANP20, CP22, DE22, FN20, FO22, KKO21, Pan20c, RSG21, SA22, SU21, SK22b, SBB22, WCF21]. non-Bayesian [ASZN21]. non-circular [AATD21]. non-Gaussian [DE22, WCF21]. non-ignorable [RSG21]. non-inferiority [CP22]. non-linear [SA22]. non-normal [FN20, Pan20c]. non-normality [FO22]. Non-Parametric [Bar20b, Sol21, ANP20]. non-response [SU21, SK22b]. non-stationary [DE22, KKO21]. noncontractual [YWT22]. Nonlinear [ARJ21, CL21, ZG22b, AES21, GG22, Hep21a, Hep21b, NSN21, Tür20]. nonlinearity [AE21]. Nonnegative [LYY21]. Nonparametric [BZ22a, CDB^{+}21, GJK21, GK20, HvZ21, LFZ21, SZ20, SK22a, SD22, YWT22, Ali22, AG21, AY21b, CLS22, DYC22, GZZ22, Gho20, KK22, Ruz22]. nonrandom [AJ22]. nonresponse [KJK20]. nonsampled [KK22]. normal [BJU20, BAEk21, Baz20, BFAR20, CSDT20, Dog22, FN20, FZM21, GZ21a, HR22, HK20, KAVR20, LES21, LTR21, MSL21, MAMA20, MES21, NZ21, ON21, PHN22, Pan20c, PD21, SS20b, SS20a, TJ20, TNN20, WWTM22]. Normality [PV20b, Bay21, EHHS20, FO22, MQ21, Par20, Sul22, WMS22]. normalizing [EHHS20]. normally [IRS20]. note [AW22, BJVV21, DLRZ21, GW22, Gho20, Mag20, RL21, SVM20, TMZL21, WLC^{+}22, YV20]. Notes [Lui22, ZL20]. np} [LK21a]. NSGA [REN22]. NSGA-II [REN22]. nuisance [GZY22]. null [uIKKH22]. number [JAP21, Tel22]. numbers [Tür20]. Numerical [CKK21, KLM^{+}22, BV22, ÖT20, SZJG21].

object [Hor22]. objective [BDD^{+}22, KE22, REN22]. observation [AZH^{+}22]. observational [ZL22]. observations [PV20a, QYC22, SI20]. occasion [SKK21, SK22b, SPS22]. occurrence [LRF20]. odds [CP22]. old [AKS^{+}21]. oncology [Yad22]. One [Wu21, BPN22, KJK20, Mis20, RYA20, RAJ21]. one-sided [RAJ21]. one-step [Mis20, RYA20]. one-truncation [BPN22]. Online [XH21]. onset [Yad22]. operating [FTC21]. operators [SM20]. optic [OP20]. optical [KKZ21]. Optimal [CLY21, CKT^{+}20, Hor22, Mis20, AJ22, BJVV21, EAJ20, GJK21, HI22, LFMMRH22, MPD20, NN20, PJL21, YHGA21]. optimality [BPP20]. optimization [ASM21, BM21, CS22, CCM20, ES22, IS20, SI20, YÖT20, dOMD^{+}21]. optimum [SBTP20, ZLSC20]. option [LLLL22, MFS21]. optional [KZG21, NS20]. order [ARJ21, AAA20, Bap20, Baz20, BPP20, DRB20, GPP21, GBHA20, HR22, LZ22, MVLMA22, MES21, NJ22, Sol21, dXLY20b]. ordered [AG21, Bey21]. ordinal [AA20b, GZ21a, GGB20, PG22, RSG21, TDD20]. ordinary [MZZF21]. organizing [CCH20]. orthogonal [EH22, GGMM22]. orthogonality [ZZWH20]. other [DP20, Lui22]. outcome [BT20, ZD22]. outcome-dependent [ZD22]. outcomes [SWC^{+}20, TDD20, Yad22]. Outlier [AMK^{+}21a, UEA20]. outliers [FBNRG21, SCK21, Tel22, TEO20, WCF21, YÖT20].

package [ATYK20, FBNRG21, JPH21, MdN22, MM22, PCWL22]. Padé [AY21b]. paired [SCHS22]. pairs [AGA20, CP22]. pairwise [Lev21]. panel [AY21a, AÖ20b, DNWG22, FZM21, Nad21]. Parallel [SCK22, Lui22]. Parameter [MFPP21, ASZN21, AA22a, Alg20, BPN22, CCBJ21, CSX20, CS22, CCM20, FLPB20, GW22, Han21, IRS20, KN22, KLM^{+}22, KP21, Kur21c, LK21a, MJ20, ÖT20, OP22, PJL21, SAF21, Sha22, TK22a, VG22, Biç20]. parameter-free [TK22a]. parameters [BA22, BPN22, CKT^{+}20, GZY22, HCTZ21, JHR21, KB21, LTR21, Mag20, MK21, MM20, MZZF21, SCK20, YHGA21]. Parametric [Bar20b, MSS20, ANP20, BCR^{+}21, DRB20, dALNPdOF21, QL21, SB20, Sol21, TF22, YZ21]. parametrization [vO21]. Pareto [DA22, JKH20]. Parisian [CHL22]. part [YZ21]. Partial [ZFG22, AG20, GZ22b, GZ22a, IMR21, KSAR21, Pol20, WS20, YZ21, ZNLW20, ZGL21]. partially [AG20, GCF^{+}20, HCZ21, RVV22, ZZWH20, ZXF22]. Particle [IS20, ASM21, RY21, SI20, dOMD^{+}21]. path [PN21]. patient [DF22]. pattern [CLL20]. patterns [Ger21, PK21, SPS22]. PBIBDs [SG21]. PBINAR [MPS21]. PCA [AMK^{+}21a]. PDF [Ras20]. Pearson [YWLD20]. Pegram [SM20]. Pena [KUA21]. Penalized [CYS21, CLJK21, LJC^{+}20, TY21, IAG21, KLM^{+}22, ONE22, YDHW21]. penalty [VLvdW22]. perfect [AA22b]. Performance [CSE^{+}21, KR21, OBN22, AG21, BS21, EAJ20, HCTZ21, NN20, ÖK21, Pol20, SZS20, UEA20]. Performances [Kur21a, Kur21b]. periodic [LG21, OB21, SB20].

Permutation [MBDK21]. personalized [KS22]. perturbed [LG21]. PH [SI20]. phase [MD22, PV20a, Yad22, YHGA21, KN22, YHGA21]. Phase-I [KN22]. phenomena [MD22, TK22a]. piecewise [MD22]. PINAR [KL21]. plan [ABJ21a, ABJ21b, ARKA22, MS20, PMK20, RAJ21]. planar [Ger20]. plans [CLY21, DR22, NN21, RGA21, SBTP20, ZLSC20]. play [SS21b]. PM [GZ21b]. Point [MG21, BSNT21, CCBJ21, CHC20, EFD21, LY20, Mil21, Pic21, QYC22, Ver21]. points [BDD^{+}22, ZT20]. Poisson [AK22, ASM22, CFM21, DL20, GZY22, KR22, LRF20, MBBS22, MV22, NM21, PASB21, SKJ20, WZ21b]. policy [WMJ21]. Pollard [SRO20]. polynomial [ZHW21]. polynomials [DY22, LJC^{+}20]. population [BPP20, BRA20, CK20, IR21, KGF20, KK22, MZ22, MYVD21, NS20, PD21, SZ20, SKV20b, SKV20a, SBB22, SSK22]. populations [Baz20, dAIIS20, RY21, TNN20]. Portfolio [NN22, AM22, BM21, dXLY20a, dXGLY20, dXLY20b]. positive [Bap22]. positives [BR21]. post [ZL22, Bar20a]. post-matching [ZL22]. Posterior [KLK21, Han20]. Power [BGM20, HM21, WMS22, CYS21, Lui22, Pan20c, RL21, RSG21, RK20, SSS20, Uya22]. powerful [DC20]. pre [SMO21]. pre-timed [SMO21]. precipitation [IRA21]. Precision [AM22, Ava21, GAS22, WY21]. Predicting [BT20]. Prediction [KK22, BJ20, BC21, CNL21, GAS22, HI22, HvZ21, MBBS22, ÖK20, SS21b, SZJG21, WS20, WL20]. Predictive [HK22, DA22, KÖ21a, LLL22, SPG^{+}22, ZBZS20]. predictor [SAD^{+}21]. predictors [CZ22, HAT^{+}22, KÖ21b]. Preface [FP20]. presence [MS20, PZ20, RAG21, RGA21, SKV20a, SU21, SBB22, SCK21]. preserving [HD22]. pressure [BT20, KKO21]. Pretesting [FO22]. prevalence [Avc21]. prevention [KAVR20]. preventive [PM20]. prices [DE22]. Pricing [CHL22, MFS21, OO21, LLLL22]. Principal [FN20, Kur21b, ZNKN22]. prior [SK22a, UKK20]. priors [ALVQ22, KLK21]. privacy [HD22, SSK22]. privacy-preserving [HD22]. probabilities [CYS21, IRS20, KR21, Mil21, NZ21]. probability [Bar20a, KTZ20, KGF20, ON21, TAA22]. problem [Gai21]. problems [AA20d, Che21, CCM20, HBM^{+}20, KLM^{+}22, SKK21, WZY^{+}21, YÖT20]. procedure [AW22, DC20, OM21, SS20a]. procedures [CCBJ21, HE21, KB21, WZ21b]. process [AH20, ABJ21b, ARKA22, BGM20, Biç20, CSDT20, CKL22, CKT^{+}20, GK22a, GK22b, GBHA20, HU20, HB22, HCTZ21, HE21, KKZ21, LK21a, LLLL22, LZ22, MMMG21, PG22, PDO^{+}20, PKA22, PIA22, RAJ21, STdL21, SBZG22, WZ21b, WJB21, YH21a, ZWY20]. processes [ASM22, BNK21, BCP21, Bey21, DRB20, DL20, EAJ20, GKM20, HORA22, JHR21, KN22, PV20a, PAT22, Pic21, RdAIS^{+}20, VWZ21, Ver21, XXH22]. Procrustes [GP20]. product [SVK22]. production [KS20a, YC22]. products [MSL21]. profile [PASB21, RAG21, ZGC21, ZG22b]. profiles [BNK21, KN22, MV22, RAG21, REN22, YHGA21]. profiling [CSE^{+}21, SCE^{+}20]. programming [AES21]. progressive [AK22, AHG20, CSL20, CKK21, KB21, LTR21, MK21, SBTP20, Sha22, WLT22, WL22, YBC22]. progressively [MG21, MJ20]. projected [NAG21b]. projection [BV22, FBNRG21]. Propagation [BCDA21]. propensity [HAT^{+}22, OBN22]. Properties [AAHR20, Alt20b, ASZN21, CZ22, DKL22, KLK21, KS20c, PKW22, WLC^{+}22]. property [XXH22]. proportion [NS20, SK22a]. proportional [CMA21, MW20]. proportions [UYMK22]. proposal [AA22a]. Proposition [AF22b]. protection [SSK22]. providers [SCE^{+}20]. Pseudo [TKF^{+}20]. psychology [MMR22]. psychometric [DKL22]. pure [RdAIS^{+}20]. pure-birth [RdAIS^{+}20]. pursuit [FBNRG21].

quadratic [YÖT20]. Quality [KS20a, SNA20, SN22, ZK20a]. Quantile [MSL21, SCK20, SCK21, AA20b, CKCG21, GGB20, KM20, LMP20, SD22, TWTT20, TWTT21, WJZ^{+}22, ZLSC20]. Quantile-based [MSL21, SCK21]. quantiles [ABHF22, ACP21, DP20, ESES20]. quantitative [SKV20b, SKV20a, SSK22]. Quanto [LLLL22]. Quasi [DDK21, CL22b, EL21, GNE21, LPP21, LWC22, NT20, SS20b, YH21b]. quasi-likelihood [YH21b]. Quasi-maximum [DDK21, LWC22]. quasi-Monte [CL22b]. quasi-random [EL21, NT20]. quasi-stationary [LPP21]. question [NS20]. questionnaire [SY22]. queue [BC21, EAJ20]. queueing [BCDA21]. queuing [BC21].

R [ATYK20, FBNRG21, JPH21, MdN22, MM22, XW20]. R/S [XW20]. R/S-bootstrapping [XW20]. radiation [KKZ21]. Rahim [PM20]. rainfall [LRF20]. random [AJ22, Bar20a, BJVV21, CK20, CGS21, CKCG21, EH22, EL21, FZZ20, GLT20, GZ21a, HI22, Hua20, Lev21, LZ22, MS22, MASA22, MYVD21, NT20, NZ21, RY21, Sha22, SK22b, SBB22, SLW20, WJZ^{+}22, YW20, ZL20]. random-effects [GLT20]. randomization [JPH21]. Randomized [NT20, BV22, NS20, SKV20b, SKV20a, SSK22, Yad22]. range [SA22, SS20b]. ranges [ACP21]. rank [AÖ20b, KE22]. ranked [AS22, AA22b, BRA20, MZ22, MLR^{+}22, Pan20b, STdL21, SZ21, VDN22]. ranking [AA22b, MK22]. rare [ONE22]. rate [AMK21b, NG22, PK22, XZJ^{+}21]. rater [TDD20]. rates [BR21, Hor22, LAZ21, SKG22]. Ratio [ESS22, QYC22, BZ22b, CP22, Che21, GGMM22, KGF20, MQ21, ON21, PN21, SVK22, WP22]. ratio-cum-regression [KGF20]. ratios [ACP21]. Rayleigh [ABHF22, HORA22, MK21]. RCINAR [ZWY20]. re [TF22]. re-estimation [TF22]. real [GNE21, SPG^{+}22]. reassessment [UKK20, ZWY21]. reconstructions [OP20]. recovered [OP20]. rectifying [RGA21]. recurrent [PRH21, SL21]. red [Özk21]. reduced [GCF^{+}20]. reduced-bias [GCF^{+}20]. reduction [KM20, ZPW21]. refinements [MVLMA22]. regarding [ZL20]. region [LLL22]. regional [SLW20]. regions [Bey21]. Regression [AÖ20a, GGMM22, HHY21, WS20, AF22a, AA22a, AES21, AA20a, Alg20, AA20b, AK21, AKS^{+}21, Ali22, AJ22, ARJ21, AY21b, BSNT21, CR21, CFM21, CLJK21, CKCG21, DLRZ21, DYC22, DA21, DF22, ESS22, EH22, FLPB20, GI20, GWZC22, GZ22c, GZZ22, GAS22, GGB20, GJK21, IRS20, KUA21, KGF20, KM20, KEE20, Kur21a, Kur21b, Lee20, LLL22, LFZ21, MVLMA22, Mil21, NIN20, dALNPdOF21, Nis21, ONE22, OBN22, OM21, ÖA22, PCWL22, PV20a, PJL21, PRH21, Pol20, RRA20, ROL^{+}20, Ric21, RVV22, SGG22, SZ21, SL21, SAD^{+}21, SCK20, SCK21, TWTT21, TÇA22, UYMK22, ULN21, dOVQP21, VW21, WLC^{+}22, WCF21, YDLX20, YZ21, ZZ21, ZG21, ZG22b, Zhe21, ZLSC20, dSdSTC21]. regressions [CNL21]. regressors [AA22a]. regular [JAP21, LAZ21]. regularization [CZP21, HvZ21]. regularized [LH20]. regularly [dXLY20b]. related [CSG22]. relationship [AZH^{+}22, ZZ21]. relative [GZ22c, KKO21]. Reliability [Bar20b, AS22, AMK21b, Han20, MK21, MSL21, SMWJ20]. removals [Sha22]. Removing [PFM21]. renewal [PKA22]. repair [Ami21]. repeated [AATD21, dOVQP21]. repetitive [CLS22]. replacement [PM20]. replicated [LAZ21, Tür20]. replicates [Gho20]. REPPlab [FBNRG21]. Representativeness [VDN22]. representing [PK21]. resample [WY21]. Resampling [CNL21, NT20]. Resampling-based [CNL21]. Rescaling [BRA20]. reserves [GHW20]. reservoir [CL22c]. residual [KL21, PZ20, Ric21, SZJG21]. Residuals [ZZFG20, AG20, IRA21]. respect [CCM20]. response [AAA20, Brz20, DLRZ21, GAS22, KS22, Mag20, MASA22, NN20, NS20, SKV20a, SU21, SK22b, SBB22, Tür20, UM22]. responses [Esk21, GZ21a, LFMMRH22, PG22, RSG21]. restricted [AK21, Gha22, KÖ21b, LW22b, Pol22, TÇA22, XZJ^{+}21]. restriction [Baz20]. restrictions [YA21, YRAB22]. results [TKF^{+}20]. retrieval [Pol22]. returns [IT21]. Reverse [XZJ^{+}21]. reversed [HM21]. review [KGR21]. Revisiting [CMMV21]. reweighted [Zhe21]. rho [TK22a]. Ridge [YRAB22, AF22a, Alg20, AK21, AKS^{+}21, FLPB20, IS20, KUA21, KEE20, KÖ21a, Kur21c, PJL21, SGG22, SCK20, SCK21, VLvdW22, YA21]. ridge-type [AF22a]. Riesz [KTZ20]. right [CKK21, PV20a]. right-censored [PV20a]. Risk [ÖK21, GHW20, GCF^{+}20, Han20, KOT20, LG21, SD22, TAA22, XXH22, dXGLY20, dXLY20b, Alt20a]. risk-adjusted [TAA22]. risks [AHG20, CSL20, WLT22, WL22]. RMCriteria [MM22]. Robust [AG20, DA21, KS20b, Nak21, OM21, RVV22, SB21a, TÇA22, WJZ^{+}22, AKS^{+}21, ASM21, BZ22b, CKK21, KN22, NIN20, ÖA22, PKW22, Pol20, SFG21, SCK21]. Robustification [RYA20]. robustness [FO22, SPG^{+}22]. ROC [ANP20, HAT^{+}22]. role [ÖT20]. room [ASM21]. root [FPAA21, GG22, Hep21b]. rooted [JAP21]. rotated [IT21]. rotation [SPS22]. rounded [ZB20]. roundoff [WN21a]. RRT [KZG21]. RSM [REN22]. rules [CKL22, SMM22]. run [CKL22, HCTZ21]. runs [CKL22, CKT^{+}20, GK22a, GK22b, Ger21, SMM22]. runs-rules [SMM22].

S [SKS21]. S-bootstrapping [XW20]. S-Gini [SKS21]. Sample [CFM21, GK22a, CZP21, Che21, CMMV21, DKL22, Ere20, Gai21, GBHA20, GBHA21, GK20, KJ20, MK21, PKW22, PV20b, SS20b, TF22, WZY^{+}21, WWTM22, WJB21, WS21b, ZPW21]. sampled [Mos22]. sampler [AA20c]. samples [ASZN21, BRA20, CLY21, LY20, MG21, SMM22, Wu21]. sampling [AS22, AA22b, ARJ20, ABPK21, ABJ21a, ABJ21b, ARKA22, BJU20, CK20, CLY21, CLS22, CL22c, DR22, KGF20, LK21a, MNL21, MZ22, MS22, MMMG21, MS20, MYVD21, NN21, NT20, OKO21, Pan20b, PN21, PMK20, RAJ21, RGA21, STdL21, SZ21, SA22, SB21b, SKK21, SK22b, SBB22, VDN22, ZD22]. sampling/importance [NT20]. SARIMA [AT21]. Saunders [BA22]. scale [GK20, MG21, MB21, PKW22, TJ20, VG22]. scales [BH20]. scaling [CSG22]. scatter [Kal21]. scattering [KKZ21]. scheme [BJU20]. schemes [CKL22, MLR^{+}22, Mis20, SMM22, Yas22]. score [GNE21, Nag21a, OBN22, VDN22]. scores [HAT^{+}22, SZS20]. scramble [SKV20a]. screening [SX22]. Second [dXLY20b, Che21, DRB20]. Second-order [dXLY20b]. select [dALNPdOF21, Wes20]. selecting [SY22]. Selection [Ami21, AES21, Alg20, ALVQ22, AM22, AY21b, CGS21, CLL20, CCBN22, EGVdR21, FPAA21, KS22, KÖ21a, LYY21, NN22, REMO22, ULN21, WJZ^{+}22, XH21, ZZ22]. self [CCH20, HYSC21]. self-organizing [CCH20]. self-updating [HYSC21]. selling [DE22]. semantic [MMR22]. Semiparametric [FZZ20, ZT20, AG20, ALB22, CST21, DY22, Esk21, LW22a, LW22b, PCWL22, SB21a, SL21, WP22, WLC^{+}22]. sensitive [GK22a, GK22b, KZG21, SKV20b, SKV20a, SSK22]. Sensitivity [ALVQ22, FPAA21, NS20, RSG21]. separation [ZZ21, ZNKN22]. Sequential [CCBJ21, FTC21, ON21, AAA20, Lui22]. serially [LES21, SM20]. series [ADK21, AS21b, Bap20, Bap22, CL21, Gho20, GTW22, dAIIS20, KKO21, KuIK20, LZ22, MPS21, OLOK22, RL21, RSG21, RK20, SNA20, SLE20, WS21a, YH21b, dSdSTC21]. server [BCDA21, BC21]. service [EAJ20]. set [AS22, AA22b, BRA20, MZ22, MLR^{+}22, STdL21, SZ21, VDN22]. SETAR} [HK22]. sets [EFD21]. setting [WWTM22, XH21, YWT22]. seven [Uya22]. several [AE21, MJ20, PD21, SS20b, SS20a, TNN20, WMS22]. Shannon [LK21b]. shape [DY22, MMR22, TJ20]. shapes [SCK22]. shared [PBLM22]. Shewhart [SMM22]. Shewhart-type [SMM22]. shift [RYA20, WZ21a]. shifts [GK22a, GK22b]. Shiryaev [LPP21]. shock [BE20]. short [CKL22]. Shrinkage [Alg20, LMP20, VG22, FLPB20, LW21, ÖK21]. side [GK22a, GK22b]. sided [RAJ21]. Sigma [KR21]. Sigma-based [KR21]. sign [CLS22, SB21b]. Signal [TP22]. signals [SMO21]. signed [KE22]. signed-rank [KE22]. simple [BCR^{+}21, KN22, RAG21, REN22]. Simplex [CR21]. Simpson [SZJG21]. simulating [Pan20c]. Simulation [FPAA21, LLZ22, NAKK20, Ali22, AAHR20, AS21b, ASM21, BCR^{+}21, CMMV21, Dog22, GZY22, GKZ^{+}22, IM20, KKZ21, KOT20, Pol20, TEO20, WMS22]. simulation-based [CMMV21, IM20]. simulations [GNE21]. Simultaneous [Par20, RAG21, ZK20b]. Single [GZ22c, AA20b, BC21, GZ22a, MNN21, MS20, Nad21, Pap20, PHS21, RGA21, Sha20, ZNLW20, ZGL21, ZXF22]. Single-index [GZ22c, AA20b, GZ22a, Nad21, ZNLW20, ZGL21, ZXF22]. single-mediator [PHS21]. single-sampling [MS20]. SingleCross [Tel22].

SingleCross-clustering [Tel22]. singular [RM20]. SIR [JAP21]. situations [SK22b]. Six [KR21]. size [CFM21, CZP21, DKL22, Har21, TF22]. sizes [WWTM22]. Skew [FZM21, CSDT20, GZ21a, HR22, HK20, IT21, KAVR20, ON21, PHN22, TJ20, WWTM22]. Skew-normal [FZM21, CSDT20, HR22]. skew-normal-Cauchy [KAVR20]. skewed [BFAR20, RYA20]. Skewing [Nis21]. skewness [PFM21, Ric21]. skill [GKZ^{+}22]. slope [AZH^{+}22, IRS20]. Small [IC21, KJ20, ZPW21, MW20, MSS20, WJB21]. Smirnov [ZZFG20]. smooth [Hep21a, Hep21b]. smoother [AS21b]. smoothing [DL20]. Sobol [Sol21]. social [MASA22, SCHS22]. Software [Bar20b]. solution [SCHS22]. solutions [GHW20]. Solving [Ric21]. Some [AAA20, EFD21, GK22b, ÖT20, SKV20a, SKK21, WZ21b, AKS^{+}21, Bar20a, BAEk21, GHW20, KS20c, Kur21a, Kur21b, OBN22, ÖK21, PAT22]. space [JGGDMRLD21, MMR22, Nag21a]. Sparse [GW20, PK21, WCF21, Kur21b, Mos22, SFG21]. Spatial [dOMD^{+}21, BJ20, DE22, FZM21, HK20, LW22b, SB21a, SMB22, Ver21, WHW21]. spatial-temporal [SB21a, SMB22]. SPC [KBR21]. SPDE [Ver21]. Spearman [TK22a]. special [KS20a]. specification [CNL21, SBZG22]. specified [Mil21, WWTM22]. spectra [Pic21]. spectral [SMB22]. spectrum [RM20]. speed [SMB22]. sphere [VWZ21]. spillover [CS20]. spline [CLJK21, KSAR21, YZ21]. splines [YDHW21]. split [SY22]. spotting [TBH22]. spread [LWC22]. square [Han21, Zhe21]. squared [MSS20]. squares [DRB20, KLM^{+}22, WY21, WLY22, ZG22b]. stability [BV22]. stabilizing [Nis21]. Stage [Bar20b, ARKA22, EAJ20, KM20, NN20, OM21, Wu21]. standard [Han21]. state [ARJ20, ABPK21, ABJ21a, ABJ21b, JGGDMRLD21, Nag21a, RAJ21, SS21b]. stationary [DRB20, DE22, HI22, KKO21, LPP21]. statistic [BGM20, KUA21, KL21, MK22, SX22, TK22a]. Statistical [AHG20, Biç20, CHC20, LW22b, MB21, MASA22, PIA22, GKZ^{+}22, HE21, KL21, NKSY22, OP20, REN22, ZB20]. statistics [ATYK20, ARJ21, ARKA22, Ere20, GK20, HR22, ML22, NJ22, SZS20]. Step [SS20a, Bey21, CSL20, GAS20, Mis20, RYA20]. step-stress [CSL20]. Step-up [SS20a]. stereo [OP20]. stimuli [MMR22]. Stochastic [Gha22, JGGDMRLD21, KKO21, KOT20, KÖ21b, AK21, JAP21, KKZ21, LX22, LYP22, RRA20, TEM20, YA21, YRAB22, YC22]. stock [CL21]. Strategies [GAS22, FO22, SCK22]. strategy [LG21, SPS22]. Stratified [Pan20b, CK20, CL22c, MYVD21, SA22]. streams [CL22c]. strength [AS22, HORA22]. stress [AS22, CSL20, GAS20]. stress-strength [AS22]. stringent [KuIK20, uIKKH22]. strongly [AATD21]. structural [AZH^{+}22, De 20, Dog22]. structure [CZP21, JM20, PK21, RYA20, SAD^{+}21, UKK20, Wes20, ZL22, ZK20b]. Structured [GZ21b]. structures [GZ21a, ROL^{+}20]. student [ML22, SN22, BA22, DDK21, NIN20]. studies [TWTT21, XWL22, ZL22]. Study [FPAA21, Ali22, ANP20, AE21, BA22, Bay21, BCR^{+}21, CKK21, Dog22, GZY22, GKZ^{+}22, KOT20, KGR21, MNL21, MPS21, NN20, NN22, Pol20, Ras20, SPG^{+}22, SMB22, TEO20, WMS22, ZWY20]. Subgroup [ZA21]. subgroups [YHGA21]. subject [KJK20]. subsampling [LLZ22]. subset [AGA20, CKCG21]. subspace [PK21]. success [ML22]. successive [SKK21, SK22b, SBB22]. sum [CKCG21, LX22]. sums [NZ21]. superposition [RdAIS^{+}20]. supplementary [SMM22]. support [FN20, MT21, MM22, Nak21, Zhe21]. supported [WJB21]. SUR/neural [IRA21]. surface [AAA20]. surgery [ASM21]. survey [BJ20, KK22, PHS21]. surveys [IC21, LFZ21]. Survival [RY21, BCR^{+}21, CST21, FZZ20, LK21b, LYZ^{+}21, NCC^{+}22, TAA22, YBC22]. survivors [FZZ20, ROL^{+}20]. swaps [CHL22]. swarm [ASM21, IS20, SI20, dOMD^{+}21]. switching [OKO21, TMZL21]. symmetric [UEA20, VWZ21, ZK20b]. symmetry [GZ22b, GZ22a, GWZC22, GZZ22, Kal21, NJ22, ZG22a, ZGL22]. synthetic [NKSY22]. system [MM22, PY20, SZJG21]. systematic [MNL21]. systems [Ami21, KS20a, PM20, SNA20].

table [KJK20]. tail [KR21, dXLY20a, dXLY20b]. tailed [GI20, QYC22]. tails [BGM20, dXLY20a, dXGLY20]. taste [MMR22]. teaching [SN22]. technique [BRA20, SG21]. techniques [GBHA20, NS20]. temperature [KKO21]. tempering [SCK22]. temporal [SB21a, SMB22]. term [FZZ20, IRS20, ROL^{+}20]. Test [DY22, YZ21, AF22b, ABPK21, BGM20, Bay21, CSL20, CP22, DR22, DKL22, EHHS20, Ere20, FS22, GZY22, GAS20, GG22, HM21, Hep21b, KuIK20, uIKKH22, MQ21, MK22, ON21, PHY20, Sul22, TK22a, XW20, ZLSC20]. Testing [EH22, GZZ22, Hep21a, Lar20, Pap20, PD21, ZGL22, ATYK20, Baz20, Ger20, GK20, NJ22, PFM21, SBTP20, WWYX21, YV20, ZD22, ZK20b]. Tests [Bay21, ABHF22, AG21, AF22b, AÖ20b, AA22b, AE21, BPN22, CYS21, CL22b, CMA21, FPAA21, Hor22, KGR21, MG21, Mil20, NdSA20, NM21, OP22, Par20, PV20b, Uya22, WMS22, ZZFG20]. Their [FO22, BR21, BE20, LX22, NYM20]. theoretic [PMK20]. theory [Brz20, Mag20, TF22, UM22]. there [AM21]. thick [BGM20]. thinning [SM20, WWYX21]. third [AAA20]. Thompson [IMR21]. Three [DR22, ASZN21, OP22]. three-parameter [ASZN21, OP22]. threshold [LLZ22, WWYX21, YH21b]. thresholding [SD22]. thyroid [HAT^{+}22]. tightened [BJU20]. tightened-normal-tightened [BJU20]. time [AK20, ADK21, AS21b, Bap20, Bap22, BCR^{+}21, CHC20, CL21, Gho20, GTW22, GZ21b, KKO21, KuIK20, MAMA20, MZZF21, MPS21, NCC^{+}22, OLOK22, SKG22, SS21b, SLE20, TBH22, TAA22, WS21a, YH21b, ZZ22, ZBZS20, dSdSTC21]. time-constrained [ZBZS20]. time-dependent [BCR^{+}21, MAMA20]. time-series [KKO21]. time-varying [GZ21b, MZZF21, NCC^{+}22, ZZ22]. timed [SMO21]. times [CST21, GHW20, NCC^{+}22, PIA22, SL21]. Toeplitz [JM20]. Too [Mil20].

Topp [AMK21b]. TOPSIS [REN22]. total [WY21, WLY22]. toxicity [YH21a]. Trace [Ava21]. traditional [AF22b]. traffic [SMO21]. transfer [KKZ21]. transformation [CST21, EHHS20, PFM21, SSS20, SL21]. transformations [EH22]. transformed [Pan20c, Pap20, SAA21]. transition [Hep21a]. transitions [Hep21b]. transmission [KM20, SKG22]. treatment [IM20, KS22, MPD20, ZA21]. treatments [KS22]. tree [MES21, SWC^{+}20]. trees [JAP21]. trend [Gho20, JGGDMRLD21, PKA22]. trial [ZL20]. trials [Med22, Yad22]. trivariate [dOMA21]. truncated [HvZ21, LTR21, MAMA20, MB21]. Truncation [AY21b, BPN22]. TSLx [Alt20a]. Tukey [MFPP21]. Tukey-type [MFPP21]. tumor [RdAIS^{+}20]. tuning [KLM^{+}22]. Twenty20 [BT20]. Two [ARKA22, Bar20b, AA22a, BPN22, CST21, Che21, DNWG22, Ere20, HZAF22, IM20, dAIIS20, IC21, KM20, KJK20, LAZ21, MG21, MJ20, MNN21, MD22, OM21, OP22, SNA20, Sha22, SKG22, SKK21, SK22b, SPS22, UYMK22, WZY^{+}21, YH21a, Biç20]. two-agent [YH21a]. two-dimensional [DNWG22]. two-level [LAZ21]. two-occasion [SKK21, SK22b, SPS22]. two-parameter [AA22a, MJ20, OP22, Sha22, Biç20]. two-phase [MD22]. two-sample [Che21, WZY^{+}21]. Two-Stage [Bar20b, ARKA22, KM20]. two-unit [SNA20]. two-way [KJK20]. Type [AS22, AF22a, AA20a, AJ22, CK20, CKK21, DKKA22, ESS22, GPP21, GAS20, KB21, LTR21, MK21, MBBS22, MFPP21, PV20a, PK22, SZ21, Sha22, SMM22, SBB22, YBC22, Yas22, YWLD20, AMK21b, CYS21, CLY21, KP21, MG21, MJ20, PHY20, PMK20, SBTP20]. type-I [MBBS22, CLY21, PHY20, PMK20]. type-II [KB21, MK21, YBC22, AMK21b, KP21, MG21, SBTP20]. types [WHW21].

unbiased [AK21, DRB20, HM21, VW21]. unblinded [TF22]. uncertainty [BCDA21]. Unconstrained [vO21]. underlying [Kal21]. unequal [KGF20, ZC22]. unidimensional [UM22]. unified [LLL22]. unit [BH20, FPAA21, GG22, Hep21b, SNA20]. unit-weighted [BH20]. units [KK22]. univariate [Pan20c]. Universal [Med22]. unknown [HCTZ21, Sha20]. unrelated [NS20]. unreliable [BCDA21]. updating [HYSC21, XH21]. use [SN22, Wes20]. used [Lui22, Uya22]. useful [ANP20, OM21]. Using [Pol22, AS21a, AMK^{+}21a, ARAB21, AM21, ARJ21, AMK21b, ABPK21, ARKA22, AY21b, ASM21, BCDA21, BJ20, Bey21, BT20, BPP20, BCR^{+}21, BZ22b, CGS21, EH22, EHHS20, FN20, FBNRG21, GZZ22, Gho20, GGMM22, GK20, HAT^{+}22, HZAF22, HvZ21, Hut21, dAIIS20, JuAH20, JHR21, KR21, KZG21, KBR21, KK22, Lee20, LJC^{+}20, MT21, MV22, NS20, NIN20, NCC^{+}22, NG22, uARS22, OBN22, OP20, PHY20, RAJ21, REN22, SSS20, SZ21, SRO20, SS20b, SKV20b, SVM20, SVK22, TJ20, TK22b, WZ21a, YWLD20, dOMD^{+}21]. utilization [Wes20].

validation [Alg20]. value [BGM20, CS22, CL22b, GCF^{+}20, KGR21, PKA22, dXGLY20, Alt20a]. value-at-risk [GCF^{+}20, dXGLY20, Alt20a]. valued [WWYX21, OB21]. VAR [Bey21, GTW22]. variability [PK22, SB21b]. Variable [EGVdR21, REMO22, ULN21, AY21a, AES21, ARJ20, BPP20, GAS22, KJK20, LYY21, uARS22, RAJ21, SKG22, SB21b, SKV20a, SSK22, dOVQP21, WZ21a, WJZ^{+}22, WCF21, ZG22a, ZZWH20]. variables [Bar20a, CGS21, HK20, KZG21, NN21, NZ21, WS20, WX21]. variance [AM22, BRA20, CK20, CCM20, Gho20, GTW22, Har21, KM20, KBR21, Nis21, SK22b, VLvdW22, WZ21a, WS21b, YW20]. variance-stabilizing [Nis21]. variances [MNL21, SS20b, SS20a]. variant [SVK22]. variation [CKL22, IR21, KJ20, TNN20]. Variational [Zha22]. variograms [DRB20]. various [SU21]. varying [ESES20, GWZC22, GZ21b, Hor22, HCZ21, LW22b, MZZF21, Mos22, NCC^{+}22, TEO20, dXLY20b, ZZFG20, Zha22, ZZWH20, ZZ22, ZXF22]. varying-coefficient [HCZ21, Zha22, ZXF22]. vector [FN20, Med22, Nak21, RAG21, RM20, Zhe21]. vectors [Baz20, Nag21a]. version [AK21, HBM^{+}20, PM20]. versus [GCF^{+}20]. via [Alg20, CZP21, CDB^{+}21, EGVdR21, HHY21, LW21, LYY21, ML22, RGA21, SD22, SX22, SLE20, TWTT20, YDHW21, YC22]. VIFs [Özk21]. VII [YWLD20]. virtual [NG22, OP20]. volatility [CS20]. volumes [DF22]. vs [OKO21]. VSI [WZ21a].

walks [RY21]. warping [Hut21]. Watson [Ali22, ESES20, NdSA20]. wavelet [AA20d, DYC22, dALNPdOF21, PASB21]. wavelet-based [AA20d]. wavelets [OO21]. way [KJK20]. weak [GBHA20, TEM20]. Weibull [ARAB21, Ami21, ABPK21, AHG20, ASP22, BJU20, DA22, KGR21, KP21, NG22, PV20a, PIA22]. weight [Pol20]. Weighted [NAKK20, NYM20, TWTT21, TDD20, AK20, BH20, CLS22, CKCG21, DRB20, ESES20, FH22, GW22, HHY21, Med22, uARS22, WY21, WLY22]. weighting [MES21]. where [Lav21]. WHIM [Lav21]. Widder [Bar20a]. Wiener [Lev21, SBZG22]. wildfires [Ver21]. wildland [TBH22]. wind [SMB22]. window [FTC21]. windows [De 20]. within [RAG21, Wes20]. without [Ric21, TP22]. Wong [PHY20]. working [Wes20]. WQS [CKCG21]. Wynn [Bar20a]. Wynn- [Bar20a].

X [AS22].

yield [NN21].

zero [ASM22, CFM21, GZY22, KR22, LX22, LRF20, SZ20]. zero-inflated [ASM22, CFM21, GZY22, KR22, SZ20]. zero-sum [LX22]. zeros [TEO20].

References

Algamal:2020:LTE

[AA20a] Zakariya Yahya Algamal and Yasin Asar. Liu-type estimator for the gamma regression model. Communications in Statistics: Simulation and Computation, 49(8):2035–2048, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2020:BSI

[AA20b] Rahim Alhamzawi and Haithem Taha Mohammad Ali. Bayesian single-index quantile regression for ordinal data. Communications in Statistics: Simulation and Computation, 49(5):1306–1320, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2020:NGS

[AA20c] Rahim Alhamzawi and Haithem Taha Mohammad Ali. A new Gibbs sampler for Bayesian lasso. Communications in Statistics: Simulation and Computation, 49(7):1855–1871, 2020. CODEN CSSCDB. ISSN 0361-0918.

Aykroyd:2020:BAW

[AA20d] Robert G. Aykroyd and Hassan Aljohani. A Bayesian approach to wavelet-based modelling of discontinuous functions applied to inverse problems. Communications in Statistics: Simulation and Computation, 49(1):207–225, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2022:APA

[AA22a] Shakeel Ahmad and Muhammad Aslam. Another proposal about the new two-parameter estimator for linear regression model with correlated regressors. Communications in Statistics: Simulation and Computation, 51(6):3054–3072, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ashour:2022:NDF

[AA22b] Samir K. Ashour and Mohamed S. Abdallah. New distribution function estimators and tests of perfect ranking in concomitant-based ranked set sampling. Communications in Statistics: Simulation and Computation, 51(3):823–848, 2022. CODEN CSSCDB. ISSN 0361-0918.

Arshad:2020:SST

[AAA20] Hafiz Muhammad Arshad, Tanvir Ahmad, and Munir Akhtar. Some sequential third order response surface designs. Communications in Statistics: Simulation and Computation, 49 (7):1872–1885, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alizadeh:2020:EEF

[AAHR20] Morad Alizadeh, Mahmoud Afshari, Bistoon Hosseini, and Thiago G. Ramires. Extended exp-G family of distributions: Properties, applications and simulation. Communications in Statistics: Simulation and Computation, 49(7):1730–1745, 2020. CODEN CSSCDB. ISSN 0361-0918.

Abdullah:2021:MNC

[AATD21] Muhammad Abdullah, Rashid Ahmed, M. H. Tahir, and Muhammad Daniyal. Minimal non-circular balanced and strongly balanced repeated measurements designs. Communications in Statistics: Simulation and Computation, 50(10): 2837–2844, 2021. CODEN CSSCDB. ISSN 0361-0918.

AbdEl-Bar:2018:EGM

[Abd18] Ahmed M. T. Abd El-Bar. An extended Gompertz–Makeham distribution with application to lifetime data. Communications in Statistics: Simulation and Computation, 47(8):2454– 2475, 2018. CODEN CSSCDB. ISSN 0361-0918. See comments [CL22a].

Ahrari:2022:GFT

[ABHF22] V. Ahrari, S. Baratpour, A. Habibirad, and V. Fakoor. Goodness of fit tests for Rayleigh distribution based on quantiles. Communications in Statistics: Simulation and Computation, 51(2):341–357, 2022. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2021:DED

[ABJ21a] Muhammad Aslam, Saminathan Balamurali, and Chi-Hyuck Jun. Determination and economic design of a generalized multiple dependent state sampling plan. Communications in Statistics: Simulation and Computation, 50(11):3465–3482, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2021:NMD

[ABJ21b] Muhammad Aslam, Saminathan Balamurali, and Chi-Hyuck Jun. A new multiple dependent state sampling plan based on the process capability index. Communications in Statistics: Simulation and Computation, 50(6):1711–1727, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Aslam:2021:DAC

[ABPK21] Muhammad Aslam, Saminathan Balamurali, Jeyadurga Periyasamypandian, and Nasrullah Khan. Designing of an attribute control chart based on modified multiple dependent state sampling using accelerated life test under Weibull distribution. Communications in Statistics: Simulation and Computation, 50(3):902–916, 2021. CODEN CSSCDB. ISSN 0361-0918.

Arachchige:2021:IER

[ACP21] Chandima N. P. G. Arachchige, Maxwell Cairns, and Luke A. Prendergast. Interval estimators for ratios of independent quantiles and interquantile ranges. Communications in Statistics: Simulation and Computation, 50(12):3914–3930, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alqawba:2021:CDD

[ADK21] Mohammed Alqawba, Norou Diawara, and Jong-Min Kim. Copula directional dependence of discrete time series marginals. Communications in Statistics: Simulation and Computation, 50(11):3733–3750, 2021. CODEN CSSCDB. ISSN 0361-0918.

Azouagh:2021:DEC

[AE21] Nabil Azouagh and Said El Melhaoui. Detecting exponential component in autoregressive models: comparative study between several tests of nonlinearity. Communications in Statistics: Simulation and Computation, 50(11):3273–3285, 2021. CODEN CSSCDB. ISSN 0361-0918.

Albassam:2021:MIN

[AES21] Mohammed Albassam, Ali El Hefnawy, and Emad E. A. Soliman. Mixed integer nonlinear goal programming approach to variable selection in linear regression. Communications in Statistics: Simulation and Computation, 50(12):4028–4040, 2021. CODEN CSSCDB. ISSN 0361-0918.

Abdel-Fattah:2022:NCB

[AF22a] Mahmoud A. Abdel-Fattah. On a new class of binomial ridgetype regression estimators. Communications in Statistics: Simulation and Computation, 51(6):3272–3290, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Alves:2022:PNA

[AF22b] Henrique J. P. Alves and Daniel F. Ferreira. Proposition of new alternative tests adapted to the traditional T^{2} test. Communications in Statistics: Simulation and Computation, 51 (5):2287–2300, 2022. CODEN CSSCDB. ISSN 0361-0918.

Abonazel:2020:RPR

[AG20] Mohamed Reda Abonazel and Ahmed Abd-Elfatah Gad. Robust partial residuals estimation in semiparametric partially linear model. Communications in Statistics: Simulation and Computation, 49(5):1223–1236, 2020. CODEN CSSCDB. ISSN 0361-0918.

Altunkaynak:2021:CPN

[AG21] Bulent Altunkaynak and Hamza Gamgam. Comparing the performance of nonparametric tests for equality of location against ordered alternatives. Communications in Statistics: Simulation and Computation, 50(1):63–84, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2020:CAP

[AGA20] Tanvir Ahmad, Steven G. Gilmour, and Hafiz Muhammad Arshad. Comparisons of augmented pairs designs and subset designs. Communications in Statistics: Simulation and Computation, 49(7):1898–1921, 2020. CODEN CSSCDB. ISSN 0361-0918.

Abbasi:2020:NAC

[AH20] Saba Abbasi and Abdul Haq. New adaptive CUSUM charts for process mean. Communications in Statistics: Simulation and Computation, 49(11):2944–2962, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Azizi:2020:SIC

[AHG20] Fariba Azizi, Firoozeh Haghighi, and Niloofar Tabibi Gilani. Statistical inference for competing risks model under progressive interval censored Weibull data. Communications in Statistics: Simulation and Computation, 49(7):1931–1944, 2020. CODEN CSSCDB. ISSN 0361-0918.

Aminnejad:2022:BOD

[AJ22] Minoo Aminnejad and Habib Jafari. Bayesian optimal designs for Cox regression model with random and nonrandom intercept based on type I censored data. Communications in Statistics: Simulation and Computation, 51(2):583–603, 2022. CODEN CSSCDB. ISSN 0361-0918.

Alevizakos:2020:DEW

[AK20] Vasileios Alevizakos and Christos Koukouvinos. A double exponentially weighted moving average chart for time between events. Communications in Statistics: Simulation and Computation, 49(10):2765–2784, 2020. CODEN CSSCDB. ISSN 0361-0918.

Alheety:2021:NVU

[AK21] Mustafa I. Alheety and B. M. Golam Kibria. A new version of unbiased ridge regression estimator under the stochastic restricted linear regression model. Communications in Statistics: Simulation and Computation, 50(6):1589–1599, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alevizakos:2022:PMC

[AK22] Vasileios Alevizakos and Christos Koukouvinos. A progressive mean control chart for COM-Poisson distribution. Communications in Statistics: Simulation and Computation, 51(3): 849–867, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ali:2021:CSN

[AKS^{+}21] Sajid Ali, Himmad Khan, Ismail Shah, Muhammad Moeen Butt, and Muhammad Suhail. A comparison of some new and old robust ridge regression estimators. Communications in Statistics: Simulation and Computation, 50(8):2213–2231, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alao:2022:ESM

[ALB22] Virgelio M. Alao, Joseph Ryan G. Lansangan, and Erniel B. Barrios. Estimation of semiparametric mixed analysis of covariance model. Communications in Statistics: Simulation and Computation, 51(5):2301–2317, 2022. CODEN CSSCDB. ISSN 0361-0918.

Algamal:2020:SPS

[Alg20] Zakariya Yahya Algamal. Shrinkage parameter selection via modified cross-validation approach for ridge regression model. Communications in Statistics: Simulation and Computation, 49(7):1922–1930, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ali:2022:MAN

[Ali22] Taha Hussein Ali. Modification of the adaptive Nadaraya– Watson kernel method for nonparametric regression (simulation study). Communications in Statistics: Simulation and Computation, 51(2):391–403, 2022. CODEN CSSCDB. ISSN 0361-0918.

Altun:2020:NAV

[Alt20a] Emrah Altun. A new approach to Value-at-Risk: GARCH-TSLx model with inference. Communications in Statistics: Simulation and Computation, 49(12):3134–3151, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Altun:2020:NGG

[Alt20b] Emrah Altun. A new generalization of geometric distribution with properties and applications. Communications in Statistics: Simulation and Computation, 49(3):793–807, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Ariyo:2022:MSB

[ALVQ22] Oludare Ariyo, Emmanuel Lesaffre, Geert Verbeke, and Adrian Quintero. Model selection for Bayesian linear mixed models with longitudinal data: Sensitivity to the choice of priors. Communications in Statistics: Simulation and Computation, 51(4):1591–1615, 2022. CODEN CSSCDB. ISSN 0361-0918.

Allaj:2021:AAH

[AM21] Erindi Allaj and Maria Elvira Mancino. On asset-allocation and high-frequency data: are there financial gains from using different covariance estimators? Communications in Statistics: Simulation and Computation, 50(12):4413–4441, 2021. CODEN CSSCDB. ISSN 0361-0918.

Avagyan:2022:PME

[AM22] Vahe Avagyan and Xiaoling Mei. Precision matrix estimation under data contamination with an application to minimum variance portfolio selection. Communications in Statistics: Simulation and Computation, 51(4):1381–1400, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Amini:2021:SBM

[Ami21] Morteza Amini. Selection of the best minimal repair systems for Weibull lifetime distribution: a Bayesian approach. Communications in Statistics: Simulation and Computation, 50 (3):832–853, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ahsan:2021:ODU

[AMK^{+}21a] Muhammad Ahsan, Muhammad Mashuri, Heri Kuswanto, Dedy Dwi Prastyo, and Hidayatul Khusna. Outlier detection using PCA mix based T^{2} control chart for continuous and categorical data. Communications in Statistics: Simulation and Computation, 50(5):1496–1523, 2021. CODEN CSSCDB. ISSN 0361-0918.

Arora:2021:BER

[AMK21b] Sangeeta Arora, Kalpana K. Mahajan, and Ritu Kumari. Bayes estimators for the reliability and hazard rate functions of Topp–Leone distribution using Type-II censored data. Communications in Statistics: Simulation and Computation, 50(8):2327–2344, 2021. CODEN CSSCDB. ISSN 0361-0918.

Alonso:2020:SIU

[ANP20] Rosa Alonso, Christos T. Nakas, and M. Carmen Pardo. A study of indices useful for the assessment of diagnostic markers in non-parametric ROC curve analysis. Communications in Statistics: Simulation and Computation, 49(8):2102–2113, 2020. CODEN CSSCDB. ISSN 0361-0918.

Acar:2020:RDM

[AÖ20a] Tugba Söküt Açar and M. Revan Özkale. Regression diagnostics methods for Liu estimator under the general linear regression model. Communications in Statistics: Simulation and Computation, 49(3):771–792, 2020. CODEN CSSCDB. ISSN 0361-0918.

Arsova:2020:ITC

[AÖ20b] Antonia Arsova and Deniz Dilan Karaman Örsal. Intersection tests for the cointegrating rank in dependent panel data. Communications in Statistics: Simulation and Computation, 49(4):918–941, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ali:2021:CHD

[ARAB21] Sajid Ali, SyedMuhammad Muslim Raza, Muhammad Aslam, and Muhammad Moeen Butt. CEV-Hybrid Dewma charts for censored data using Weibull distribution. Communications in Statistics: Simulation and Computation, 50(2):446–461, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2020:NVC

[ARJ20] Muhammad Aslam, Muhammad Ali Raza, and Chi-Hyuck Jun. A new variable control chart under generalized multiple dependent state sampling. Communications in Statistics: Simulation and Computation, 49(9):2321–2332, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Amiri:2021:NRU

[ARJ21] M. Amiri, R. Roozegar, and A. Jamalizadeh. Nonlinear regression using order statistics from the multivariate generalized hyperbolic distributions. Communications in Statistics: Simulation and Computation, 50(5):1348–1363, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Aslam:2022:TSS

[ARKA22] Muhammad Aslam, Gadde Srinivasa Rao, Nasrullah Khan, and Liaquat Ahmad. Two-stage sampling plan using process loss index under neutrosophic statistics. Communications in Statistics: Simulation and Computation, 51(6):2831–2841, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ahmed:2021:ECD

[AS21a] Aneel Ahmed and Javid Shabbir. On estimation of coefficient of dispersion using the auxiliary information. Communications in Statistics: Simulation and Computation, 50(11): 3590–3606, 2021. CODEN CSSCDB. ISSN 0361-0918.

Asai:2021:SSL

[AS21b] Manabu Asai and Mike K. P. So. A simulation smoother for long memory time series with correlated and heteroskedastic additive noise. Communications in Statistics: Simulation and Computation, 50(2):388–399, 2021. CODEN CSSCDB. ISSN 0361-0918.

Akgul:2022:ISS

[AS22] Fatma Gül Akgül and Birdal Senoglu. Inferences for stressstrength reliability of Burr Type X distributions based on ranked set sampling. Communications in Statistics: Simulation and Computation, 51(6):3324–3340, 2022. CODEN CSSCDB. ISSN 0361-0918.

Azizi:2021:RSO

[ASM21] Mohammad Javad Azizi, Farshad Seifi, and Samira Moghadam. A robust simulation optimization algorithm using kriging and particle swarm optimization: Application to surgery room optimization. Communications in Statistics: Simulation and Computation, 50(7):2025–2041, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aly:2022:AEC

[ASM22] Aya A. Aly, Nesma A. Saleh, and Mahmoud A. Mahmoud. An adaptive EWMA control chart for monitoring zero-inflated Poisson processes. Communications in Statistics: Simulation and Computation, 51(4):1564–1577, 2022. CODEN CSSCDB. ISSN 0361-0918.

Azizi:2022:IAB

[ASP22] Afsaneh Azizi, Abdolreza Sayyareh, and Hanieh Panahi. Inference about the bivariate new extended Weibull distribution based on complete and censored data. Communications in Statistics: Simulation and Computation, 51(3):738–756, 2022. CODEN CSSCDB. ISSN 0361-0918.

Afify:2021:TPE

[ASZN21] Ahmed Z. Afify, Adriano K. Suzuki, Chunfang Zhang, and Mazen Nassar. On three-parameter exponential distribution: properties, Bayesian and non-Bayesian estimation based on complete and censored samples. Communications in Statistics: Simulation and Computation, 50(11):3799–3819, 2021. CODEN CSSCDB. ISSN 0361-0918.

Arlt:2021:ASM

[AT21] Josef Arlt and Peter Trcka. Automatic SARIMA modeling and forecast accuracy. Communications in Statistics: Simulation and Computation, 50(10):2949–2970, 2021. CODEN CSSCDB. ISSN 0361-0918.

Albayrak:2020:PKR

[ATYK20] Muammer Albayrak, Kemal Turhan, Yasemin Yavuz, and Zeliha Aydin Kasap. kaphom: an R package for testing the homogeneity of intra-class kappa statistics. Communications in Statistics: Simulation and Computation, 49(12):3283–3298, 2020. CODEN CSSCDB. ISSN 0361-0918.

Avagyan:2021:DTE

[Ava21] Vahe Avagyan. D-Trace estimation of a precision matrix with eigenvalue control. Communications in Statistics: Simulation and Computation, 50(4):1231–1247, 2021. CODEN CSSCDB. ISSN 0361-0918.

Avci:2021:BAM

[Avc21] Esin Avci. Bayesian approach to the meta-analysis of multicategory prevalence. Communications in Statistics: Simulation and Computation, 50(5):1541–1559, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Ananda:2022:NLC

[AW22] Malwane M. A. Ananda and Samaradasa Weerahandi. A note on the limitations of the CAT procedure with application to mixed-effects models. Communications in Statistics: Simulation and Computation, 51(6):3138–3148, 2022. CODEN CSSCDB. ISSN 0361-0918.

Akay:2021:HCM

[AY21a] Özlem Akay and Güzin Yüksel. Hierarchical clustering of mixed variable panel data based on new distance. Communications in Statistics: Simulation and Computation, 50(6): 1695–1710, 2021. CODEN CSSCDB. ISSN 0361-0918.

Aydin:2021:TLS

[AY21b] Dursun Aydin and Ersin Yilmaz. Truncation level selection in nonparametric regression using Padé approximation. Communications in Statistics: Simulation and Computation, 50 (3):744–763, 2021. CODEN CSSCDB. ISSN 0361-0918.

AlMamuno:2022:IIO

[AZH^{+}22] Abu Sayed Md. Al Mamuno, Yong Zulina Zubairi, Abdul Ghapor Hussin, A. H. M. Rahmatullah Imon, Sohel Rana, and Jalmar Carrasco. Identification of influential observation in linear structural relationship model with known slope. Communications in Statistics: Simulation and Computation, 51(1):72–83, 2022. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2022:MLE

[BA22] Narayanaswamy Balakrishnan and Farouq Mohammad A. Alam. Maximum likelihood estimation of the parameters of Student’s t Birnbaum–Saunders distribution: a comparative study. Communications in Statistics: Simulation and Computation, 51(3):793–822, 2022. CODEN CSSCDB. ISSN 0361-0918.

Barakat:2021:SGF

[BAEk21] H. M. Barakat, A. W. Aboutahoun, and N. N. El-kadar. On some generalized families arising from mixture normal distribution with applications. Communications in Statistics: Simulation and Computation, 50(1):198–216, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bapat:2020:NCB

[Bap20] Sudeep R. Bapat. A new correlation for bivariate time series with a higher order of integration. Communications in Statistics: Simulation and Computation, 49(10):2546–2558, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bapat:2022:AMP

[Bap22] Sudeep R. Bapat. An alternative measure of positive correlation for bivariate time series. Communications in Statistics: Simulation and Computation, 51(6):3252–3258, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Barabesi:2020:CPD

[Bar20a] Lucio Barabesi. The computation of the probability density and distribution functions for some families of random variables by means of the Wynn-ρ accelerated Post–Widder formula. Communications in Statistics: Simulation and Computation, 49(5):1333–1351, 2020. CODEN CSSCDB. ISSN 0361-0918.

Barghout:2020:TSN

[Bar20b] May Barghout. A two-stage non-parametric software reliability model. Communications in Statistics: Simulation and Computation, 49(5):1159–1180, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bayoud:2021:TNN

[Bay21] Husam Awni Bayoud. Tests of normality: new test and comparative study. Communications in Statistics: Simulation and Computation, 50(12):4442–4463, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Bazyari:2020:GTO

[Baz20] Abouzar Bazyari. A general testing for order restriction on mean vectors of multivariate normal populations. Communications in Statistics: Simulation and Computation, 49(8): 1982–1998, 2020. CODEN CSSCDB. ISSN 0361-0918.

Basak:2021:BIP

[BC21] Arpita Basak and Amit Choudhury. Bayesian inference and prediction in single server M/M/1 queuing model based on queue length. Communications in Statistics: Simulation and Computation, 50(6):1576–1588, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bachi:2021:PEU

[BCDA21] Katia Bachi, Cédric Chauvière, Hacène Djellout, and Karim Abbas. Propagation of epistemic uncertainty in queueing models with unreliable server using chaos expansions. Communications in Statistics: Simulation and Computation, 50 (4):1039–1061, 2021. CODEN CSSCDB. ISSN 0361-0918.

Barbeito:2021:BCI

[BCP21] Inés Barbeito, Ricardo Cao, and Dimitris Politis. Bootstrap confidence intervals for conditional density function in Markov processes. Communications in Statistics: Simulation and Computation, 50(12):4315–4337, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Bower:2021:CSC

[BCR^{+}21] Hannah Bower, Michael J. Crowther, Mark J. Rutherford, Therese M.-L. Andersson, Mark Clements, Xing-Rong Liu, Paul W. Dickman, and Paul C. Lambert. Capturing simple and complex time-dependent effects using flexible parametric survival models: a simulation study. Communications in Statistics: Simulation and Computation, 50(11):3777–3793, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bodevan:2022:MOA

[BDD^{+}22] Emerson C. Bodevan, Luiz H. Duczmal, Anderson R. Duarte, Pedro H. L. Silva, and Gladston J. P. Moreira. Multi-objective approach for multiple clusters detection in data points events. Communications in Statistics: Simulation and Computation, 51(3):1313–1332, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bozbulut:2020:GES

[BE20] Ali Riza Bozbulut and Serkan Eryilmaz. Generalized extreme shock models and their applications. Communications in Statistics: Simulation and Computation, 49(1):110–120, 2020. CODEN CSSCDB. ISSN 0361-0918.

Beyaztas:2021:CMS

[Bey21] Beste H. Beyaztas. Construction of multi-step forecast regions of VAR processes using ordered block bootstrap. Communications in Statistics: Simulation and Computation, 50 (7):2107–2125, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bekker:2020:CMA

[BFAR20] A. Bekker, J. T. Ferreira, M. Arashi, and B. W. Rowland. Computational methods applied to a skewed generalized normal family. Communications in Statistics: Simulation and Computation, 49(11):2930–2943, 2020. CODEN CSSCDB. ISSN 0361-0918.

Balagyozyan:2020:PTT

[BGM20] Aram Balagyozyan, Christos Giannikos, and Barry K. Ma. Power and thick tails: an ARCH process example with extreme value as test statistic. Communications in Statistics: Simulation and Computation, 49(2):556–564, 2020. CODEN CSSCDB. ISSN 0361-0918.

Beauducel:2020:FMI

[BH20] André Beauducel and Norbert Hilger. On the fit of models implied by unit-weighted scales. Communications in Statistics: Simulation and Computation, 49(11):3054–3064, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bicer:2020:SIG

[Biç20] Hayrinisa Demirci Biçer. Statistical inference for geometric process with the Two-Parameter Lindley Distribution. Communications in Statistics: Simulation and Computation, 49 (11):2979–3000, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bakar:2020:APS

[BJ20] K. Shuvo Bakar and Huidong Jin. Areal prediction of survey data using Bayesian spatial generalised linear models. Communications in Statistics: Simulation and Computation, 49 (11):2963–2978, 2020. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2020:DTN

[BJU20] S. Balamurali, P. Jeyadurga, and M. Usha. Designing of tightened-normal-tightened sampling scheme under Weibull and gamma distributions for mean life assurance. Communications in Statistics: Simulation and Computation, 49(7): 1704–1729, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bhowmik:2021:NOD

[BJVV21] Arpan Bhowmik, Seema Jaggi, Eldho Varghese, and Cini Varghese. A note on optimal directional neighbour designs with random block effect. Communications in Statistics: Simulation and Computation, 50(3):865–880, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Bnouachir:2021:ECB

[BM21] Najla Bnouachir and Abdallah Mkhadri. Efficient clusterbased portfolio optimization. Communications in Statistics: Simulation and Computation, 50(11):3241–3255, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Bahrami:2021:MMP

[BNK21] Hassan Bahrami, Seyed Taghi Akhavan Niaki, and Majid Khedmati. Monitoring multivariate profiles in multistage processes. Communications in Statistics: Simulation and Computation, 50(11):3436–3464, 2021. CODEN CSSCDB. ISSN 0361-0918.

Bhatt:2022:ATP

[BPN22] Milind B. Bhatt, Shantilal R. Patel, and Pratik B. Nandy. Asymptotic tests for parameters of two one-truncation parameter family of distributions. Communications in Statistics: Simulation and Computation, 51(4):1616–1636, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bhushan:2020:OIM

[BPP20] Shashi Bhushan, Abhay Pratap Pandey, and Arvind Pandey. On optimality of imputation methods for estimation of population mean using higher order moment of an auxiliary variable. Communications in Statistics: Simulation and Computation, 49(6):1560–1574, 2020. CODEN CSSCDB. ISSN 0361-0918.

Bickel:2021:CFD

[BR21] David R. Bickel and Abbas Rahal. Correcting false discovery rates for their bias toward false positives. Communications in Statistics: Simulation and Computation, 50(11):3699–3713, 2021. CODEN CSSCDB. ISSN 0361-0918.

Biswas:2020:RBT

[BRA20] Ankur Biswas, Anil Rai, and Tauqueer Ahmad. Rescaling bootstrap technique for variance estimation for ranked set samples in finite population. Communications in Statistics: Simulation and Computation, 49(10):2704–2718, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Brzezinska:2020:IRT

[Brz20] Justyna Brzezińska. Item response theory models in the measurement theory. Communications in Statistics: Simulation and Computation, 49(12):3299–3313, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Bulinskaya:2021:MAA

[BS21] Ekaterina V. Bulinskaya and Boris I. Shigida. Modeling and asymptotic analysis of insurance company performance. Communications in Statistics: Simulation and Computation, 50 (9):2743–2756, 2021. CODEN CSSCDB. ISSN 0361-0918.

Basalamah:2021:MIC

[BSNT21] Doaa Basalamah, Khamis K. Said, Wei Ning, and Yubin Tian. Modified information criterion for linear regression changepoint model with its applications. Communications in Statistics: Simulation and Computation, 50(1):180–197, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Bhattacharjee:2020:POM

[BT20] Dibyojyoti Bhattacharjee and Priyanka Talukdar. Predicting outcome of matches using pressure index: evidence from Twenty20 cricket. Communications in Statistics: Simulation and Computation, 49(11):3028–3040, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Bulgakova:2022:NSR

[BV22] T. E. Bulgakova and A. V. Voytishek. On numerical stability of randomized projection functional algorithms. Communications in Statistics: Simulation and Computation, 51(4): 1637–1646, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bhowmik:2022:GFD

[BVJV22] Arpan Bhowmik, Eldho Varghese, Seema Jaggi, and Cini Varghese. On the generation of factorial designs with minimum level changes. Communications in Statistics: Simulation and Computation, 51(6):3400–3409, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Bedouhene:2022:NMB

[BZ22a] Kahina Bedouhene and Nabil Zougab. Nonparametric multiplicative bias correction for von Mises kernel circular density estimator. Communications in Statistics: Simulation and Computation, 51(3):774–792, 2022. CODEN CSSCDB. ISSN 0361-0918.

Bulut:2022:ICR

[BZ22b] Hasan Bulut and Tolga Zaman. An improved class of robust ratio estimators by using the minimum covariance determinant estimation. Communications in Statistics: Simulation and Computation, 51(5):2457–2463, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chaturvedi:2021:SPE

[CCBJ21] Ajit Chaturvedi, Saibal Chattopadhyay, Sudeep R. Bapat, and Neeraj Joshi. Sequential point estimation procedures for the parameter of a family of distributions. Communications in Statistics: Simulation and Computation, 50(9):2678–2704, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chiew:2022:CDS

[CCBN22] Esther Chiew, Katherine Cauthen, Nathanael Brown, and Linda Nozick. Comparison of distribution selection methods. Communications in Statistics: Simulation and Computation, 51(4):1982–2005, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:SOC

[CCH20] Jen-Hao Chen, Yen-Chang Chang, and Wen-Liang Hung. A self-organizing clustering algorithm for functional data. Communications in Statistics: Simulation and Computation, 49 (5):1237–1263, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chowdhury:2020:COP

[CCM20] M. Chowdhury, M. Chen, and S. Mandal. A class of optimization problems on minimizing variance based criteria in respect of parameter estimators of a linear model. Communications in Statistics: Simulation and Computation, 49(10): 2719–2731, 2020. CODEN CSSCDB. ISSN 0361-0918.

Couto:2021:NDM

[CDB^{+}21] Ricardo Couto, Luiz H. Duczmal, Denise Burgarelli, Felipe Álvares, and Gladston J. P. Moreira. Nonparametric dependence modeling via cluster analysis: a financial contagion application. Communications in Statistics: Simulation and Computation, 50(2):537–556, 2021. CODEN CSSCDB. ISSN 0361-0918.

Channouf:2021:SSC

[CFM21] Nabil Channouf, Marc Fredette, and Brenda MacGibbon. Sample size calculations for hierarchical Poisson and zeroinflated Poisson regression models. Communications in Statistics: Simulation and Computation, 50(4):937–956, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chavent:2021:CCV

[CGS21] Marie Chavent, Robin Genuer, and Jérôme Saracco. Combining clustering of variables and feature selection using random forests. Communications in Statistics: Simulation and Computation, 50(2):426–445, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:SIM

[CHC20] Jinbao Chen, Yawen Hou, and Zheng Chen. Statistical inference methods for cumulative incidence function curves at a fixed point in time. Communications in Statistics: Simulation and Computation, 49(1):79–94, 2020. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2021:ELR

[Che21] Conghua Cheng. Empirical likelihood ratio under infinite second moment for two-sample problems. Communications in Statistics: Simulation and Computation, 50(10):3069–3076, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:PCD

[CHL22] Wenting Chen, Xin-Jiang He, and Sha Lin. Pricing credit default swaps with Parisian and Parasian default mechanics. Communications in Statistics: Simulation and Computation, 51(2):421–431, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cekim:2020:TEP

[CK20] Hatice Oncel Cekim and Cem Kadilar. In-type estimators for the population variance in stratified random sampling. Communications in Statistics: Simulation and Computation, 49(7):1665–1677, 2020. CODEN CSSCDB. ISSN 0361-0918.

Curtin:2021:RSI

[CKCG21] Paul Curtin, Joshua Kellogg, Nadja Cech, and Chris Gennings. A random subset implementation of weighted quantile sum (WQS_{RS}) regression for analysis of high-dimensional mixtures. Communications in Statistics: Simulation and Computation, 50(4):1119–1134, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chaturvedi:2021:NSR

[CKK21] Ajit Chaturvedi, Taruna Kumari, and Narendra Kumar. Numerical study of robust Bayesian analysis of generalized inverted family of distributions based on progressive type II right censoring. Communications in Statistics: Simulation and Computation, 50(11):3207–3240, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Chew:2022:ERR

[CKL22] XinYing Chew, Khai Wah Khaw, and Ming Ha Lee. The efficiency of run rules schemes for the multivariate coefficient of variation in short runs process. Communications in Statistics: Simulation and Computation, 51(6):2942–2962, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Chong:2020:ODM

[CKT^{+}20] Zhi Lin Chong, Michael Boon Chong Khoo, Wei Lin Teoh, Wai Chung Yeong, and Sok Li Lim. Optimal design of the modified group runs (MGR) X̄ chart when process parameters are estimated. Communications in Statistics: Simulation and Computation, 49(1):244–260, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chin:2021:NHF

[CL21] Wen Cheong Chin and Min Cherng Lee. Nonlinear highfrequency stock market time series: Modeling and combine forecast evaluations. Communications in Statistics: Simulation and Computation, 50(7):2126–2144, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Castellares:2022:CEG

[CL22a] Fredy Castellares and Artur J. Lemonte. Comments on “An extended Gompertz–Makeham distribution with application to lifetime data”. Communications in Statistics: Simulation and Computation, 51(5):2769–2777, 2022. CODEN CSSCDB. ISSN 0361-0918. See [Abd18].

Chiu:2022:IVA

[CL22b] Sung Nok Chiu and Kwong Ip Liu. Improving p-value approximation and level accuracy of Monte Carlo tests by quasi-Monte Carlo methods. Communications in Statistics: Simulation and Computation, 51(3):1272–1288, 2022. CODEN CSSCDB. ISSN 0361-0918.

Collins:2022:SRS

[CL22c] David Collins and Yan Lu. A stratified reservoir sampling algorithm in streams and large datasets. Communications in Statistics: Simulation and Computation, 51(4):1767–1782, 2022. CODEN CSSCDB. ISSN 0361-0918.

Choi:2021:PSM

[CLJK21] Junsouk Choi, JungJun Lee, Jae-Hwan Jhong, and Ja-Yong Koo. Penalized I-spline monotone regression estimation. Communications in Statistics: Simulation and Computation, 50(11):3714–3732, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:CML

[CLL20] Nan Chen, Meijuan Li, and Hongyun Liu. Comparison of maximum likelihood approach, Diggle–Kenward selection model, pattern mixture model with MAR and MNAR dropout data. Communications in Statistics: Simulation and Computation, 49(7):1746–1767, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chen:2022:NGW

[CLS22] Jen-Hsiang Chen, Shin-Li Lu, and Shey-Huei Sheu. A nonparametric generally weighted moving average sign chart based on repetitive sampling. Communications in Statistics: Simulation and Computation, 51(3):1137–1156, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2021:OCB

[CLY21] Lee-Shen Chen, TaChen Liang, and Ming-Chung Yang. Optimal curtailed Bayesian sampling plans for exponential distributions with Type-I hybrid censored samples. Communications in Statistics: Simulation and Computation, 50(3): 764–777, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cockeran:2021:GFT

[CMA21] Marike Cockeran, Simos George Meintanis, and James S. Allison. Goodness-of-fit tests in the Cox proportional hazards model. Communications in Statistics: Simulation and Computation, 50(12):4132–4143, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chiodini:2021:RSA

[CMMV21] Paola Maddalena Chiodini, Giancarlo Manzi, Bianca Maria Martelli, and Flavio Verrecchia. Revisiting sample allocation methods: a simulation-based comparison. Communications in Statistics: Simulation and Computation, 50(8):2197–2212, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cribari-Neto:2021:RBP

[CNL21] Francisco Cribari-Neto and Fábio P. Lima. Resampling-based prediction intervals in beta regressions under correct and incorrect model specification. Communications in Statistics: Simulation and Computation, 50(5):1398–1416, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2022:FTA

[CP22] Chao Chen and Haiyan Pan. A fiducial test for assessing the non-inferiority of odds ratio in matched-pairs design. Communications in Statistics: Simulation and Computation, 51 (6):2978–2991, 2022. CODEN CSSCDB. ISSN 0361-0918.

Carrasco:2021:SRM

[CR21] Jalmar M. F. Carrasco and Nancy Reid. Simplex regression models with measurement error. Communications in Statistics: Simulation and Computation, 50(11):3420–3435, 2021. CODEN CSSCDB. ISSN 0361-0918.

Choi:2020:BVS

[CS20] Ji-Eun Choi and Dong Wan Shin. Bootstrapping volatility spillover index. Communications in Statistics: Simulation and Computation, 49(1):66–78, 2020. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2022:IMP

[CS22] Maolin Cheng and Guojun Shi. Improved methods for parameter estimation of gray model GM(1,1) based on new background value optimization and model application. Communications in Statistics: Simulation and Computation, 51(2): 647–669, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2020:IGP

[CSDT20] Xudan Chen, Xinli Sun, Xiong Ding, and Jue Tang. The inverse Gaussian process with a skew-normal distribution as a degradation model. Communications in Statistics: Simulation and Computation, 49(11):2827–2843, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chen:2021:PCP

[CSE^{+}21] Yanjun Chen, Damla Sentürk, Jason P. Estes, Luis F. Campos, Connie M. Rhee, Lorien S. Dalrymple, Kamyar Kalantar-Zadeh, and Danh V. Nguyen. Performance characteristics of profiling methods and the impact of inadequate case-mix adjustment. Communications in Statistics: Simulation and Computation, 50(6):1854–1871, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cuadras:2022:CMF

[CSG22] Carles M. Cuadras and Sonia Salvo-Garrido. Comparing multiple factor analysis and related metric scaling. Communications in Statistics: Simulation and Computation, 51(5):2343– 2362, 2022. CODEN CSSCDB. ISSN 0361-0918.

Cai:2020:BAD

[CSL20] Jing Cai, Yimin Shi, and Bin Liu. Bayesian analysis for dependent competing risks model with masked causes of failure in step-stress accelerated life test under progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 49(9):2302–2320, 2020. CODEN CSSCDB. ISSN 0361-0918.

Chang:2021:CID

[CST21] Yu-Mei Chang, Pao-Sheng Shen, and Yu-Hsin Tang. Confidence interval for the difference between two median survival times with semiparametric transformation models. Communications in Statistics: Simulation and Computation, 50(3): 633–649, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2020:IPE

[CSX20] Maolin Cheng, Guojun Shi, and Mingyin Xiang. On the improvement of the parameter estimation of the grey model GM(1,1) and model application. Communications in Statistics: Simulation and Computation, 49(5):1367–1384, 2020. CODEN CSSCDB. ISSN 0361-0918.

Cavus:2021:PPA

[CYS21] Mustafa Cavus, Berna Yazici, and Ahmet Sezer. Penalized power approach to compare the power of the tests when Type I error probabilities are different. Communications in Statistics: Simulation and Computation, 50(7):1912–1926, 2021. CODEN CSSCDB. ISSN 0361-0918.

Chwila:2022:PEB

[CZ22] Adam Chwila and Tomasz Zadło. On properties of empirical best predictors. Communications in Statistics: Simulation and Computation, 51(1):220–253, 2022. CODEN CSSCDB. ISSN 0361-0918.

Chen:2021:CSR

[CZP21] Chen Chen, Jie Zhou, and Jianxin Pan. Correlation structure regularization via entropy loss function for high-dimension and low-sample-size data. Communications in Statistics: Simulation and Computation, 50(4):993–1008, 2021. CODEN CSSCDB. ISSN 0361-0918.

Cui:2021:MDE

[CZS21] Lihong Cui, Qianhui Zhang, and Jianjun Sun. Multiwavelet density estimation for biased data. Communications in Statistics: Simulation and Computation, 50(1):234–253, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Dogru:2021:RMR

[DA21] Fatma Zehra Dogru and Olcay Arslan. Robust mixture regression modeling based on the generalized M (GM)-estimation method. Communications in Statistics: Simulation and Computation, 50(9):2643–2665, 2021. CODEN CSSCDB. ISSN 0361-0918.

Deng:2022:BPA

[DA22] M. Deng and M. S. Aminzadeh. Bayesian predictive analysis for Weibull–Pareto composite model with an application to insurance data. Communications in Statistics: Simulation and Computation, 51(5):2683–2709, 2022. CODEN CSSCDB. ISSN 0361-0918.

Inacio:2020:CTP

[dAIIS20] Marco Henrique de Almeida Inácio, Rafael Izbicki, and Luis Ernesto Salasar. Comparing two populations using Bayesian Fourier series density estimation. Communications in Statistics: Simulation and Computation, 49(1):261–282, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lima-Filho:2021:KCC

[dALFB21] Luiz Medeiros de Araujo Lima-Filho and Fábio Mariano Bayer. Kumaraswamy control chart for monitoring double bounded environmental data. Communications in Statistics: Simulation and Computation, 50(9):2513–2528, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Neto:2021:WSP

[dALNPdOF21] Eufrásio de Andrade Lima Neto, Aluísio Pinheiro, and Adenice Gomes de Oliveira Ferreira. On wavelet to select the parametric form of a regression model. Communications in Statistics: Simulation and Computation, 50(9):2619–2642, 2021. CODEN CSSCDB. ISSN 0361-0918.

Davidson:2020:MPF

[DC20] Heather Davidson and Robert A. Cribbie. A more powerful familywise error control procedure for evaluating mean equivalence. Communications in Statistics: Simulation and Computation, 49(11):2914–2929, 2020. CODEN CSSCDB. ISSN 0361-0918.

Djeddour-Djaballah:2021:QML

[DDK21] K. Djeddour-Djaballah and L. Kerar. Quasi-maximum likelihood estimation of GARCH with Student distributed noise. Communications in Statistics: Simulation and Computation, 50(5):1249–1271, 2021. CODEN CSSCDB. ISSN 0361-0918.

DeGaetano:2020:FGM

[De 20] Davide De Gaetano. Forecasting with GARCH models under structural breaks: an approach based on combinations across estimation windows. Communications in Statistics: Simulation and Computation, 49(10):2559–2582, 2020. CODEN CSSCDB. ISSN 0361-0918.

DeOliveira:2022:NSN

[DE22] Victor De Oliveira and Mark D. Ecker. A non-stationary non-Gaussian hedonic spatial model for house selling prices. Communications in Statistics: Simulation and Computation, 51(6):2888–2905, 2022. CODEN CSSCDB. ISSN 0361-0918.

Duran:2022:MRA

[DF22] Ahmet Duran and Mohammed Farrukh. Multiple regression analysis for dynamics of patient volumes. Communications in Statistics: Simulation and Computation, 51(6):2906–2923, 2022. CODEN CSSCDB. ISSN 0361-0918.

Dag:2022:DCE

[DKKA22] Osman Dag, Merve Kasikci, Erdem Karabulut, and Reha Alpar. Diverse classifiers ensemble based on GMDH-type neural network algorithm for binary classification. Communications in Statistics: Simulation and Computation, 51(5):2440–2456, 2022. CODEN CSSCDB. ISSN 0361-0918.

Draxler:2022:GTF

[DKL22] Clemens Draxler, Andreas Kurz, and Artur J. Lemonte. The gradient test and its finite sample size properties in a conditional maximum likelihood and psychometric modeling context. Communications in Statistics: Simulation and Computation, 51(6):3185–3203, 2022. CODEN CSSCDB. ISSN 0361-0918.

Duan:2020:FSF

[DL20] Qihong Duan and Junrong Liu. Filtering and smoothing formulas of AR(p)-modulated Poisson processes. Communications in Statistics: Simulation and Computation, 49(6):1575– 1591, 2020. CODEN CSSCDB. ISSN 0361-0918.

Dey:2021:NRM

[DLRZ21] Dipak K. Dey, Junfeng Liu, Nalini Ravishanker, and Edwards Qiang Zhang. A note on response mean confidence band for linear regression models. Communications in Statistics: Simulation and Computation, 50(3):778–785, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Du:2022:ETD

[DNWG22] Yujing Du, Hon Keung Tony Ng, Jun Wang, and Wei Gao. On estimation of two-dimensional dynamic panel model with confounders. Communications in Statistics: Simulation and Computation, 51(2):505–517, 2022. CODEN CSSCDB. ISSN 0361-0918.

Dogan:2022:SSC

[Dog22] Ilkay Dogan. A simulation study comparing model fit measures of structural equation modeling with multivariate contaminated normal distribution. Communications in Statistics: Simulation and Computation, 51(5):2526–2536, 2022. CODEN CSSCDB. ISSN 0361-0918.

deOliveira:2021:GBD

[dOMA21] Ricardo Puziol de Oliveira, Josmar Mazucheli, and Jorge Alberto Achcar. A generalization of Basu-Dhar’s bivariate geometric distribution to the trivariate case. Communications in Statistics: Simulation and Computation, 50(12):4388–4412, 2021. CODEN CSSCDB. ISSN 0361-0918.

deOliveira:2021:SCA

[dOMD^{+}21] Dênis Ricardo Xavier de Oliveira, Gladston Moreira, Anderson Ribeiro Duarte, André Cançado, and Eduardo Luz. Spatial cluster analysis using particle swarm optimization and dispersion function. Communications in Statistics: Simulation and Computation, 50(8):2368–2385, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Valadares:2021:NCR

[dOVQP21] Danilo Gilberto de Oliveira Valadares, Roberto Costa Quinino, and Magda Carvalho Pires. The need to conduct repeated classifications in a logistic regression model with misclassification in the dependent variable. Communications in Statistics: Simulation and Computation, 50(5):1459–1472, 2021. CODEN CSSCDB. ISSN 0361-0918.

Dedduwakumara:2020:CIQ

[DP20] Dilanka S. Dedduwakumara and Luke A. Prendergast. Confidence intervals for quantiles from histograms and other grouped data. Communications in Statistics: Simulation and Computation, 49(6):1546–1559, 2020. CODEN CSSCDB. ISSN 0361-0918.

Divecha:2022:TEL

[DR22] Jyoti Divecha and D. P. Raykundaliya. Three economical life test acceptance sampling plans. Communications in Statistics: Simulation and Computation, 51(6):3305–3323, 2022. CODEN CSSCDB. ISSN 0361-0918.

Das:2020:AUW

[DRB20] Sourav Das, Tata Subba Rao, and Georgi N. Boshnakov. An asymptotically unbiased weighted least squares estimation criterion for parametric variograms of second order stationary geostatistical processes. Communications in Statistics: Simulation and Computation, 49(7):1839–1854, 2020. CODEN CSSCDB. ISSN 0361-0918.

da-Silva:2021:DMA

[dSdSTC21] Cibele Q. da Silva, Paulo H. D. da Silva, Osiris Turnes, and Leandro T. Correia. Dynamic model averaging adapted to dynamic regression models for time series of counts. Communications in Statistics: Simulation and Computation, 50(2): 501–524, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ding:2020:MMB

[DSL20] Hao Ding, Ziwei Su, and Xiaoqian Liu. A modified multinomial baseline logit model with logit functions having different covariates. Communications in Statistics: Simulation and Computation, 49(11):2861–2875, 2020. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:AVR

[dXGLY20] Guo dong Xing, Xiaoli Gan, Xiaohu Li, and Shanchao Yang. On the asymptotics of value-at-risk for portfolio loss under bivariate Eyraud–Farlie–Gumbel–Morgenstern copula and heavy tails. Communications in Statistics: Simulation and Computation, 49(9):2462–2471, 2020. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:ATC

[dXLY20a] Guo dong Xing, Xiaohu Li, and Shanchao Yang. On the asymptotics of tail conditional expectation for portfolio loss under bivariate Eyraud–Farlie–Gumbel–Morgenstern copula and heavy tails. Communications in Statistics: Simulation and Computation, 49(8):2049–2058, 2020. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:SOA

[dXLY20b] Guo dong Xing, Xiaohu Li, and Shanchao Yang. Second-order asymptotics of tail distortion risk measure for portfolio loss in the multivariate regularly varying model. Communications in Statistics: Simulation and Computation, 49(2):491–503, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ding:2022:TSC

[DY22] Jianhua Ding and Ping Yu. Test shape constraints in semiparametric model with Bernstein polynomials. Communications in Statistics: Simulation and Computation, 51(5):2667– 2682, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ding:2022:BEB

[DYC22] Liwang Ding, Shengwei Yao, and Ping Chen. On Berry– Esseen bound of wavelet estimators in nonparametric regression model under asymptotically negatively associated assumptions. Communications in Statistics: Simulation and Computation, 51(3):924–940, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ebadi:2020:CCM

[EAJ20] Mohsen Ebadi and Amir Ahmadi-Javid. Control charts for monitoring multi-stage service processes with optimal queue performance. Communications in Statistics: Simulation and Computation, 49(9):2472–2484, 2020. CODEN CSSCDB. ISSN 0361-0918.

Elsawah:2021:SIB

[EFD21] A. M. Elsawah, Kai-Tai Fang, and Yu Hui Deng. Some interesting behaviors of good lattice point sets. Communications in Statistics: Simulation and Computation, 50(11):3650–3668, 2021. CODEN CSSCDB. ISSN 0361-0918.

Epprecht:2021:VSF

[EGVdR21] Camila Epprecht, Dominique Guégan, Álvaro Veiga, and Joel Correa da Rosa. Variable selection and forecasting via automated methods for linear models: LASSO/adaLASSO and Autometrics. Communications in Statistics: Simulation and Computation, 50(1):103–122, 2021. CODEN CSSCDB. ISSN 0361-0918.

El-Horbaty:2022:TAR

[EH22] Yahia S. El-Horbaty. Testing the absence of random effects in the nested-error regression model using orthogonal transformations. Communications in Statistics: Simulation and Computation, 51(5):2736–2746, 2022. CODEN CSSCDB. ISSN 0361-0918.

Enomoto:2020:MNT

[EHHS20] Rie Enomoto, Zofia Hanusz, Ayako Hara, and Takashi Seo. Multivariate normality test using normalizing transformation for Mardia’s multivariate kurtosis. Communications in Statistics: Simulation and Computation, 49(3):684–698, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Ermakov:2021:DMQ

[EL21] Sergej M. Ermakov and Svetlana N. Leora. Decrease of the mean of the quasi-random integration error. Communications in Statistics: Simulation and Computation, 50(11): 3581–3589, 2021. CODEN CSSCDB. ISSN 0361-0918.

Erem:2020:BTS

[Ere20] Aysegul Erem. Bivariate two sample test based on exceedance statistics. Communications in Statistics: Simulation and Computation, 49(9):2389–2401, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ermakov:2022:GGO

[ES22] Sergej M. Ermakov and Dmitriy N. Semenchikov. Genetic global optimization algorithms. Communications in Statistics: Simulation and Computation, 51(4):1503–1512, 2022. CODEN CSSCDB. ISSN 0361-0918.

ElShekhAhmed:2020:AWN

[ESES20] Hazem I. El Shekh Ahmed, Raid B. Salha, and Hossam O. El-Sayed. Adaptive weighted Nadaraya–Watson estimation of the conditional quantiles by varying bandwidth. Communications in Statistics: Simulation and Computation, 49(5): 1105–1117, 2020. CODEN CSSCDB. ISSN 0361-0918.

Eskandari:2021:AES

[Esk21] Farzad Eskandari. Asymptotic efficient semiparametric empirical Bayes estimation of multinomial responses. Communications in Statistics: Simulation and Computation, 50(1): 123–141, 2021. CODEN CSSCDB. ISSN 0361-0918.

Eappen:2022:RRT

[ESS22] Christin Variathu Eappen, Stephen A. Sedory, and Sarjinder Singh. Ratio and regression type estimators of a new measure of coefficient of dispersion. Communications in Statistics: Simulation and Computation, 51(4):1899–1920, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Fischer:2021:PRR

[FBNRG21] Daniel Fischer, Alain Berro, Klaus Nordhausen, and Anne Ruiz-Gazen. REPPlab: an R package for detecting clusters and outliers using exploratory projection pursuit. Communications in Statistics: Simulation and Computation, 50(11): 3397–3419, 2021. CODEN CSSCDB. ISSN 0361-0918.

Faisal:2022:BIW

[FH22] Shahla Faisal and Christian Heumann. Bootstrap inference for weighted nearest neighbors imputation. Communications in Statistics: Simulation and Computation, 51(6):2842–2857, 2022. CODEN CSSCDB. ISSN 0361-0918.

Firinguetti-Limone:2020:BES

[FLPB20] Luis Firinguetti-Limone and Manuel Pereira-Barahona. Bayesian estimation of the shrinkage parameter in ridge regression. Communications in Statistics: Simulation and Computation, 49(12):3314–3327, 2020. CODEN CSSCDB. ISSN 0361-0918.

Farokhnia:2020:PCA

[FN20] Moez Farokhnia and S. T. A. Niaki. Principal component analysis-based control charts using support vector machines for multivariate non-normal distributions. Communications in Statistics: Simulation and Computation, 49(7):1815–1838, 2020. CODEN CSSCDB. ISSN 0361-0918.

Flores:2022:PSH

[FO22] Pablo Flores and Jordi Ocaña. Pretesting strategies for homoscedasticity when comparing means. their robustness facing non-normality. Communications in Statistics: Simulation and Computation, 51(1):280–292, 2022. CODEN CSSCDB. ISSN 0361-0918.

Filipiak:2020:PAL

[FP20] Katarzyna Filipiak and Simo Puntanen. Preface — advances on linear models and inference: computational aspects. Communications in Statistics: Simulation and Computation, 49 (3):565–566, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ferrer-Perez:2021:SUR

[FPAA21] Hugo Ferrer-Pérez, María-Isabel Ayuda, and Antonio Aznar. The sensitivity of unit root tests to the initial condition and to the lag length selection: a Monte Carlo simulation study. Communications in Statistics: Simulation and Computation, 50(4):1062–1072, 2021. CODEN CSSCDB. ISSN 0361-0918.

Fernando:2022:DGF

[FS22] Gayara Fernando and Roshini Sooriyarachchi. The development of a goodness-of-fit test for high level binary multilevel models. Communications in Statistics: Simulation and Computation, 51(5):2710–2730, 2022. CODEN CSSCDB. ISSN 0361-0918.

Su:2021:DFD

[fSLC21] Zhi fang Su, Ken-Yu Lin, and Mei-Yuan Chen. Dual-frailty default intensity model: estimations and an application. Communications in Statistics: Simulation and Computation, 50(5):1315–1332, 2021. CODEN CSSCDB. ISSN 0361-0918.

Fang:2021:SDE

[FTC21] Yongfei Fang, Yubin Tian, and Xia Cai. Sequential designs for estimating the operating window. Communications in Statistics: Simulation and Computation, 50(8):2386–2398, 2021. CODEN CSSCDB. ISSN 0361-0918.

Fan:2021:ACE

[FY21] Qingliang Fan and Wanlu Yu. Adaptive k-class estimation in high-dimensional linear models. Communications in Statistics: Simulation and Computation, 50(12):3885–3913, 2021. CODEN CSSCDB. ISSN 0361-0918.

Farzammehr:2021:SNG

[FZM21] Mohadeseh Alsadat Farzammehr, Mohammad Reza Zadkarami, and Geoffrey J. McLachlan. Skew-normal generalized spatial panel data model. Communications in Statistics: Simulation and Computation, 50(11):3286–3314, 2021. CODEN CSSCDB. ISSN 0361-0918.

Feng:2020:SRC

[FZZ20] Yan Feng, Xiaobing Zhao, and Xian Zhou. Semiparametric random censorship models for survival data with long-term survivors. Communications in Statistics: Simulation and Computation, 49(11):2876–2896, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ghitany:2022:GLD

[GABG22] Mohamed E. Ghitany, Suja M. Aboukhamseen, Abrar A. Baqer, and Ramesh C. Gupta. Gompertz–Lindley distribution and associated inference. Communications in Statistics: Simulation and Computation, 51(5):2599–2618, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Gaigall:2021:NAM

[Gai21] Daniel Gaigall. On a new approach to the multi-sample goodness-of-fit problem. Communications in Statistics: Simulation and Computation, 50(10):2971–2989, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ghaly:2020:ACA

[GAS20] A. A. Abdel Ghaly, H. M. Aly, and R. N. Salah. Applying the copula approach on step stress accelerated life test under type II censoring. Communications in Statistics: Simulation and Computation, 49(1):159–177, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gangsei:2022:LRB

[GAS22] Lars Erik Gangsei, Trygve Almøy, and Solve Sæbø. Linear regression with bivariate response variable containing missing data. strategies to increase prediction precision. Communications in Statistics: Simulation and Computation, 51(2): 527–538, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Barrios:2020:DAC

[GBHA20] José María González-Barrios and Ricardo Hoyos-Argüelles. Distributions associated to the counting techniques of the dsample copula of order m and weak convergence of the sample process. Communications in Statistics: Simulation and Computation, 49(10):2505–2532, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Barrios:2021:ECA

[GBHA21] José M. González-Barrios and Ricardo Hoyos-Argüelles. Estimating checkerboard approximations with sample d-copulas. Communications in Statistics: Simulation and Computation, 50(12):3992–4027, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2020:CHV

[GCF^{+}20] M. Ivette Gomes, Frederico Caeiro, Fernanda Figueiredo, Lgia Henriques-Rodrigues, and Dinis Pestana. Corrected-Hill versus partially reduced-bias value-at-risk estimation. Communications in Statistics: Simulation and Computation, 49 (4):867–885, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gera:2020:ATP

[Ger20] Amos E. Gera. Acceptance testing for a planar array. Communications in Statistics: Simulation and Computation, 49 (5):1198–1209, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gera:2021:RP

[Ger21] Amos E. Gera. From runs to patterns. Communications in Statistics: Simulation and Computation, 50(12):4300–4314, 2021. CODEN CSSCDB. ISSN 0361-0918.

Guris:2022:GDN

[GG22] Selahattin Güris and Burak Güris. GLS detrending in nonlinear unit root test. Communications in Statistics: Simulation and Computation, 51(3):1096–1102, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghasemzadeh:2020:BQR

[GGB20] Siamak Ghasemzadeh, Mojtaba Ganjali, and Taban Baghfalaki. Bayesian quantile regression for joint modeling of longitudinal mixed ordinal and continuous data. Communications in Statistics: Simulation and Computation, 49(2):375– 395, 2020. CODEN CSSCDB. ISSN 0361-0918.

Giancristofaro:2022:RAC

[GGMM22] R. Arboretti Giancristofaro, M. Gastaldi, L. Martinello, and C. Meneguzzer. Regression analysis with compositional data using orthogonal log-ratio coordinates. Communications in Statistics: Simulation and Computation, 51(4):1932–1945, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghapani:2022:SRL

[Gha22] Fatemeh Ghapani. Stochastic restricted Liu estimator in linear mixed measurement error models. Communications in Statistics: Simulation and Computation, 51(3):1220–1233, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2020:NUE

[Gho20] Sucharita Ghosh. A note on using the empirical moment generating function to estimate the variance of nonparametric trend estimates from independent time series replicates. Communications in Statistics: Simulation and Computation, 49 (9):2287–2301, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gao:2020:ESS

[GHW20] Zhongqin Gao, Jingmin He, and Bingbing Wang. Exact solutions of some exit times for the diffusion risk model with liquid reserves, credit and debit interest. Communications in Statistics: Simulation and Computation, 49(10):2693–2703, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gad:2020:ALR

[GI20] Ahmed M. Gad and Wafaa I. M. Ibrahim. An adaptive linear regression approach for modeling heavy-tailed longitudinal data. Communications in Statistics: Simulation and Computation, 49(5):1181–1197, 2020. CODEN CSSCDB. ISSN 0361-0918.

Goudarzi:2021:NBO

[GJK21] Manizheh Goudarzi, Habib Jafari, and Soleiman Khazaei. Nonparametric Bayesian optimal designs for exponential regression model. Communications in Statistics: Simulation and Computation, 50(6):1809–1819, 2021. CODEN CSSCDB. ISSN 0361-0918.

Goyal:2020:NMS

[GK20] Manish Goyal and Narinder Kumar. Nonparametric multiple sample scale testing using U-statistics. Communications in Statistics: Simulation and Computation, 49(11):3019–3027, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2022:SBS

[GK22a] M. P. Gadre and V. C. Kakade. Sample based side sensitive group runs control chart to detect shifts in the process mean. Communications in Statistics: Simulation and Computation, 51(1):190–200, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2022:SSS

[GK22b] Mukund P. Gadre and Vikas C. Kakade. Some side sensitive group runs based control charts to detect shifts in the process median. Communications in Statistics: Simulation and Computation, 51(4):1474–1487, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghasami:2020:APG

[GKM20] Safdar Ghasami, Zahra Khodadadi, and Mohsen Maleki. Autoregressive processes with generalized hyperbolic innovations. Communications in Statistics: Simulation and Computation, 49(12):3080–3092, 2020. CODEN CSSCDB. ISSN 0361-0918.

Guo:2022:DLS

[GKZ^{+}22] Zhaoyu Guo, Irwin Khuu, Kevin Zhu, Jeffrey S. Rosenthal, and Frederic P. Schoenberg. Distinguishing luck from skill through statistical simulation: a case study. Communications in Statistics: Simulation and Computation, 51(5):2537–2559, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gan:2020:CCE

[GLT20] Shengjun Gan, Changli Lu, and Yongge Tian. Computation and comparison of estimators under different linear randomeffects models. Communications in Statistics: Simulation and Computation, 49(5):1210–1222, 2020. CODEN CSSCDB. ISSN 0361-0918.

Gammoudi:2021:GQM

[GNE21] Imed Gammoudi, Asma Nani, and Mohamed El Ghourabi. Generalized quasi maximum likelihood estimation for generalized autoregressive score models: simulations and real applications. Communications in Statistics: Simulation and Computation, 50(11):3338–3363, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gorecki:2020:GPC

[GP20] Tomasz Górecki and Paweł Piasecki. Generalization of the Procrustes coefficient to functional data. Communications in Statistics: Simulation and Computation, 49(3):808–816, 2020. CODEN CSSCDB. ISSN 0361-0918.

Georghiou:2021:MNB

[GPP21] Costas Georghiou, Andreas N. Philippou, and Zaharias M. Psillakis. On the modes of the negative binomial distribution of order k, type I. Communications in Statistics: Simulation and Computation, 50(4):1217–1230, 2021. CODEN CSSCDB. ISSN 0361-0918.

Grzesiek:2022:CCB

[GTW22] Aleksandra Grzesiek, Marek Teuerle, and Agnieszka Wyłomańska. Cross-codifference for bidimensional VAR(1) time series with infinite variance. Communications in Statistics: Simulation and Computation, 51(3):1355–1380, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gu:2020:SCC

[GW20] Xiaolan Gu and Qiusheng Wang. Sparse canonical correlation analysis algorithm with alternating direction method of multipliers. Communications in Statistics: Simulation and Computation, 49(9):2372–2388, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ghitany:2022:NPA

[GW22] M. E. Ghitany and Shaochen Wang. A note on parameter asymptotics for weighted Lindley distribution. Communications in Statistics: Simulation and Computation, 51(4):1513– 1524, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:MSM

[GWZC22] Yujie Gai, Yusheng Wei, Jun Zhang, and Aixian Chen. Measuring the symmetry of model errors for varying coefficient regression models based on correlation coefficient. Communications in Statistics: Simulation and Computation, 51(5): 2235–2251, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ghalani:2021:ICS

[GZ21a] Mohammad Reza Ghalani and Mohammad Reza Zadkarami. Investigation of covariance structures in modelling longitudinal ordinal responses with skew normal random effect. Communications in Statistics: Simulation and Computation, 50 (1):254–269, 2021. CODEN CSSCDB. ISSN 0361-0918.

Guo:2021:SLT

[GZ21b] Xiao Guo and Hai Zhang. Structured learning of time-varying networks with application to PM_{2.5} data. Communications in Statistics: Simulation and Computation, 50(5):1364–1382, 2021. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:DSM

[GZ22a] Yujie Gai and Jun Zhang. Detection of the symmetry of model errors for partial linear single-index models. Communications in Statistics: Simulation and Computation, 51(6):3410–3427, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:DSA

[GZ22b] Yujie Gai and Jun Zhang. Detection the symmetry or asymmetry of model errors in partial linear models. Communications in Statistics: Simulation and Computation, 51(5):2217– 2234, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:SIR

[GZ22c] Yujie Gai and Jun Zhang. Single-index relative error regression models. Communications in Statistics: Simulation and Computation, 51(5):2252–2267, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gao:2022:IMN

[GZY22] Siyu Gao, Qianqian Zhang, and Jing Yu. The impact of misspecification of nuisance parameters on test for homogeneity in zero-inflated Poisson model: a simulation study. Communications in Statistics: Simulation and Computation, 51(1): 84–98, 2022. CODEN CSSCDB. ISSN 0361-0918.

Gai:2022:TSM

[GZZ22] Yujie Gai, Xuehu Zhu, and Jun Zhang. Testing symmetry of model errors for nonparametric regression models by using correlation coefficient^{1}. Communications in Statistics: Simulation and Computation, 51(4):1436–1453, 2022. CODEN CSSCDB. ISSN 0361-0918.

Han:2020:BER

[Han20] Ming Han. E-Bayesian estimations of the reliability and its eposterior risk under different loss functions. Communications in Statistics: Simulation and Computation, 49(6):1527–1545, 2020. CODEN CSSCDB. ISSN 0361-0918.

Han:2021:BEP

[Han21] Ming Han. E-Bayesian estimations of parameter and its evaluation standard: E-MSE (expected mean square error) under different loss functions. Communications in Statistics: Simulation and Computation, 50(7):1971–1988, 2021. CODEN CSSCDB. ISSN 0361-0918.

Harwell:2021:ESV

[Har21] Michael Harwell. An effect size for variance heterogeneity in meta-analysis. Communications in Statistics: Simulation and Computation, 50(7):1955–1970, 2021. CODEN CSSCDB. ISSN 0361-0918.

Han:2022:RAU

[HAT^{+}22] Seungbong Han, Adin-Cristian Andrei, Kam-Wah Tsui, Sung-Cheol Yun, and Jong Ho Yoon. ROC analysis using covariate balancing propensity scores with an application to biochemical predictors for thyroid cancer. Communications in Statistics: Simulation and Computation, 51(2):374–390, 2022. CODEN CSSCDB. ISSN 0361-0918.

Haq:2022:DCC

[HB22] Abdul Haq and Lubna Bibi. The dual CUSUM charts with auxiliary information for process mean. Communications in Statistics: Simulation and Computation, 51(1):164–189, 2022. CODEN CSSCDB. ISSN 0361-0918.

Holling:2020:ENV

[HBM^{+}20] Heinz Holling, Walailuck Böhning, Ehsan Masoudi, Dankmar Böhning, and Patarawan Sangnawakij. Evaluation of a new version of I^{2} with emphasis on diagnostic problems. Communications in Statistics: Simulation and Computation, 49(4): 942–972, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hu:2021:GCP

[HCTZ21] XueLong Hu, Philippe Castagliola, Anan Tang, and Jianlan Zhong. Guaranteed conditional performance of the median run length based EWMA X̄ chart with unknown process parameters. Communications in Statistics: Simulation and Computation, 50(12):4280–4299, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hu:2021:FIC

[HCZ21] Guozhi Hu, Weihu Cheng, and Jie Zeng. Focused information criterion and model averaging for varying-coefficient partially linear models with longitudinal data. Communications in Statistics: Simulation and Computation, 50(8):2399–2417, 2021. CODEN CSSCDB. ISSN 0361-0918.

Han:2020:MLE

[HD20] Zifei Han and Victor De Oliveira. Maximum likelihood estimation of Gaussian copula models for geostatistical count data. Communications in Statistics: Simulation and Computation, 49(8):1957–1981, 2020. CODEN CSSCDB. ISSN 0361-0918.

He:2022:MAP

[HD22] Baihua He and Fangli Dong. Model averaging with privacypreserving. Communications in Statistics: Simulation and Computation, 51(4):1401–1414, 2022. CODEN CSSCDB. ISSN 0361-0918.

Huang:2021:MDP

[HE21] Xin-Wei Huang and Takeshi Emura. Model diagnostic procedures for copula-based Markov chain models for statistical process control. Communications in Statistics: Simulation and Computation, 50(8):2345–2367, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hepsag:2021:TCN

[Hep21a] Aycan Hepsag. Testing for cointegration in nonlinear asymmetric smooth transition error correction models. Communications in Statistics: Simulation and Computation, 50(2): 400–412, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hepsag:2021:URT

[Hep21b] Aycan Hepsag. A unit root test based on smooth transitions and nonlinear adjustment. Communications in Statistics: Simulation and Computation, 50(3):625–632, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Huang:2021:REI

[HHY21] Mian Huang, Kang He, and Weixin Yao. Regression estimation via information-weighted composite models with different dimensions. Communications in Statistics: Simulation and Computation, 50(6):1613–1621, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hamaz:2022:OPS

[HI22] Abdelghani Hamaz and Mohamed Ibazizen. An optimal prediction in stationary random fields based on a new interpolation approach. Communications in Statistics: Simulation and Computation, 51(4):1746–1758, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hosseini:2020:ALI

[HK20] Fatemeh Hosseini and Omid Karimi. Approximate likelihood inference in spatial generalized linear mixed models with closed skew normal latent variables. Communications in Statistics: Simulation and Computation, 49(1):121–134, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hamdi:2022:PDC

[HK22] Fayçal Hamdi and Abderaouf Khalfi. Predictive density criterion for SETAR models. Communications in Statistics: Simulation and Computation, 51(2):443–459, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Hanyuda:2021:PCU

[HM21] Bunto Hanyuda and Hidetoshi Murakami. Power comparisons of the unbiased Berk–Jones test and the unbiased reversed Berk–Jones test. Communications in Statistics: Simulation and Computation, 50(4):1009–1024, 2021. CODEN CSSCDB. ISSN 0361-0918.

Homburg:2020:CEA

[Hom20] Annika Homburg. Criteria for evaluating approximations of count distributions. Communications in Statistics: Simulation and Computation, 49(12):3152–3170, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hormozinejad:2022:OTA

[Hor22] Farshin Hormozinejad. Optimal tests for arbitrarily varying object and the familywise error rates. Communications in Statistics: Simulation and Computation, 51(3):993–1001, 2022. CODEN CSSCDB. ISSN 0361-0918.

Hossain:2022:DNC

[HORA22] M. Pear Hossain, M. Hafidz Omar, Muhammad Riaz, and Sheikh Y. Arafat. On designing a new control chart for Rayleigh distributed processes with an application to monitor glass fiber strength. Communications in Statistics: Simulation and Computation, 51(6):3168–3184, 2022. CODEN CSSCDB. ISSN 0361-0918.

Han:2021:EJD

[HPZ^{+}21] Sung Won Han, Sunghoon Park, Hua Zhong, Eun-Seok Ryu, Pei Wang, Sehee Jung, Jayeon Lim, Jeewhan Yoon, and SungHwan Kim. Estimation of joint directed acyclic graphs with lasso family for gene networks. Communications in Statistics: Simulation and Computation, 50(9):2793–2807, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hasanalipour:2022:ISN

[HR22] P. Hasanalipour and M. Razmkhah. Inference on skew-normal distribution based on Fisher information in order statistics. Communications in Statistics: Simulation and Computation, 51(4):1525–1541, 2022. CODEN CSSCDB. ISSN 0361-0918.

Haq:2020:EGC

[HU20] Abdul Haq and Zain Ul Abidin. An enhanced GWMA chart for process mean. Communications in Statistics: Simulation and Computation, 49(4):847–866, 2020. CODEN CSSCDB. ISSN 0361-0918.

Huang:2020:CFR

[Hua20] Bai Huang. Combined fixed and random effects estimators. Communications in Statistics: Simulation and Computation, 49(8):1945–1956, 2020. CODEN CSSCDB. ISSN 0361-0918.

Hutson:2021:IEL

[Hut21] Alan D. Hutson. Inference for L-estimators of location using a bootstrap warping approach. Communications in Statistics: Simulation and Computation, 50(7):2145–2150, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Hartog:2021:NBL

[HvZ21] Jarno Hartog and Harry van Zanten. Nonparametric Bayesian label prediction on a large graph using truncated Laplacian regularization. Communications in Statistics: Simulation and Computation, 50(12):3821–3838, 2021. CODEN CSSCDB. ISSN 0361-0918.

Hung:2021:MSU

[HYSC21] Wen-Liang Hung, Jenn-Hwai Yang, I-Wen Song, and Yen-Chang Chang. A modified self-updating clustering algorithm for application to dengue gene expression data. Communications in Statistics: Simulation and Computation, 50(2):483– 500, 2021. CODEN CSSCDB. ISSN 0361-0918.

Harfouche:2022:DMA

[HZAF22] L. Harfouche, N. Zougab, S. Adjabi, and B. Funke. Discrete multivariate associated kernel estimators using two multiplicative bias correction methods. Communications in Statistics: Simulation and Computation, 51(2):404–420, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Ijaz:2021:EPL

[IAG21] Musarrat Ijaz, Zahid Asghar, and Asma Gul. Ensemble of penalized logistic models for classification of high-dimensional data. Communications in Statistics: Simulation and Computation, 50(7):2072–2088, 2021. CODEN CSSCDB. ISSN 0361-0918.

Islam:2021:SAE

[IC21] Sadikul Islam and Hukum Chandra. Small area estimation combining data from two surveys. Communications in Statistics: Simulation and Computation, 50(6):1728–1749, 2021. CODEN CSSCDB. ISSN 0361-0918.

Iachina:2020:SBC

[IM20] Maria Iachina and Peter Morling. A simulation-based comparison of two methods for determining the treatment effect in children diagnosed with hyperkinetic disorder. Communications in Statistics: Simulation and Computation, 49(6): 1385–1396, 2020. CODEN CSSCDB. ISSN 0361-0918.

Izaguirre:2021:HTE

[IMR21] Alejandro Izaguirre and Gabriel Montes-Rojas. Horvitz– Thompson estimator under partial information with an application to network degree distribution. Communications in Statistics: Simulation and Computation, 50(2):343–366, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ivkovic:2021:BCI

[IR21] Ivana Ivković and Vesna Rajić. Better confidence intervals for the population coefficient of variation. Communications in Statistics: Simulation and Computation, 50(12):4215–4262, 2021. CODEN CSSCDB. ISSN 0361-0918.

Iriany:2021:CGM

[IRA21] Atiek Iriany, Diana Rosyida, and Arifin Arifin. A comparison of GSTAR-SUR models and a hybrid GSTAR-SUR/neural network model on residuals of precipitation forecasting. Communications in Statistics: Simulation and Computation, 50 (9):2782–2792, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ivkovic:2020:CPC

[IRS20] Ivana Ivković, Vesna Rajić, and Jelena Stanojević. Coverage probabilities of confidence intervals for the slope parameter of linear regression model when the error term is not normally distributed. Communications in Statistics: Simulation and Computation, 49(1):147–158, 2020. CODEN CSSCDB. ISSN 0361-0918.

Inan:2020:PSO

[IS20] Deniz Inan and Nuriye Sancar. Particle swarm optimization based ridge logistic estimator. Communications in Statistics: Simulation and Computation, 49(3):669–683, 2020. CODEN CSSCDB. ISSN 0361-0918.

Iqbal:2021:BIM

[IT21] Farhat Iqbal and Kostas Triantafyllopoulos. Bayesian inference of multivariate rotated GARCH models with skew returns. Communications in Statistics: Simulation and Computation, 50(10):3105–3123, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jamshidi:2021:DNI

[JAP21] Babak Jamshidi, Sayed Mohammad Reza Alavi, and Gholam Ali Parham. The distribution of the number of the infected individuals in a stochastic SIR model on regular rooted trees. Communications in Statistics: Simulation and Computation, 50(5):1427–1444, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jorge-Gonzalez:2021:SDT

[JGGDMRLD21] Elisa Jorge-González, Enrique González-Dávila, Raquel Martín-Rivero, and Domingo Lorenzo-Díaz. Stochastic and deterministic trend in state space models. Communications in Statistics: Simulation and Computation, 50(10):2809–2822, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2021:EMP

[JHR21] Wenjun Jiang, H. P. Hong, and Jiandong Ren. Estimation of model parameters of dependent processes constructed using Lévy Copulas. Communications in Statistics: Simulation and Computation, 50(3):691–707, 2021. CODEN CSSCDB. ISSN 0361-0918.

Jayakumar:2020:NGP

[JKH20] K. Jayakumar, Bindu Krishnan, and G. G. Hamedani. On a new generalization of Pareto distribution and its applications. Communications in Statistics: Simulation and Computation, 49(5):1264–1284, 2020. CODEN CSSCDB. ISSN 0361-0918.

John:2020:CEB

[JM20] Mateusz John and Adam Mieldzioc. The comparison of the estimators of banded Toeplitz covariance structure under the high-dimensional multivariate model. Communications in Statistics: Simulation and Computation, 49(3):734–752, 2020. CODEN CSSCDB. ISSN 0361-0918.

Jin:2021:AMR

[JPH21] Man Jin, Adam Polis, and Jonathan Hartzel. Algorithms for minimization randomization and the implementation with an R package. Communications in Statistics: Simulation and Computation, 50(10):3077–3087, 2021. CODEN CSSCDB. ISSN 0361-0918.

Javaid:2020:NMH

[JuAH20] Amjad Javaid, Muhammad Noor ul Amin, and Muhammad Hanif. A new Max-HEWMA control chart using auxiliary information. Communications in Statistics: Simulation and Computation, 49(5):1285–1305, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kalina:2021:CME

[Kal21] Jan Kalina. Common multivariate estimators of location and scatter capture the symmetry of the underlying distribution. Communications in Statistics: Simulation and Computation, 50(10):2845–2857, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kahrari:2020:BPM

[KAVR20] Fereshte Kahrari, Reinaldo Boris Arellano-Valle, and Majid Rezaei. Bias prevention of maximum likelihood estimates for skew-normal-Cauchy distribution. Communications in Statistics: Simulation and Computation, 49(1):1–15, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kohansal:2021:EPK

[KB21] Akram Kohansal and Hassan S. Bakouch. Estimation procedures for Kumaraswamy distribution parameters under adaptive type-II hybrid progressive censoring. Communications in Statistics: Simulation and Computation, 50(12):4059–4078, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kim:2021:CCM

[KBR21] Jong-Min Kim, Jaiwook Baik, and Mitch Reller. Control charts of mean and variance using copula Markov SPC and conditional distribution by copula. Communications in Statistics: Simulation and Computation, 50(1):85–102, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Kwessi:2022:ANN

[KE22] Eddy Kwessi and Lloyd J. Edwards. Artificial neural networks with a signed-rank objective function and applications. Communications in Statistics: Simulation and Computation, 51(6):3363–3388, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kizilaslan:2020:IFR

[KEE20] Busenur Kizilaslan, Erol Egrioglu, and Atif Ahmet Evren. Intuitionistic fuzzy ridge regression functions. Communications in Statistics: Simulation and Computation, 49(3):699– 708, 2020. CODEN CSSCDB. ISSN 0361-0918.

Khan:2020:RRE

[KGF20] Hina Khan, Sat Gupta, and Habiba Farhat. A ratio-cum-regression estimator of population mean in unequal probability sampling design. Communications in Statistics: Simulation and Computation, 49(2):484–490, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Krit:2021:GFT

[KGR21] Meryam Krit, Olivier Gaudoin, and Emmanuel Remy. Goodness-of-fit tests for the Weibull and extreme value distributions: A review and comparative study. Communications in Statistics: Simulation and Computation, 50(7):1888–1911, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kim:2021:DIC

[Kim21] Chanmin Kim. Deviance information criteria for mixtures of distributions. Communications in Statistics: Simulation and Computation, 50(10):2935–2948, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kazemi:2020:SSI

[KJ20] M. R. Kazemi and A. A. Jafari. Small sample inference for the common coefficient of variation. Communications in Statistics: Simulation and Computation, 49(1):226–243, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Kim:2020:LLM

[KJK20] Seongyong Kim, Saebom Jeon, and Daeyoung Kim. On log– linear modeling for an incomplete two-way contingency table with one variable subject to nonresponse. Communications in Statistics: Simulation and Computation, 49(4):973–988, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kordasiabi:2022:PNU

[KK22] S. Rahnamay Kordasiabi and S. Khazaei. Prediction of the nonsampled units in survey design with the finite population using Bayesian nonparametric mixture model. Communications in Statistics: Simulation and Computation, 51(6):3457– 3470, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kargapolova:2021:SMJ

[KKO21] N. A. Kargapolova, E. I. Khlebnikova, and V. A. Ogorodnikov. Stochastic models of joint non-stationary time-series of air temperature, relative humidity and atmospheric pressure. Communications in Statistics: Simulation and Computation, 50(12):3972–3983, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kargin:2021:MCS

[KKZ21] B. A. Kargin, E. G. Kablukova, and P. Zhen. Monte Carlo simulation of the optical radiation transfer process in stochastic scattering media. Communications in Statistics: Simulation and Computation, 50(12):3984–3991, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kim:2021:RCS

[KL21] Hanwool Kim and Sangyeol Lee. On residual CUSUM statistic for PINAR(1) model in statistical design and diagnostic of control chart. Communications in Statistics: Simulation and Computation, 50(5):1290–1314, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kang:2021:PPU

[KLK21] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Posterior properties under matching priors for generalized gamma distribution. Communications in Statistics: Simulation and Computation, 50(4):1202–1216, 2021. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2022:NME

[KLM^{+}22] Christos Koukouvinos, Angeliki Lappa, Marilena Mitrouli, Paraskevi Roupa, and Ondrej Turek. Numerical methods for estimating the tuning parameter in penalized least squares problems. Communications in Statistics: Simulation and Computation, 51(4):1542–1563, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kim:2020:ITV

[KM20] Tae-Hwan Kim and Christophe Muller. Inconsistency transmission and variance reduction in two-stage quantile regression. Communications in Statistics: Simulation and Computation, 49(4):1044–1077, 2020. CODEN CSSCDB. ISSN 0361-0918.

Khedmati:2022:PRP

[KN22] Majid Khedmati and Seyed Taghi Akhavan Niaki. Phase-I robust parameter estimation of simple linear profiles in multistage processes. Communications in Statistics: Simulation and Computation, 51(2):460–485, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kuran:2021:MRC

[KÖ21a] Özge Kuran and M. Revan Özkale. Marginal ridge conceptual predictive model selection criterion in linear mixed models. Communications in Statistics: Simulation and Computation, 50(2):581–607, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kuran:2021:SRL

[KÖ21b] Özge Kuran and M. Revan Özkale. Stochastic restricted Liu predictors in linear mixed models. Communications in Statistics: Simulation and Computation, 50(9):2561–2580, 2021. CODEN CSSCDB. ISSN 0361-0918.

Khalfi:2020:SRA

[KOT20] Linda Khalfi and Megdouda Ourbih-Tari. Stochastic risk analysis in Monte Carlo simulation: a case study. Communications in Statistics: Simulation and Computation, 49(11): 3041–3053, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kumari:2021:BPE

[KP21] Neera Kumari and Ranjita Pandey. On Bayesian parameter estimation of beta log Weibull distribution under Type-II censoring. Communications in Statistics: Simulation and Computation, 50(3):650–668, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kalaivani:2021:PEE

[KR21] S. Kalaivani and J. Ravichandran. Performance evaluation of exponential distribution using Six Sigma-based tail probabilities. Communications in Statistics: Simulation and Computation, 50(11):3125–3145, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2022:ZIH

[KR22] C. Satheesh Kumar and Rakhi Ramachandran. On zeroinflated hyper-Poisson distribution and its applications. Communications in Statistics: Simulation and Computation, 51 (3):868–881, 2022. CODEN CSSCDB. ISSN 0361-0918.

Korzenowski:2020:QMS

[KS20a] A. L. Korzenowski and W. L. Simões. Quality monitoring by special charts for highly customized production systems. Communications in Statistics: Simulation and Computation, 49(8):2181–2192, 2020. CODEN CSSCDB. ISSN 0361-0918.

Krzysko:2020:RMF

[KS20b] Mirosław Krzyśko and Lukasz Smaga. Robust multivariate functional discriminant coordinates. Communications in Statistics: Simulation and Computation, 49(3):717–733, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2020:SPA

[KS20c] C. Satheesh Kumar and S. Sreejakumari. On some properties and applications of intervened cluster negative binomial distribution. Communications in Statistics: Simulation and Computation, 49(9):2357–2371, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kulasekera:2022:MRB

[KS22] K. B. Kulasekera and Chathura Siriwardhana. Multiresponse based personalized treatment selection with data from crossover designs for multiple treatments. Communications in Statistics: Simulation and Computation, 51(2):554– 569, 2022. CODEN CSSCDB. ISSN 0361-0918.

Kazemi:2021:EPL

[KSAR21] M. Kazemi, D. Shahsavani, M. Arashi, and P. C. Rodrigues. Estimation in partial linear model with spline modal function. Communications in Statistics: Simulation and Computation, 50(11):3256–3272, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kessentini:2020:RPD

[KTZ20] Sameh Kessentini, Mariem Tounsi, and Raoudha Zine. The Riesz probability distribution: Generation and EM algorithm. Communications in Statistics: Simulation and Computation, 49(8):2114–2133, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kashif:2021:IDP

[KUA21] Muhammad Kashif, Muhammad Aman Ullah, and Muhammad Aslam. Influential diagnostics with Pena’s statistic for the modified ridge regression. Communications in Statistics: Simulation and Computation, 50(12):3870–3884, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Khan:2020:MST

[KuIK20] Waqar Muhammad Khan and Asad ul Islam Khan. Most stringent test of independence for time series. Communications in Statistics: Simulation and Computation, 49(11): 2808–2826, 2020. CODEN CSSCDB. ISSN 0361-0918.

Kurnaz:2021:PSHa

[Kur21a] Fatma Sevinç Kurnaz. Performances of some high dimensional regression methods. Communications in Statistics: Simulation and Computation, 50(6):1820–1836, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Kurnaz:2021:PSHb

[Kur21b] Fatma Sevinç Kurnaz. Performances of some high dimensional regression methods: sparse principal component regression. Communications in Statistics: Simulation and Computation, 50(9):2529–2543, 2021. CODEN CSSCDB. ISSN 0361-0918.

Kurtoglu:2021:MRP

[Kur21c] Fikriye Kurtoglu. Modified ridge parameter estimators for log-gamma model: Monte Carlo evidence with a graphical investigation. Communications in Statistics: Simulation and Computation, 50(7):2168–2183, 2021. CODEN CSSCDB. ISSN 0361-0918.

Khalil:2021:MES

[KZG21] Sadia Khalil, Qi Zhang, and Sat Gupta. Mean estimation of sensitive variables under measurement errors using optional RRT models. Communications in Statistics: Simulation and Computation, 50(5):1417–1426, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Larsson:2020:TIE

[Lar20] Rolf Larsson. Testing for INAR effects. Communications in Statistics: Simulation and Computation, 49(10):2745–2764, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lavine:2021:WFA

[Lav21] Michael Lavine. WHIM: function approximation where it matters. Communications in Statistics: Simulation and Computation, 50(12):3839–3869, 2021. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:CIE

[LAZ21] Pengfei Li, Oludotun J. Akinlawon, and Shengli Zhao. Controlling individual and experimentwise error rates in replicated regular two-level factorial experiments. Communications in Statistics: Simulation and Computation, 50(6):1770– 1790, 2021. CODEN CSSCDB. ISSN 0361-0918.

Lee:2020:ACL

[Lee20] Seung-Hwan Lee. Assessing the censored linear regression model using martingale approximation. Communications in Statistics: Simulation and Computation, 49(2):408–418, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lin:2021:EUC

[LES21] Wei-Cheng Lin, Takeshi Emura, and Li-Hsien Sun. Estimation under copula-based Markov normal mixture models for serially correlated data. Communications in Statistics: Simulation and Computation, 50(12):4483–4515, 2021. CODEN CSSCDB. ISSN 0361-0918.

Levada:2021:CFB

[Lev21] Alexandre L. M. Levada. Closed-form Bayesian image denoising: improving the adaptive Wiener filter through pairwise Gaussian–Markov random fields. Communications in Statistics: Simulation and Computation, 50(4):1094–1118, 2021. CODEN CSSCDB. ISSN 0361-0918.

Lopez-Fidalgo:2022:EOE

[LFMMRH22] Jesús López-Fidalgo, Raúl Martín-Martín, and Mercedes Rodríguez-Hernández. Estimators and D-optimal experimental designs for mixtures of binary responses. Communications in Statistics: Simulation and Computation, 51(4):1488–1502, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lu:2021:NRE

[LFZ21] Yan Lu, Ye Fu, and Guoyi Zhang. Nonparametric regression estimators in dual frame surveys. Communications in Statistics: Simulation and Computation, 50(3):854–864, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Long:2021:PRM

[LG21] Yang Long and Deng Guohe. A perturbed risk model with constant interest and periodic barrier dividend strategy. Communications in Statistics: Simulation and Computation, 50 (8):2467–2481, 2021. CODEN CSSCDB. ISSN 0361-0918.

Li:2020:LRG

[LH20] Baohua Li and Lixia Hu. Laplacian regularized generalized Dirichlet mixture distribution for data clustering. Communications in Statistics: Simulation and Computation, 49(1): 16–28, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lee:2020:PLD

[LJC^{+}20] JungJun Lee, Jae-Hwan Jhong, Young-Rae Cho, SungHwan Kim, and Ja-Yong Koo. Penalized log-density estimation using Legendre polynomials. Communications in Statistics: Simulation and Computation, 49(11):2844–2860, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Lee:2021:DSN

[LK21a] Ming Ha Lee and Michael B. C. Khoo. Double sampling np chart with estimated process parameter. Communications in Statistics: Simulation and Computation, 50(8):2232–2250, 2021. CODEN CSSCDB. ISSN 0361-0918.

Levene:2021:ESJ

[LK21b] Mark Levene and Aleksejus Kononovicius. Empirical survival Jensen-Shannon divergence as a goodness-of-fit measure for maximum likelihood estimation and curve fitting. Communications in Statistics: Simulation and Computation, 50(11): 3751–3767, 2021. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:ELB

[LLL22] Xiaohui Liu, Yuzi Liu, and Fucai Lu. Empirical likelihoodbased unified confidence region for a predictive regression model. Communications in Statistics: Simulation and Computation, 51(5):2122–2139, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:QOP

[LLLL22] Wenhan Li, Lixia Liu, Cuixiang Li, and Guiwen Lv. Quanto option pricing with a jump diffusion process. Communications in Statistics: Simulation and Computation, 51(5):2095–2109, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:SAS

[LLZ22] Dong Li, Muyi Li, and Lianbin Zeng. Simulation and application of subsampling for threshold autoregressive movingaverage models. Communications in Statistics: Simulation and Computation, 51(5):2110–2121, 2022. CODEN CSSCDB. ISSN 0361-0918.

Liu:2020:SEI

[LMP20] Huilan Liu, Junjie Ma, and Changgen Peng. Shrinkage estimation for identification of linear components in composite quantile additive models. Communications in Statistics: Simulation and Computation, 49(10):2678–2692, 2020. CODEN CSSCDB. ISSN 0361-0918.

Le:2020:IMC

[LPNP20] Hoa Le, Uyen Pham, Phuong Nguyen, and The Bao Pham. Improvement on Monte Carlo estimation of HPD intervals.

Communications in Statistics: Simulation and Computation, 49(8):2164–2180, 2020. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:AEF

[LPP21] Kexuan Li, Aleksey S. Polunchenko, and Andrey Pepelyshev. Analytic evaluation of the fractional moments for the quasistationary distribution of the Shiryaev martingale on an interval. Communications in Statistics: Simulation and Computation, 50(9):2705–2720, 2021. CODEN CSSCDB. ISSN 0361-0918.

Louzada:2020:EPD

[LRF20] Francisco Louzada, Pedro Luiz Ramos, and Paulo Henrique Ferreira. Exponential-Poisson distribution: estimation and applications to rainfall and aircraft data with zero occurrence. Communications in Statistics: Simulation and Computation, 49(4):1024–1043, 2020. CODEN CSSCDB. ISSN 0361-0918.

Lodhi:2021:EPT

[LTR21] Chandrakant Lodhi, Yogesh Mani Tripathi, and Manoj Kumar Rastogi. Estimating the parameters of a truncated normal distribution under progressive type II censoring. Communications in Statistics: Simulation and Computation, 50 (9):2757–2781, 2021. CODEN CSSCDB. ISSN 0361-0918.

Lui:2022:NPC

[Lui22] Kung-Jong Lui. Notes on power comparison between the sequential parallel comparison and other commonly-used designs. Communications in Statistics: Simulation and Computation, 51(4):1783–1800, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:IMM

[LW21] Zhouping Li and Hui Wang. Improving K-means method via shrinkage estimation and LVQ algorithm. Communications in Statistics: Simulation and Computation, 50(11):3166–3181, 2021. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:CFE

[LW22a] Changbiao Liu and Jun Wang. A computationally fast estimator for semiparametric multinomial choice model. Communications in Statistics: Simulation and Computation, 51 (6):3355–3362, 2022. CODEN CSSCDB. ISSN 0361-0918.

Luo:2022:SIS

[LW22b] Guowang Luo and Mixia Wu. Statistical inference for semiparametric varying -coefficient spatial autoregressive models under restricted conditions. Communications in Statistics: Simulation and Computation, 51(5):2268–2286, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Liu:2022:CFQ

[LWC22] Weiyi Liu, Die Wan, and Wei Chen. A closed-form quasimaximum likelihood estimator of bid–ask spread. Communications in Statistics: Simulation and Computation, 51(3): 1066–1079, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:ABS

[LX22] Shengnan Li and Xiaoming Xu. Anticipated backward stochastic differential equations and their applications to zero-sum stochastic differential games. Communications in Statistics: Simulation and Computation, 51(5):2363–2380, 2022. CODEN CSSCDB. ISSN 0361-0918.

Lu:2020:ELD

[LY20] Weixue Lu and Shijuan Yang. Empirical likelihood of the distribution function in the finite point under φ-mixing samples. Communications in Statistics: Simulation and Computation, 49(10):2533–2545, 2020. CODEN CSSCDB. ISSN 0361-0918.

Li:2022:EFL

[LYP22] Bin Li, Qingzhao Yu, and Lu Peng. Ensemble of fast learning stochastic gradient boosting. Communications in Statistics: Simulation and Computation, 51(1):40–52, 2022. CODEN CSSCDB. ISSN 0361-0918.

Li:2021:NEV

[LYY21] Ning Li, Hu Yang, and Jing Yang. Nonnegative estimation and variable selection via adaptive elastic-net for highdimensional data. Communications in Statistics: Simulation and Computation, 50(12):4263–4279, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Liya:2021:GEE

[LYZ^{+}21] Fu Liya, Zhuoran Yang, Jun Zhang, Anle Long, and Yan Zhou. Generalized estimating equations for analyzing multivariate survival data. Communications in Statistics: Simulation and Computation, 50(10):3060–3068, 2021. CODEN CSSCDB. ISSN 0361-0918.

Liu:2022:FOR

[LZ22] Jie Liu and Haixiang Zhang. First-order random coefficient INAR process with dependent counting series. Communications in Statistics: Simulation and Computation, 51(6):3341– 3354, 2022. CODEN CSSCDB. ISSN 0361-0918.

Majid:2021:ADL

[MAAA21] Abdul Majid, Muhammad Aslam, Saima Altaf, and Muhammad Amanullah. Addressing the distributed lag models with heteroscedastic errors. Communications in Statistics: Simulation and Computation, 50(12):4464–4482, 2021. CODEN CSSCDB. ISSN 0361-0918.

Magis:2020:NCC

[Mag20] David Magis. A note on converting calibrated parameters of item response theory models. Communications in Statistics: Simulation and Computation, 49(12):3065–3079, 2020. CODEN CSSCDB. ISSN 0361-0918.

Manoharan:2020:IML

[MAMA20] Thirunanthini Manoharan, Jayanthi Arasan, Habshah Midi, and Mohd Bakri Adam. Influential measures on log-normal model for left-truncated and case-k interval censored data with time-dependent covariate. Communications in Statistics: Simulation and Computation, 49(6):1445–1466, 2020. CODEN CSSCDB. ISSN 0361-0918.

Mogouie:2022:SMB

[MASA22] Hamed Mogouie, Gholam Ali Raissi Ardali, Ehsan Bahrami Samani, and Amirhossein Amiri. Statistical monitoring of binary response attributed social networks considering random effects. Communications in Statistics: Simulation and Computation, 51(3):973–992, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mitra:2021:SIB

[MB21] Debanjan Mitra and Narayanaswamy Balakrishnan. Statistical inference based on left truncated and interval censored data from log-location-scale family of distributions. Communications in Statistics: Simulation and Computation, 50(4): 1073–1093, 2021. CODEN CSSCDB. ISSN 0361-0918.

Monfared:2022:EPB

[MBBS22] M. Mohammadi Monfared, Reza Arabi Belaghi, M. H. Behzadi, and Sukhdev Singh. Estimation and prediction based on type-I hybrid censored data from the Poisson–Exponential distribution. Communications in Statistics: Simulation and Computation, 51(5):2560–2585, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mefleh:2021:PBB

[MBDK21] Aline Mefleh, Romain Biard, Clément Dombry, and Zaher Khraibani. Permutation bootstrap and the block maxima method. Communications in Statistics: Simulation and Computation, 50(1):295–311, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mir:2022:PMT

[MD22] Youness Mir and François Dubeau. A piecewise model for two-phase growth phenomena. Communications in Statistics: Simulation and Computation, 51(1):99–115, 2022. CODEN CSSCDB. ISSN 0361-0918.

MeSilva:2022:PMR

[MdN22] W. V. M.e Silva and F. F. do Nascimento. MCMC4Extremes: an R package for Bayesian inference for extremes and its extensions. Communications in Statistics: Simulation and Computation, 51(2):432–442, 2022. CODEN CSSCDB. ISSN 0361-0918.

Medvedev:2022:UMV

[Med22] Ilia N. Medvedev. Universal modification of vector weighted method of dependent trials. Communications in Statistics: Simulation and Computation, 51(2):518–526, 2022. CODEN CSSCDB. ISSN 0361-0918.

Momeni:2021:ENM

[MES21] Reza Momeni, Javad Etminan, and M. Khanjari Sadegh. Estimation of normal means in the tree order model by the weighting methods. Communications in Statistics: Simulation and Computation, 50(1):282–294, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Mostel:2021:PET

[MFPP21] Linda Möstel, Matthias Fischer, Fabian Pfälzner, and Marius Pfeuffer. Parameter estimation of Tukey-type distributions: a comparative analysis. Communications in Statistics: Simulation and Computation, 50(4):957–992, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mehrdoust:2021:PMA

[MFS21] Farshid Mehrdoust, Somayeh Fallah, and Oldouz Samimi. Pricing multi-asset American option under Heston–CIR diffusion model with jumps. Communications in Statistics: Simulation and Computation, 50(11):3182–3193, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ma:2021:PET

[MG21] Yanbin Ma and Wenhao Gui. Point estimation and two new goodness of fit tests for the scale family based on general progressively Type-II censored samples. Communications in Statistics: Simulation and Computation, 50(12):4187–4214, 2021. CODEN CSSCDB. ISSN 0361-0918.

Milosevic:2020:AEG

[Mil20] Bojana Milosević. Asymptotic efficiency of goodness-of-fit tests based on Too–Lin characterization. Communications in Statistics: Simulation and Computation, 49(8):2082–2101, 2020. CODEN CSSCDB. ISSN 0361-0918.

Miller:2021:MPP

[Mil21] Curtis Miller. Marginal probabilities and point estimation for conditionally specified logistic regression. Communications in Statistics: Simulation and Computation, 50(12):4338–4363, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mishra:2020:OOS

[Mis20] Nutan Mishra. Optimal one-step censoring schemes under entropy criterion. Communications in Statistics: Simulation and Computation, 49(8):2068–2081, 2020. CODEN CSSCDB. ISSN 0361-0918.

Malekzadeh:2020:IEM

[MJ20] A. Malekzadeh and A. A. Jafari. Inference on the equality means of several two-parameter exponential distributions under progressively Type II censoring. Communications in Statistics: Simulation and Computation, 49(12):3196–3211, 2020. CODEN CSSCDB. ISSN 0361-0918.

Maiti:2021:EPR

[MK21] Kousik Maiti and Suchandan Kayal. Estimation of parameters and reliability characteristics for a generalized Rayleigh distribution under progressive type-II censored sample. Communications in Statistics: Simulation and Computation, 50 (11):3669–3698, 2021. CODEN CSSCDB. ISSN 0361-0918.

Murakami:2022:TSR

[MK22] Hidetoshi Murakami and Soshi Kawada. A test statistic with a ranking method based on the Jeffreys divergence measure. Communications in Statistics: Simulation and Computation, 51(1):266–279, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mitra:2022:ECB

[ML22] Sinjini Mitra and Kenny Le. The effect of cognitive and behavioral factors on student success in a bottleneck business statistics course via deeper analytics. Communications in Statistics: Simulation and Computation, 51(6):2779–2808, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mehmood:2022:HCC

[MLR^{+}22] Rashid Mehmood, Muhammed Hisyam Lee, Muhammad Riaz, Baber Zaman, and Iftikhar Ali. Hotelling T^{2} control chart based on bivariate ranked set schemes. Communications in Statistics: Simulation and Computation, 51(1):1–28, 2022. CODEN CSSCDB. ISSN 0361-0918.

Menezes:2020:IML

[MM20] André Felipe Berdusco Menezes and Josmar Mazucheli. Improved maximum likelihood estimators for the parameters of the Johnson S_{B} distribution. Communications in Statistics: Simulation and Computation, 49(6):1511–1526, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

MeloAlbuquerque:2022:PRD

[MM22] Pedro Henrique Melo Albuquerque and Gustavo Monteiro. RMCriteria: a decision making support system package for R. Communications in Statistics: Simulation and Computation, 51(6):3259–3271, 2022. CODEN CSSCDB. ISSN 0361-

Malela-Majika:2021:NDS

[MMMG21] J. C. Malela-Majika, C. M. Motsepa, and M. A. Graham. A new double sampling X̄ control chart for monitoring an abrupt change in the process location. Communications in Statistics: Simulation and Computation, 50(3):917–935, 2021. CODEN CSSCDB. ISSN 0361-0918.

Marozzi:2022:CIE

[MMR22] Marco Marozzi and Fernando Marmolejo-Ramos. Composite indicators in experimental psychology. an example with the semantic space of taste and shape stimuli stimuli. Communications in Statistics: Simulation and Computation, 51(1): 154–163, 2022. CODEN CSSCDB. ISSN 0361-0918.

Magnussen:2021:DCM

[MNL21] Steen Magnussen and Thomas Nord-Larsen. Designconsistent model-based variances with systematic sampling: a case study with the Danish national forest inventory. Communications in Statistics: Simulation and Computation, 50 (1):38–48, 2021. CODEN CSSCDB. ISSN 0361-0918.

Maneerat:2021:BCI

[MNN21] Patcharee Maneerat, Sa-Aat Niwitpong, and Suparat Niwitpong. Bayesian confidence intervals for a single mean and the difference between two means of delta-lognormal distributions. Communications in Statistics: Simulation and Computation, 50(10):2906–2934, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mostafaiy:2022:EVC

[Mos22] Behdad Mostafaiy. On estimation in varying coefficient models for sparse and irregularly sampled functional data. Communications in Statistics: Simulation and Computation, 51 (3):1030–1045, 2022. CODEN CSSCDB. ISSN 0361-0918.

Mandal:2020:COB

[MPD20] B. N. Mandal, Rajender Parsad, and Sukanta Dash. Construction of A-optimal balanced treatment incomplete block designs: an algorithmic approach. Communications in Statistics: Simulation and Computation, 49(6):1653–1664, 2020. CODEN CSSCDB. ISSN 0361-0918.

Monteiro:2021:BMT

[MPS21] Magda Monteiro, Isabel Pereira, and Manuel G. Scotto. Bivariate models for time series of counts: A comparison study between PBINAR models and dynamic factor models. Communications in Statistics: Simulation and Computation, 50 (7):1873–1887, 2021. CODEN CSSCDB. ISSN 0361-0918.

Marange:2021:NEL

[MQ21] Chioneso Show Marange and Yongsong Qin. A new empirical likelihood ratio goodness of fit test for normality based on moment constraints. Communications in Statistics: Simulation and Computation, 50(6):1561–1575, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mostofi:2020:DSS

[MS20] Amin Ghalamfarsa Mostofi and Sheyda Shirkani. Designing a single-sampling plan for attributes in the presence of classification errors. Communications in Statistics: Simulation and Computation, 49(7):1768–1782, 2020. CODEN CSSCDB. ISSN 0361-0918.

Makalic:2022:EAS

[MS22] Enes Makalic and Daniel F. Schmidt. An efficient algorithm for sampling from sin^{k}(x) for generating random correlation matrices. Communications in Statistics: Simulation and Computation, 51(5):2731–2735, 2022. CODEN CSSCDB. ISSN 0361-0918.

Malekzadeh:2021:QBR

[MSL21] Ahad Malekzadeh and Kamyar Sabri-Laghaie. Quantilebased reliability comparison of products: Applied to lognormal distribution. Communications in Statistics: Simulation and Computation, 50(8):2251–2265, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Moretti:2020:PBM

[MSS20] Angelo Moretti, Natalie Shlomo, and Joseph W. Sakshaug. Parametric bootstrap mean squared error of a small area multivariate EBLUP. Communications in Statistics: Simulation and Computation, 49(6):1474–1486, 2020. CODEN CSSCDB. ISSN 0361-0918.

Maboudou-Tchao:2021:HDD

[MT21] Edgard M. Maboudou-Tchao. High-dimensional data monitoring using support machines. Communications in Statistics: Simulation and Computation, 50(7):1927–1942, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Morales:2022:EAM

[MV22] Victor Hugo Morales and Jose Alberto Vargas. The effect of aggregating multivariate count data using Poisson profiles. Communications in Statistics: Simulation and Computation, 51(5):2646–2666, 2022. CODEN CSSCDB. ISSN 0361-0918.

Melo:2022:HOA

[MVLMA22] Tatiane F. N. Melo, Tiago M. Vargas, Artur J. Lemonte, and Germán Moreno-Arenas. Higher-order asymptotic refinements in the multivariate Dirichlet regression model. Communications in Statistics: Simulation and Computation, 51 (1):53–71, 2022. CODEN CSSCDB. ISSN 0361-0918.

Moretti:2020:ESA

[MW20] Angelo Moretti and Adam Whitworth. Evaluations of small area composite estimators based on the iterative proportional fitting algorithm. Communications in Statistics: Simulation and Computation, 49(12):3093–3110, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Mradula:2021:EEP

[MYVD21] Mradula, Subhash Kumar Yadav, Rahul Varshney, and Madhulika Dube. Efficient estimation of population mean under stratified random sampling with linear cost function. Communications in Statistics: Simulation and Computation, 50 (12):4364–4387, 2021. CODEN CSSCDB. ISSN 0361-0918.

Mahdizadeh:2022:IEP

[MZ22] M. Mahdizadeh and Ehsan Zamanzade. On interval estimation of the population mean in ranked set sampling. Communications in Statistics: Simulation and Computation, 51(5): 2747–2768, 2022. CODEN CSSCDB. ISSN 0361-0918.

Meng:2021:BET

[MZZF21] Lixin Meng, Jiwei Zhang, Xue Zhang, and Guozhong Feng. Bayesian estimation of time-varying parameters in ordinary differential equation models with noisy time-varying covariates. Communications in Statistics: Simulation and Computation, 50(3):708–723, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nademi:2021:SIP

[Nad21] Arash Nademi. The single-index panel data models with heterogeneous link function: mixture approach. Communications in Statistics: Simulation and Computation, 50(8):2418–2431, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nagakura:2021:CES

[Nag21a] Daisuke Nagakura. Computing exact score vectors for linear Gaussian state space models. Communications in Statistics: Simulation and Computation, 50(8):2313–2326, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Nunez-Antonio:2021:MPG

[NAG21b] Gabriel Núñez-Antonio and Emiliano Geneyro. A multivariate projected gamma model for directional data. Communications in Statistics: Simulation and Computation, 50(9): 2721–2742, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nakayama:2021:RSV

[Nak21] Yugo Nakayama. Robust support vector machine for highdimensional imbalanced data. Communications in Statistics: Simulation and Computation, 50(5):1524–1540, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Najarzadegan:2020:WBG

[NAKK20] Hossein Najarzadegan, Mohammad Hossein Alamatsaz, Iraj Kazemi, and Debasis Kundu. Weighted bivariate geometric distribution: Simulation and estimation. Communications in Statistics: Simulation and Computation, 49(9):2419–2443, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ngwa:2022:GST

[NCC^{+}22] Julius S. Ngwa, Howard J. Cabral, Debbie M. Cheng, David R. Gagnon, Michael P. LaValley, and L. Adrienne Cupples. Generating survival times with time-varying covariates using the Lambert W function. Communications in Statistics: Simulation and Computation, 51(1):135–153, 2022. CODEN CSSCDB. ISSN 0361-0918.

Nascimento:2020:DBH

[NdSA20] Abraão D. C. Nascimento, Raquel C. da Silva, and Getúlio J. A. Amaral. Distance-based hypothesis tests on the Watson distribution. Communications in Statistics: Simulation and Computation, 49(9):2225–2238, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nissas:2022:HDD

[NG22] Wassila Nissas and Soufiane Gasmi. A hybrid decision dependent maintenance model of failure rate and virtual age classes using modified Weibull intensity. Communications in Statistics: Simulation and Computation, 51(5):2397–2411, 2022. CODEN CSSCDB. ISSN 0361-0918.

Nduka:2020:EAR

[NIN20] Uchenna C. Nduka, Iheanyi S. Iwueze, and Elezar C. Nwogu. Efficient algorithms for robust estimation in autoregressive regression models using Student’s t distribution. Communications in Statistics: Simulation and Computation, 49(2): 355–374, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nishida:2021:SMV

[Nis21] Kiheiji Nishida. Skewing methods for variance-stabilizing local linear regression estimation. Communications in Statistics: Simulation and Computation, 50(7):2089–2106, 2021. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2022:EOS

[NJ22] Hadi Alizadeh Noughabi and Jalil Jarrahiferiz. Extropy of order statistics applied to testing symmetry. Communications in Statistics: Simulation and Computation, 51(6):3389–3399, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ng:2022:EES

[NKSY22] Peh Sang Ng, Michael B. C. Khoo, Sajal Saha, and Wai Chung Yeong. Economic and economic-statistical designs of auxiliary information based X̄, synthetic and EWMA charts. Communications in Statistics: Simulation and Computation, 51(3):1157–1185, 2022. CODEN CSSCDB. ISSN 0361-0918.

Novoa-Munoz:2021:GFT

[NM21] Francisco Novoa-Muñoz. Goodness-of-fit tests for the bivariate Poisson distribution. Communications in Statistics: Simulation and Computation, 50(7):1998–2014, 2021. CODEN CSSCDB. ISSN 0361-0918.

Nandy:2020:SPS

[NN20] Karabi Nandy and Rajesh Ranjan Nandy. A study of the performance of 2-stage adaptive optimal designs in a logistic dose-response model. Communications in Statistics: Simulation and Computation, 49(5):1118–1141, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nezhad:2021:DVS

[NN21] Mohammad Saber Fallah Nezhad and Marziyeh Nesaee. Developing variables sampling plans based on EWMA yield index. Communications in Statistics: Simulation and Computation, 50(4):1135–1153, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ngerng:2022:PSA

[NN22] Miang Hong Ngerng and Sherilynn S. F. Ngerng. Portfolio selection algorithm under financial crisis: a case study with Bursa Malaysia. Communications in Statistics: Simulation and Computation, 51(5):2586–2598, 2022. CODEN CSSCDB. ISSN 0361-0918.

Narjis:2020:EPP

[NS20] Ghulam Narjis and Javid Shabbir. Estimation of population proportion and sensitivity level using optional unrelated question randomized response techniques. Communications in Statistics: Simulation and Computation, 49(12):3212–3226, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nasirzadeh:2021:LFN

[NSN21] R. Nasirzadeh, A. R. Soltani, and A. R. Nematollahi. Longitudinal functional nonlinear marginal mixed effect models. Communications in Statistics: Simulation and Computation, 50(12):4079–4099, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ning:2020:RQR

[NT20] Jianhui Ning and Huiqiang Tao. Randomized quasi-random sampling/importance resampling. Communications in Statistics: Simulation and Computation, 49(12):3367–3379, 2020. CODEN CSSCDB. ISSN 0361-0918.

Nourbakhsh:2020:WET

[NYM20] Mohammadreza Nourbakhsh, Gholamhossein Yari, and Yaser Mehrali. Weighted entropies and their estimations. Communications in Statistics: Simulation and Computation, 49(5): 1142–1158, 2020. CODEN CSSCDB. ISSN 0361-0918.

Noonan:2021:ABC

[NZ21] Jack Noonan and Anatoly Zhigljavsky. Approximations for the boundary crossing probabilities of moving sums of normal random variables. Communications in Statistics: Simulation and Computation, 50(11):3547–3568, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ozdemir:2022:CEL

[ÖA22] Senay Özdemir and Olcay Arslan. Combining empirical likelihood and robust estimation methods for linear regression models. Communications in Statistics: Simulation and Computation, 51(3):941–954, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ouzzani:2021:MPI

[OB21] Fares Ouzzani and Mohamed Bentarzi. On mixture periodic Integer-Valued ARCH models. Communications in Statistics: Simulation and Computation, 50(12):3931–3957, 2021. CODEN CSSCDB. ISSN 0361-0918.

Olmus:2022:PES

[OBN22] Hülya Olmus, Esra Bespinar, and Ezgi Nazman. Performance evaluation of some propensity score matching methods by using binary logistic regression model. Communications in Statistics: Simulation and Computation, 51(4):1647–1660, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ozkale:2020:FPM

[ÖK20] M. Revan Özkale and Özge Kuran. A further prediction method in linear mixed models: Liu prediction. Communications in Statistics: Simulation and Computation, 49(12): 3171–3195, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ozbay:2021:RPS

[ÖK21] Nimet Özbay and Selahattin Kaçiranlar. Risk performance of some shrinkage estimators. Communications in Statistics:

Simulation and Computation, 50(2):323–342, 2021. CODEN CSSCDB. ISSN 0361-0918.

Osmundsen:2021:MMS

[OKO21] Kjartan Kloster Osmundsen, Tore Selland Kleppe, and Atle Oglend. MCMC for Markov-switching models — Gibbs sampling vs. marginalized likelihood. Communications in Statistics: Simulation and Computation, 50(3):669–690, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Obeidat:2022:BMT

[OLOK22] Mohammed Obeidat, Juxin Liu, Nathaniel Osgood, and Geoff Klassen. Bayesian methods for time series of count data. Communications in Statistics: Simulation and Computation, 51(2):486–504, 2022. CODEN CSSCDB. ISSN 0361-0918.

Orenti:2021:RRA

[OM21] Annalisa Orenti and Ettore Marubini. Robust regression analysis: a useful two stage procedure. Communications in Statistics: Simulation and Computation, 50(1):16–37, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Opperman:2021:SPR

[ON21] Logan Opperman and Wei Ning. Sequential probability ratio test for skew normal distribution. Communications in Statistics: Simulation and Computation, 50(10):2823–2836, 2021. CODEN CSSCDB. ISSN 0361-0918.

Olmus:2022:CPL

[ONE22] Hülya Olmus, Ezgi Nazman, and Semra Erbas. Comparison of penalized logistic regression models for rare event case. Communications in Statistics: Simulation and Computation, 51(4):1578–1590, 2022. CODEN CSSCDB. ISSN 0361-0918.

Okorn:2021:LMG

[OO21] Rok Okorn and Matjaz Omladic. Lévy modeled GMWB: Pricing with wavelets. Communications in Statistics: Simulation and Computation, 50(2):413–425, 2021. CODEN CSSCDB. ISSN 0361-0918.

Osborne:2020:DSD

[OP20] Daniel E. Osborne and Vic Patrangenaru. Detecting statistical differences between average glaucoma indices of the eye recovered from virtual 3D reconstructions of the eye’s optic nerve head using 2D stereo images. Communications in Statistics: Simulation and Computation, 49(8):2193–2205, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ozonur:2022:GFT

[OP22] Deniz Ozonur and Sudhir Paul. Goodness of fit tests of the two-parameter gamma distribution against the threeparameter generalized gamma distribution. Communications in Statistics: Simulation and Computation, 51(3):687–697, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ozbay:2020:EMM

[ÖT20] Nimet Özbay and Selma Toker. Efficiency of Mansson’s method: Some numerical findings about the role of biasing parameter in the estimation of distributed lag model. Communications in Statistics: Simulation and Computation, 49 (9):2333–2346, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ozkale:2021:RIC

[Özk21] M. Revan Özkale. The red indicator and corrected VIFs in generalized linear models. Communications in Statistics: Simulation and Computation, 50(12):4144–4170, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Pan:2020:GCC

[Pan20a] Yingli Pan. Generalized case-cohort analysis for constrained estimation in the Cox’s model. Communications in Statistics: Simulation and Computation, 49(1):45–65, 2020. CODEN CSSCDB. ISSN 0361-0918.

Panahbehagh:2020:SRC

[Pan20b] Bardia Panahbehagh. Stratified and ranked composite sampling. Communications in Statistics: Simulation and Computation, 49(2):504–515, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pant:2020:TPM

[Pan20c] Mohan D. Pant. The t-transformed power method distributions for simulating univariate and multivariate non-normal distributions. Communications in Statistics: Simulation and Computation, 49(4):825–846, 2020. CODEN CSSCDB. ISSN 0361-0918.

Papanastassiou:2020:TSM

[Pap20] Demetrios Papanastassiou. Testing for a single mean with transformed data. Communications in Statistics: Simulation and Computation, 49(2):436–452, 2020. CODEN CSSCDB. ISSN 0361-0918.

Park:2020:SMT

[Par20] Hyo-Il Park. Simultaneous multivariate tests under the normality assumption. Communications in Statistics: Simulation and Computation, 49(7):1886–1897, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Piri:2021:WAP

[PASB21] Sepehr Piri, Abdel-Salam G. Abdel-Salam, and Edward L. Boone. A wavelet approach for profile monitoring of Poisson distribution with application. Communications in Statistics: Simulation and Computation, 50(2):525–536, 2021. CODEN CSSCDB. ISSN 0361-0918.

Pekalp:2022:DBS

[PAT22] Mustafa Hilmi Pekalp, Halil Aydogdu, and Kamil Feridun Türkman. Discriminating between some lifetime distributions in geometric counting processes. Communications in Statistics: Simulation and Computation, 51(3):715–737, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Pandey:2022:CAS

[PBLM22] Arvind Pandey, Shashi Bhushan, Ralte Lalpawimawha, and Praveen Kumar Misra. Comparison of additive shared frailty models under Lindley baseline distribution. Communications in Statistics: Simulation and Computation, 51(3):1121–1136, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pan:2022:IPB

[PCWL22] Chun Pan, Bo Cai, Lianming Wang, and Xiaoyan Lin. ICBayes: a package of Bayesian semiparametric regression for intervel-censored data. Communications in Statistics: Simulation and Computation, 51(4):1852–1866, 2022. CODEN CSSCDB. ISSN 0361-0918.

Park:2021:THS

[PD21] Junyong Park and Andrei Draganescu. Testing homogeneity of several normal population means based on an interval hypothesis. Communications in Statistics: Simulation and Computation, 50(12):4114–4131, 2021. CODEN CSSCDB. ISSN 0361-0918.

Park:2020:IBC

[PDO^{+}20] Chanseok Park, Sanku Dey, Linhan Ouyang, Jai-Hyun Byun, and Mark Leeds. Improved bootstrap confidence intervals for the process capability index C_{pk}. Communications in Statistics: Simulation and Computation, 49(10):2583–2603, 2020. CODEN CSSCDB. ISSN 0361-0918.

Parra-Frutos:2021:RSK

[PFM21] Isabel Parra-Frutos and Lourdes Molera. Removing skewness and kurtosis by transformation when testing for mean equality. Communications in Statistics: Simulation and Computation, 50(9):2598–2618, 2021. CODEN CSSCDB. ISSN 0361-0918.

Pal:2022:ECP

[PG22] Surajit Pal and Susanta Kumar Gauri. Evaluating capability of a process with ordinal responses. Communications in Statistics: Simulation and Computation, 51(6):2871–2887, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pakbaz:2022:MAG

[PHN22] F. Pakbaz, F. Hosseini, and A. R. Nematollahi. Modeling additive genetic effects in animal models by closed skew normal distribution. Communications in Statistics: Simulation and Computation, 51(3):1186–1198, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pham:2021:DBS

[PHS21] Thanh Pham, Trung Ha, and Julia N. Soulakova. Designbased single-mediator approach for complex survey data. Communications in Statistics: Simulation and Computation, 50(3):822–831, 2021. CODEN CSSCDB. ISSN 0361-0918.

Pakgohar:2020:GFT

[PHY20] A. Pakgohar, A. Habibirad, and F. Yousefzadeh. Goodness of fit test using Lin-Wong divergence based on Type-I censored data. Communications in Statistics: Simulation and Computation, 49(9):2485–2504, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pekalp:2022:SID

[PIA22] Mustafa Hilmi Pekalp, Gültaç Eroglu Inan, and Halil Aydogdu. Statistical inference for doubly geometric process with Weibull interarrival times. Communications in Statistics: Simulation and Computation, 51(6):3428–3440, 2022. CODEN CSSCDB. ISSN 0361-0918.

Picinbono:2021:CFB

[Pic21] Bernard Picinbono. Coincidence functions and Bartlett spectra of point processes. Communications in Statistics: Simulation and Computation, 50(9):2581–2597, 2021. CODEN CSSCDB. ISSN 0361-0918.

Phrueksawatnon:2021:DOR

[PJL21] Piyada Phrueksawatnon, Jirawan Jitthavech, and Vichit Lorchirachoonkul. Determining the optimal ridge parameter in logistic regression. Communications in Statistics: Simulation and Computation, 50(11):3569–3580, 2021. CODEN CSSCDB. ISSN 0361-0918.

Park:2021:SKS

[PK21] Heewon Park and Sadanori Konishi. Sparse kernel subspace method for classifying and representing patterns from data with complex structure. Communications in Statistics: Simulation and Computation, 50(10):3088–3104, 2021. CODEN CSSCDB. ISSN 0361-0918.

Park:2022:EBB

[PK22] Junhui Park and Seung-Ho Kang. Effects of between-batch variability on the type I error rate in biosimilar development. Communications in Statistics: Simulation and Computation, 51(1):323–340, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pekalp:2022:EMV

[PKA22] Mustafa Hilmi Pekalp, Melike Özlem Karaduman, and Halil Aydogdu. Estimation of the mean value function for gamma trend renewal process. Communications in Statistics: Simulation and Computation, 51(6):3441–3456, 2022. CODEN CSSCDB. ISSN 0361-0918.

Park:2022:IFS

[PKW22] Chanseok Park, Haewon Kim, and Min Wang. Investigation of finite-sample properties of robust location and scale estimators. Communications in Statistics: Simulation and Computation, 51(5):2619–2645, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pasha:2020:GVB

[PM20] M. A. Pasha and M. Bameni Moghadam. A generalized version of Ben–Daya–Rahim (2000) and Rahim–Banerjee (1993) cost models in economic design of X-control charts for systems with early replacement and preventive maintenance under decreasing integrated hazard. Communications in Statistics: Simulation and Computation, 49(1):178–193, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Prajapati:2020:NDT

[PMK20] Deepak Prajapati, Sharmistha Mitra, and Debasis Kundu. A new decision theoretic sampling plan for exponential distribution under Type-I censoring. Communications in Statistics: Simulation and Computation, 49(2):453–471, 2020. CODEN CSSCDB. ISSN 0361-0918.

Patummasut:2021:REP

[PN21] Mena Patummasut and Sudarat Nidsunkid. A ratio estimator in path sampling. Communications in Statistics: Simulation and Computation, 50(2):376–387, 2021. CODEN CSSCDB. ISSN 0361-0918.

Polat:2020:EDW

[Pol20] Esra Polat. The effects of different weight functions on partial robust M-regression performance: a simulation study. Communications in Statistics: Simulation and Computation, 49 (4):1089–1104, 2020. CODEN CSSCDB. ISSN 0361-0918.

Polajnar:2022:UEN

[Pol22] Emil Polajnar. Using elastic net restricted kernel canonical correlation analysis for cross-language information retrieval. Communications in Statistics: Simulation and Computation, 51(6):2924–2941, 2022. CODEN CSSCDB. ISSN 0361-0918.

Pietrosanu:2021:HMB

[PRH21] Matthew Pietrosanu, Rhonda J. Rosychuk, and X. Joan Hu. Handling missing birthdates in marginal regression analysis with recurrent events. Communications in Statistics: Simulation and Computation, 50(1):142–152, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Panza:2020:MWR

[PV20a] Carlos Arturo Panza and Jose Alberto Vargas. Monitoring Weibull regression models with type i right-censored observations in phase II processes. Communications in Statistics: Simulation and Computation, 49(2):335–354, 2020. CODEN CSSCDB. ISSN 0361-0918.

Psaradakis:2020:NTD

[PV20b] Zacharias Psaradakis and Marián Vávra. Normality tests for dependent data: large-sample and bootstrap approaches. Communications in Statistics: Simulation and Computation, 49(2):283–304, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pan:2020:IAI

[PY20] Tianshu Pan and Yue Yin. Improving the accuracy of identifying the lognormal curve in the Johnson system. Communications in Statistics: Simulation and Computation, 49(5): 1321–1332, 2020. CODEN CSSCDB. ISSN 0361-0918.

Pourjafar:2020:EMR

[PZ20] Hojjat Pourjafar and Vali Zardasht. Estimation of the mean residual life function in the presence of measurement errors. Communications in Statistics: Simulation and Computation, 49(2):532–555, 2020. CODEN CSSCDB. ISSN 0361-0918.

Qu:2021:CDE

[QL21] Leming Qu and Yang Lu. Copula density estimation by finite mixture of parametric copula densities. Communications in Statistics: Simulation and Computation, 50(11):3315–3337, 2021. CODEN CSSCDB. ISSN 0361-0918.

Qin:2022:RDC

[QYC22] Ruibing Qin, Xiaoqin Yang, and Zhanshou Chen. Ratio detections for change point in heavy tailed observations. Communications in Statistics: Simulation and Computation, 51 (5):2487–2510, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rahimi:2021:SMM

[RAG21] Sajede Baratian Rahimi, Amirhossein Amiri, and Reza Ghashghaei. Simultaneous monitoring of mean vector and covariance matrix of multivariate simple linear profiles in the presence of within profile autocorrelation. Communications in Statistics: Simulation and Computation, 50(6):1791–1808, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rao:2021:VSP

[RAJ21] G. Srinivasa Rao, Muhammad Aslam, and Chi-Hyuck Jun. A variable sampling plan using generalized multiple dependent state based on a one-sided process capability index. Communications in Statistics: Simulation and Computation, 50(9): 2666–2677, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rasekhi:2020:SME

[Ras20] Mahdi Rasekhi. A study on methods for estimating the PDF and the CDF in the exponentiated gamma distribution. Communications in Statistics: Simulation and Computation, 49 (8):1999–2013, 2020. CODEN CSSCDB. ISSN 0361-0918.

Regis:2022:LMM

[RBN^{+}22] Marta Regis, Alberto Brini, Nazanin Nooraee, Reinder Haakma, and Edwin R. van den Heuvel. The t linear mixed model: model formulation, identifiability and estimation. Communications in Statistics: Simulation and Computation, 51(5):2318–2342, 2022. CODEN CSSCDB. ISSN 0361-0918.

Raqab:2021:IEH

[RBRAa21] Mohammad Z. Raqab, Omar M. Bdair, Manoj K. Rastogi, and Fahad M. Al-aboud. Inference for an exponentiated half logistic distribution with application to cancer hybrid censored data. Communications in Statistics: Simulation and Computation, 50(4):1178–1201, 2021. CODEN CSSCDB. ISSN 0361-0918.

RodriguesLiska:2022:MLB

[RCSS22] Gilberto Rodrigues Liska, Marcelo Ângelo Cirillo, Fortunato Silva de Menezes, and Julio Silvio de Sousa Bueno Filho. Machine learning based on extended generalized linear model applied in mixture experiments. Communications in Statistics: Simulation and Computation, 51(5):2511–2525, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Rodrigues:2020:BSP

[RdAIS^{+}20] Josemar Rodrigues, Marco Henrique de Almeida Inacio, Adriano K. Suzuki, Fernando Raimundo da Silva, and Narayanaswamy Balakrishnan. Bayesian superposition of pure-birth destructive cure processes for tumor latency. Communications in Statistics: Simulation and Computation, 49 (12):3240–3253, 2020. CODEN CSSCDB. ISSN 0361-0918.

Rezazadeh:2022:VSF

[REMO22] H. Rezazadeh, F. Eskandari, M. Bameni Moghadam, and E. Ormoz. Variable selection in finite mixture of generalized estimating equations. Communications in Statistics: Simulation and Computation, 51(6):3237–3251, 2022. CODEN CSSCDB. ISSN 0361-0918.

Roshanbin:2022:MOE

[REN22] N. Roshanbin, M. J. Ershadi, and S. T. A. Niaki. Multiobjective economic-statistical design of simple linear profiles using a combination of NSGA-II, RSM, and TOPSIS. Communications in Statistics: Simulation and Computation, 51 (4):1704–1720, 2022. CODEN CSSCDB. ISSN 0361-0918.

Razmkhah:2021:EDR

[RGA21] Mansooreh Razmkhah, Bahram Sadeghpour Gildeh, and Jafar Ahmadi. An economic design of rectifying single sampling plans via maxima nomination sampling in the presence of inspection errors. Communications in Statistics: Simulation and Computation, 50(1):217–233, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ricker:2021:SLR

[Ric21] Martin Ricker. Solving linear regression without skewness of the residuals’ distribution. Communications in Statistics: Simulation and Computation, 50(8):2482–2495, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Roozegar:2020:ACB

[RK20] Rasool Roozegar and Debasis Kundu. On absolutely continuous bivariate generalized exponential power series distribution. Communications in Statistics: Simulation and Computation, 49(7):1678–1703, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ramos:2021:NEG

[RL21] Pedro Luiz Ramos and Francisco Louzada. A note on the exponential geometric power series distribution. Communications in Statistics: Simulation and Computation, 50(11): 3794–3798, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rodrigues:2020:NAV

[RM20] Paulo Canas Rodrigues and Rahim Mahmoudvand. A new approach for the vector forecast algorithm in singular spectrum analysis. Communications in Statistics: Simulation and Computation, 49(3):591–605, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ramires:2020:FBM

[ROL^{+}20] Thiago G. Ramires, Edwin M. M. Ortega, Artur J. Lemonte, Niel Hens, and Gauss M. Cordeiro. A flexible bimodal model with long-term survivors and different regression structures. Communications in Statistics: Simulation and Computation, 49(10):2639–2660, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ramanathan:2020:SFR

[RRA20] T. V. Ramanathan, Neelabh Rohan, and Bovas Abraham. A stochastic frontier regression model with dynamic frontier. Communications in Statistics: Simulation and Computation, 49(6):1415–1428, 2020. CODEN CSSCDB. ISSN 0361-0918.

Razie:2021:AML

[RSG21] Farzaneh Razie, Ehsan Bahrami Samani, and Mojtaba Ganjali. Analysis of mixed longitudinal (k, l)-inflated power series, ordinal and continuous responses with sensitivity analysis to non-ignorable missing mechanism. Communications in Statistics: Simulation and Computation, 50(8):2286–2312, 2021. CODEN CSSCDB. ISSN 0361-0918.

Ruzankin:2022:CNM

[Ruz22] Pavel S. Ruzankin. A class of nonparametric mode estimators. Communications in Statistics: Simulation and Computation, 51(6):3291–3304, 2022. CODEN CSSCDB. ISSN 0361-0918.

Rodriguez:2022:REP

[RVV22] Daniela Rodríguez, Marina Valdora, and Pablo Vena. Robust estimation in partially linear regression models with monotonicity constraints. Communications in Statistics: Simulation and Computation, 51(4):2039–2052, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Rytova:2021:SAP

[RY21] Anastasiia Rytova and Elena Yarovaya. Survival analysis of particle populations in branching random walks. Communications in Statistics: Simulation and Computation, 50(10): 3031–3045, 2021. CODEN CSSCDB. ISSN 0361-0918.

Rahman:2020:RCC

[RYA20] Ayu Abdul Rahman, Sharipah Soaad Syed Yahaya, and Abdu Mohammed Ali Atta. Robustification of CUSUM control structure for monitoring location shift of skewed distributions based on modified one-step M-estimator. Communications in Statistics: Simulation and Computation, 49(11):3001–3018, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shabbir:2022:EIR

[SA22] Javid Shabbir and Aneel Ahmed. Estimation of interquartile range in stratified sampling under non-linear cost function. Communications in Statistics: Simulation and Computation, 51(4):1891–1898, 2022. CODEN CSSCDB. ISSN 0361-0918.

Saghir:2021:MEC

[SAA21] Aamir Saghir, Liaquat Ahmad, and Muhammad Aslam. Modified EWMA control chart for transformed gamma data. Communications in Statistics: Simulation and Computation, 50(10):3046–3059, 2021. CODEN CSSCDB. ISSN 0361-0918.

Stephenson:2021:DLP

[SAD^{+}21] Matthew Stephenson, R. Ayesha Ali, Gerarda A. Darlington, Flavio S. Schenkel, and E. James Squires. DSLRIG: Leveraging predictor structure in logistic regression. Communications in Statistics: Simulation and Computation, 50(6):1600–1612, 2021. CODEN CSSCDB. ISSN 0361-0918.

Saghir:2021:EMD

[SAF21] Aamir Saghir, Saddam Akber Abbasi, and Alireza Faraz. The exact method for designing the Maxwell chart with estimated parameter. Communications in Statistics: Simulation and Computation, 50(1):270–281, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sadoun:2020:EEP

[SB20] Mohamed Sadoun and Mohamed Bentarzi. Efficient estimation in periodic INAR(1) model: parametric case. Communications in Statistics: Simulation and Computation, 49(8): 2014–2034, 2020. CODEN CSSCDB. ISSN 0361-0918.

Santos:2021:RIS

[SB21a] Julius Santos and Erniel Barrios. Robust inference in semiparametric spatial-temporal models. Communications in Statistics: Simulation and Computation, 50(8):2266–2285, 2021. CODEN CSSCDB. ISSN 0361-0918.

Shirke:2021:VSI

[SB21b] Digambar T. Shirke and Mahesh S. Barale. A variable sampling interval sign chart for variability based on deciles. Communications in Statistics: Simulation and Computation, 50 (5):1484–1495, 2021. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:FLT

[SBB22] G. N. Singh, D. Bhattacharyya, and A. Bandyopadhyay. Formulation of logarithmic type estimators to estimate population mean in successive sampling in presence of random non response and measurement errors. Communications in Statistics: Simulation and Computation, 51(3):901–923, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Sen:2020:IOL

[SBTP20] Tanmay Sen, Ritwik Bhattacharya, Yogesh Mani Tripathi, and Biswabrata Pradhan. Inference and optimum life testing plans based on Type-II progressive hybrid censored generalized exponential data. Communications in Statistics: Simulation and Computation, 49(12):3254–3282, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sun:2022:MSA

[SBZG22] Li Sun, Narayanaswamy Balakrishnan, Fang-Chao Zhao, and Xiao-Hui Gu. Mis-specification analysis of the impact of covariates on the diffusion coefficient in Wiener degradation process. Communications in Statistics: Simulation and Computation, 51(6):3204–3222, 2022. CODEN CSSCDB. ISSN 0361-0918.

Senturk:2020:ICM

[SCE^{+}20] Damla Sentürk, Yanjun Chen, Jason P. Estes, Luis F. Campos, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Danh V. Nguyen. Impact of case-mix measurement error on estimation and inference in profiling of health care providers. Communications in Statistics: Simulation and Computation, 49(8): 2206–2224, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shah:2022:MDS

[SCHS22] Said Farooq Shah, Salman Arif Cheema, Zawar Hussain, and Ejaz Ali Shah. Masking data: a solution to social desirability bias in paired comparison experiments. Communications in Statistics: Simulation and Computation, 51(6):3149–3167, 2022. CODEN CSSCDB. ISSN 0361-0918.

Suhail:2020:QBE

[SCK20] Muhammad Suhail, Sohail Chand, and B. M. Golam Kibria. Quantile based estimation of biasing parameters in ridge regression model. Communications in Statistics: Simulation and Computation, 49(10):2732–2744, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Suhail:2021:QBR

[SCK21] Muhammad Suhail, Sohail Chand, and B. M. Golam Kibria. Quantile-based robust ridge M-estimator for linear regression model in presence of multicollinearity and outliers. Communications in Statistics: Simulation and Computation, 50(11): 3194–3206, 2021. CODEN CSSCDB. ISSN 0361-0918.

Strait:2022:PTS

[SCK22] Justin Strait, Oksana Chkrebtii, and Sebastian Kurtek. Parallel tempering strategies for model-based landmark detection on shapes. Communications in Statistics: Simulation and Computation, 51(4):1415–1435, 2022. CODEN CSSCDB. ISSN 0361-0918.

Shirazi:2022:NEQ

[SD22] Esmaeil Shirazi and Hassan Doosti. Nonparametric estimation of a quantile density function under L_{p} risk via block thresholding method. Communications in Statistics: Simulation and Computation, 51(2):539–553, 2022. CODEN CSS-CDB. ISSN 0361-0918.

Shahriari:2021:RSL

[SFG21] Shirin Shahriari, Susana Faria, and A. Manuela Gonçalves. A robust sparse linear approach for contaminated data. Communications in Statistics: Simulation and Computation, 50 (6):1662–1678, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sharma:2021:NTC

[SG21] Kush Sharma and Davinder Kumar Garg. A new technique for comparing PBIBDs. Communications in Statistics: Simulation and Computation, 50(4):1154–1164, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sanchez:2022:CDR

[SGG22] Ainara Rodríguez Sánchez, Román Salmerón Gómez, and Catalina García. The coefficient of determination in the ridge regression. Communications in Statistics: Simulation and Computation, 51(1):201–219, 2022. CODEN CSSCDB. ISSN 0361-0918.

Shang:2020:EFS

[Sha20] Han Lin Shang. Estimation of a functional single index model with dependent errors and unknown error density. Communications in Statistics: Simulation and Computation, 49(12): 3111–3133, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sharafi:2022:ITP

[Sha22] Maryam Sharafi. Inference of the two-parameter Lindley distribution based on progressive type II censored data with random removals. Communications in Statistics: Simulation and Computation, 51(4):1967–1981, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sancar:2020:IIO

[SI20] Nuriye Sancar and Deniz Inan. Identification of influential observations based on binary particle swarm optimization in the Cox PH model. Communications in Statistics: Simulation and Computation, 49(3):567–590, 2020. CODEN CSSCDB. ISSN 0361-0918.

Seo:2022:NPE

[SK22a] Jung In Seo and Yongku Kim. Nonparametric prior elicitation for a binomial proportion. Communications in Statistics: Simulation and Computation, 51(6):2809–2821, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Singh:2022:CCE

[SK22b] Garib Nath Singh and Mohd Khalid. A composite class of estimators to deal with the issue of variance estimation under the situations of random non-response in two-occasion successive sampling. Communications in Statistics: Simulation and Computation, 51(4):1454–1473, 2022. CODEN CSSCDB. ISSN 0361-0918.

Shende:2022:MMI

[SKG22] Subhash Shende, Mohan Kale, and Nikhil Gupte. Modeling maternal infant HIV transmission having variable hazard rates with two lag time distributions. Communications in Statistics: Simulation and Computation, 51(6):2858–2870, 2022. CODEN CSSCDB. ISSN 0361-0918.

Sunecher:2020:BMC

[SKJ20] Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. BINMA(1) model with COM-Poisson innovations: Estimation and application. Communications in Statistics: Simulation and Computation, 49(6):1631–1652, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2021:SIM

[SKK21] Garib Nath Singh, Mohd Khalid, and Jong-Min Kim. Some imputation methods to deal with the problems of missing data in two-occasion successive sampling. Communications in Statistics: Simulation and Computation, 50(2):557–580, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sreelakshmi:2021:JEL

[SKS21] N. Sreelakshmi, Sudheesh K. Kattumannil, and Rituparna Sen. Jackknife empirical likelihood-based inference for S-Gini indices. Communications in Statistics: Simulation and Computation, 50(6):1645–1661, 2021. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:SAA

[SKV20a] G. N. Singh, Amod Kumar, and Gajendra K. Vishwakarma. Some alternative additive randomized response models for estimation of population mean of quantitative sensitive variable in the presence of scramble variable. Communications in Statistics: Simulation and Computation, 49(11):2785–2807, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:EPM

[SKV20b] Garib Nath Singh, Amod Kumar, and Gajendra K. Vishwakarma. Estimation of population mean of sensitive quantitative character using blank cards in randomized device. Communications in Statistics: Simulation and Computation, 49(6):1603–1630, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shen:2021:MRR

[SL21] Pao-Sheng Shen and Yi Liu. Marginal regression of recurrent gap times based on semiparametric transformation cure model. Communications in Statistics: Simulation and Computation, 50(5):1383–1397, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sun:2020:BIT

[SLE20] Li-Hsien Sun, Chang-Shang Lee, and Takeshi Emura. A Bayesian inference for time series via copula-based Markov chain models. Communications in Statistics: Simulation and Computation, 49(11):2897–2913, 2020. CODEN CSSCDB. ISSN 0361-0918.

Song:2020:MRE

[SLW20] Jingwen Song, Zhenzhou Lu, and Pengfei Wei. Measuring regional effects of model inputs with random forest. Communications in Statistics: Simulation and Computation, 49(9): 2444–2461, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shirozhan:2020:IMB

[SM20] Masoumeh Shirozhan and Mehrnaz Mohammadpour. An INAR(1) model based on the Pegram and thinning operators with serially dependent innovation. Communications in Statistics: Simulation and Computation, 49(10):2617–2638, 2020. CODEN CSSCDB. ISSN 0361-0918.

Shahnavaz:2022:HSA

[SMB22] Ali Shahnavaz, Ali M. Mosammam, and Mohammad Hassan Behzadi. Half-spectral analysis of spatial-temporal data: The case study of Iranian daily wind speed data. Communications in Statistics: Simulation and Computation, 51(3):1103–1120, 2022. CODEN CSSCDB. ISSN 0361-0918.

Shongwe:2022:STM

[SMM22] S. C. Shongwe and J.-C. Malela-Majika. Shewhart-type monitoring schemes with supplementary w-of-w runs-rules to monitor the mean of autocorrelated samples. Communications in Statistics: Simulation and Computation, 51(1):293–322, 2022. CODEN CSSCDB. ISSN 0361-0918.

Simoes:2021:ACM

[SMO21] Maria L. Simões and Paula Milheiro-Oliveira. An algorithm for the computation of mean delay of traffic at pre-timed control signals. Communications in Statistics: Simulation and Computation, 50(12):4100–4113, 2021. CODEN CSSCDB. ISSN 0361-0918.

Shi:2020:DDB

[SMWJ20] Jian Shi, Yixuan Meng, Shaoping Wang, and Zongxia Jiao. A dynamic delay-based reliability evaluation model for communication networks. Communications in Statistics: Simulation and Computation, 49(6):1397–1414, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sivena:2022:IQH

[SN22] Sofia Sivena and Yiannis Nikolaidis. Improving the quality of higher education teaching through the exploitation of student evaluations and the use of control charts. Communications in Statistics: Simulation and Computation, 51(3):1289–1312, 2022. CODEN CSSCDB. ISSN 0361-0918.

Salmasnia:2020:IQM

[SNA20] Ali Salmasnia, Mohammadreza Namdar, and Behnam Abdzadeh. An integrated quality and maintenance model for two-unit series systems. Communications in Statistics: Simulation and Computation, 49(4):886–917, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Solis:2021:NPE

[Sol21] Maikol Solís. Non-parametric estimation of the first-order Sobol indices with bootstrap bandwidth. Communications in Statistics: Simulation and Computation, 50(9):2497–2512, 2021. CODEN CSSCDB. ISSN 0361-0918.

Salmaso:2022:DEM

[SPG^{+}22] Luigi Salmaso, Luca Pegoraro, Rosa Arboretti Giancristofaro, Riccardo Ceccato, Alberto Bianchi, Silvio Restello, and Davide Scarabottolo. Design of experiments and machine learning to improve robustness of predictive maintenance with application to a real case study. Communications in Statistics: Simulation and Computation, 51(2):570–582, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:GES

[SPS22] Garib Nath Singh, Awadhesh K. Pandey, and Chandraketu Singh. Generalized estimation strategy for mean estimation on current occasion in two-occasion rotation patterns. Communications in Statistics: Simulation and Computation, 51 (4):1661–1684, 2022. CODEN CSSCDB. ISSN 0361-0918.

Silva:2020:IMU

[SRO20] Eliud Silva, Aram Ramos, and Karen G. Olvera. Inferences on mortality using the Heligman–Pollard model: the Mexican case. Communications in Statistics: Simulation and Computation, 49(12):3227–3239, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:SMC

[SS20a] Parminder Singh and Navdeep Singh. Step-up multiple comparison procedure for comparing several normal variances. Communications in Statistics: Simulation and Computation, 49(6):1592–1602, 2020. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:CSN

[SS20b] Rajvir Chauhan Singh and Parminder Singh. Comparing several normal variances with a control using sample quasi range. Communications in Statistics: Simulation and Computation, 49(2):396–407, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sarada:2021:BDA

[SS21a] Y. Sarada and R. Shenbagam. Bi-dimensional availability function and its application. Communications in Statistics: Simulation and Computation, 50(5):1333–1347, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Shi:2021:DTF

[SS21b] Jian Shi and Kai Song. A discrete-time and finite-state Markov chain based in-play prediction model for NBA basketball matches. Communications in Statistics: Simulation and Computation, 50(11):3768–3776, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sasiwannapong:2022:ECB

[SSBA22] Sirasak Sasiwannapong, Saowanit Sukparungsee, Piyapatr Busababodhin, and Yupaporn Areepong. The efficiency of constructed bivariate copulas for MEWMA and Hotelling’s T^{2} control charts. Communications in Statistics: Simulation and Computation, 51(4):1837–1851, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:MRD

[SSK22] Garib Nath Singh, Chandraketu Singh, and Amod Kumar. A modified randomized device for estimation of population mean of quantitative sensitive variable with measure of privacy protection. Communications in Statistics: Simulation and Computation, 51(4):1867–1890, 2022. CODEN CSSCDB. ISSN 0361-0918.

Salinas:2020:CUP

[SSS20] Veronica I. Salinas, Stephen A. Sedory, and Sarjinder Singh. Calibration using power transformation. Communications in Statistics: Simulation and Computation, 49(9):2256–2286, 2020. CODEN CSSCDB. ISSN 0361-0918.

Santore:2021:ACC

[STdL21] Fabiola Santore, Cesar Augusto Taconeli, and Idemauro Antonio Rodrigues de Lara. An adaptive control chart for the process location based on ranked set sampling. Communications in Statistics: Simulation and Computation, 50(11): 3364–3382, 2021. CODEN CSSCDB. ISSN 0361-0918.

Singh:2021:ECV

[SU21] Garib Nath Singh and Mahamood Usman. Efficient combination of various estimators in the presence of non-response. Communications in Statistics: Simulation and Computation, 50(8):2432–2466, 2021. CODEN CSSCDB. ISSN 0361-0918.

Sulewski:2022:MLG

[Sul22] Piotr Sulewski. Modified Lilliefors goodness-of-fit test for normality. Communications in Statistics: Simulation and Computation, 51(3):1199–1219, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2022:CEM

[SVK22] Neha Singh, Gajendra K. Vishwakarma, and Jong Min Kim. Computing the effect of measurement errors on efficient variant of the product and ratio estimators of mean using auxiliary information. Communications in Statistics: Simulation and Computation, 51(2):604–625, 2022. CODEN CSSCDB. ISSN 0361-0918.

Singh:2020:NEC

[SVM20] Rajesh Singh, Gautam Kumar Vishwakarma, and Prabhakar Mishra. A note on the estimators for coefficient of dispersion using auxiliary information. Communications in Statistics: Simulation and Computation, 49(9):2347–2356, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Speiser:2020:BTD

[SWC^{+}20] Jaime Lynn Speiser, Bethany J. Wolf, Dongjun Chung, Constantine J. Karvellas, David G. Koch, and Valerie L. Durkalski. BiMM tree: a decision tree method for modeling clustered and longitudinal binary outcomes. Communications in Statistics: Simulation and Computation, 49(4):1004–1023, 2020. CODEN CSSCDB. ISSN 0361-0918.

Song:2022:GFS

[SX22] Won Chul Song and Jun Xie. Group feature screening via the F statistic. Communications in Statistics: Simulation and Computation, 51(4):1921–1931, 2022. CODEN CSSCDB. ISSN 0361-0918.

Stuart:2022:CEM

[SY22] Matthew Stuart and Cindy Yu. A computationally efficient method for selecting a split questionnaire design. Communications in Statistics: Simulation and Computation, 51(5): 2464–2486, 2022. CODEN CSSCDB. ISSN 0361-0918.

Satter:2020:NIE

[SZ20] Faysal Satter and Yichuan Zhao. Nonparametric interval estimation for the mean of a zero-inflated population. Communications in Statistics: Simulation and Computation, 49(8): 2059–2067, 2020. CODEN CSSCDB. ISSN 0361-0918.

Sazak:2021:MML

[SZ21] Hakan Savas Sazak and Melis Zeybek. The modified maximum likelihood regression type estimators using bivariate ranked set sampling. Communications in Statistics: Simulation and Computation, 50(11):3614–3649, 2021. CODEN CSSCDB. ISSN 0361-0918.

Song:2021:NMS

[SZJG21] Zhao-Li Song, Qian Zhao, Ping Jiang, and Bo Guo. A numerical method for system residual life prediction based on Simpson formula. Communications in Statistics: Simulation and Computation, 50(12):4171–4186, 2021. CODEN CSSCDB. ISSN 0361-0918.

Song:2020:MSP

[SZS20] Kai Song, Qingrong Zou, and Jian Shi. Modelling the scores and performance statistics of NBA basketball games. Communications in Statistics: Simulation and Computation, 49 (10):2604–2616, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tighkhorshid:2022:RAE

[TAA22] Elahe Tighkhorshid, Amirhossein Amiri, and Farzad Amirkhani. A risk-adjusted EWMA chart with dynamic probability control limits for monitoring survival time. Communications in Statistics: Simulation and Computation, 51(3):1333–1354, 2022. CODEN CSSCDB. ISSN 0361-0918.

Thomson:2022:TFS

[TBH22] Trevor J. Thomson, W. John Braun, and X. Joan Hu. On the time to first spotting in wildland fires. Communications in Statistics: Simulation and Computation, 51(5):2381–2396, 2022. CODEN CSSCDB. ISSN 0361-0918.

Toka:2022:RER

[TÇA22] Onur Toka, Meral Çetin, and Olcay Arslan. Robust estimation in restricted linear regression. Communications in Statistics: Simulation and Computation, 51(3):1015–1029, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tran:2020:WIR

[TDD20] Duyet Tran, Anil Dolgun, and Haydar Demirhan. Weighted inter-rater agreement measures for ordinal outcomes. Communications in Statistics: Simulation and Computation, 49 (4):989–1003, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tellaroli:2022:SCA

[Tel22] P. Tellaroli. SingleCross-clustering: an algorithm for finding elongated clusters with automatic estimation of outliers and number of clusters. Communications in Statistics: Simulation and Computation, 51(5):2412–2428, 2022. CODEN CSSCDB. ISSN 0361-0918.

Taremi:2020:EWI

[TEM20] Mohammad Taremi, Farzad Eskandari, and Mohammad Bameni Moghadam. Effects of weak identification on the MD estimator in dynamic stochastic general equilibrium models.

Communications in Statistics: Simulation and Computation, 49(10):2661–2677, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tuzen:2020:SSC

[TEO20] Fatih Tüzen, Semra Erbas, and Hülya Olmus. A simulation study for count data models under varying degrees of outliers and zeros. Communications in Statistics: Simulation and Computation, 49(4):1078–1088, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tarima:2022:DTF

[TF22] Sergey Tarima and Nancy Flournoy. Distribution theory following blinded and unblinded sample size re-estimation under parametric models. Communications in Statistics: Simulation and Computation, 51(4):2053–2064, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tamandi:2020:FMM

[TJ20] Mostafa Tamandi and Ahad Jamalizadeh. Finite mixture modeling using shape mixtures of the skew scale mixtures of normal distributions. Communications in Statistics: Simulation and Computation, 49(12):3345–3366, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tanil:2022:MSR

[TK22a] Halil Tanil and Agah Kozan. A modified Spearman’s rho parameter-free test statistic for early detection of newly emerging phenomena. Communications in Statistics: Simulation and Computation, 51(6):2822–2830, 2022. CODEN CSSCDB. ISSN 0361-0918.

Tee:2022:EIC

[TK22b] Liivika Tee and Meelis Käärik. Estimating IBNR claim counts using different levels of data aggregation. Communications in Statistics: Simulation and Computation, 51(6):3013–3036, 2022. CODEN CSSCDB. ISSN 0361-0918.

Takam:2020:PML

[TKF^{+}20] Patrice Soh Takam, Eugene Kouassi, Renaud Fadonougbo, Jean Marcelin Bosson Brou, and Mbodja Mougoué. Pseudo maximum likelihood estimation and asymptotic results of the GARCH(1, 2) model under dependent innovations. Communications in Statistics: Simulation and Computation, 49(8): 2134–2163, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tong:2021:NEC

[TMZL21] Jinying Tong, Qingting Meng, Zhenzhong Zhang, and Yunfang Lu. A note on ergodicity for CIR model with Markov switching. Communications in Statistics: Simulation and Computation, 50(5):1445–1458, 2021. CODEN CSSCDB. ISSN 0361-0918.

Thangjai:2020:AGC

[TNN20] Warisa Thangjai, Sa-Aat Niwitpong, and Suparat Niwitpong. Adjusted generalized confidence intervals for the common coefficient of variation of several normal populations. Communications in Statistics: Simulation and Computation, 49(1): 194–206, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tunno:2022:SDD

[TP22] Ferebee Tunno and Miranda Perry. Signal discrimination without denoising. Communications in Statistics: Simulation and Computation, 51(2):626–646, 2022. CODEN CSSCDB. ISSN 0361-0918.

Turksen:2020:NML

[Tür20] Özlem Türksen. A nonlinear modeling with linear fuzzy numbers for replicated response measures. Communications in Statistics: Simulation and Computation, 49(3):629–646, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tian:2020:LBQ

[TWTT20] Yuzhu Tian, Liyong Wang, Manlai Tang, and Maozai Tian. Likelihood-based quantile mixed effects models for longitudinal data with multiple features via MCEM algorithm. Communications in Statistics: Simulation and Computation, 49 (2):317–334, 2020. CODEN CSSCDB. ISSN 0361-0918.

Tian:2021:WCQ

[TWTT21] Yuzhu Tian, Liyong Wang, Manlai Tang, and Maozai Tian. Weighted composite quantile regression for longitudinal mixed effects models with application to AIDS studies. Communications in Statistics: Simulation and Computation, 50(6):1837–1853, 2021. CODEN CSSCDB. ISSN 0361-0918.

Tan:2021:PEL

[TY21] Xiaoyan Tan and Li Yan. Penalized empirical likelihood for generalized linear models with longitudinal data. Communications in Statistics: Simulation and Computation, 50(2): 608–623, 2021. CODEN CSSCDB. ISSN 0361-0918.

Noor-ul-Amin:2022:EWM

[uARS22] Muhammad Noor ul Amin, Afshan Riaz, and Amna Safeer. Exponentially weighted moving average control chart using auxiliary variable with measurement error. Communications in Statistics: Simulation and Computation, 51(3):1002–1014, 2022. CODEN CSSCDB. ISSN 0361-0918.

Uzabaci:2020:EOD

[UEA20] Ender Uzabaci, Ilker Ercan, and Ozlem Alpu. Evaluation of outlier detection method performance in symmetric multivariate distributions. Communications in Statistics: Simulation and Computation, 49(2):516–531, 2020. CODEN CSSCDB. ISSN 0361-0918.

Khan:2022:MST

[uIKKH22] Asad ul Islam Khan, Waqar Muhammad Khan, and Mehmood Hussan. Most stringent test of null of cointegration: a Monte Carlo comparison. Communications in Statistics: Simulation and Computation, 51(4):2020–2038, 2022. CODEN CSSCDB. ISSN 0361-0918.

Ulas:2020:IDM

[UKK20] Efehan Ulas, Filiz Karaman, and Tuba Koc. Impact of different model structure and prior distribution in continual reassessment method. Communications in Statistics: Simulation and Computation, 49(1):95–109, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Utazirubanda:2021:VSG

[ULN21] Jean Claude Utazirubanda, Tomás M. León, and Papa Ngom. Variable selection with group LASSO approach: Application to Cox regression with frailty model. Communications in Statistics: Simulation and Computation, 50(3):881–901, 2021. CODEN CSSCDB. ISSN 0361-0918.

UlHassan:2022:DUM

[UM22] Mahmood Ul Hassan and Frank Miller. Discrimination with unidimensional and multidimensional item response theory models for educational data. Communications in Statistics: Simulation and Computation, 51(6):2992–3012, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

UlIslam:2022:CCM

[UR22] Tanweer Ul Islam and Mahvish Rizwan. Comparison of correlation measures for nominal data. Communications in Statistics: Simulation and Computation, 51(3):698–714, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Uyanto:2022:MCP

[Uya22] Stanislaus S. Uyanto. Monte Carlo power comparison of seven most commonly used heteroscedasticity tests. Communications in Statistics: Simulation and Computation, 51(4):2065– 2082, 2022. CODEN CSSCDB. ISSN 0361-0918.

Uozumi:2022:CID

[UYMK22] Ryuji Uozumi, Shinjo Yada, Kazushi Maruo, and Atsushi Kawaguchi. Confidence intervals for difference between two binomial proportions derived from logistic regression. Communications in Statistics: Simulation and Computation, 51 (6):3223–3236, 2022. CODEN CSSCDB. ISSN 0361-0918.

Verma:2022:RRS

[VDN22] Vivek Verma, Radhakanta Das, and Dilip C. Nath. Representativeness of ranked set sampling based on Bayesian score. Communications in Statistics: Simulation and Computation, 51(3):1080–1095, 2022. CODEN CSSCDB. ISSN 0361-0918.

Verdoy:2021:ESM

[Ver21] Pablo Juan Verdoy. Enhancing the SPDE modeling of spatial point processes with INLA, applied to wildfires. Choosing the best mesh for each database. Communications in Statistics: Simulation and Computation, 50(10):2990–3030, 2021. CODEN CSSCDB. ISSN 0361-0918.

Vishwakarma:2022:SES

[VG22] Gajendra K. Vishwakarma and Shubham Gupta. Shrinkage estimator for scale parameter of gamma distribution. Communications in Statistics: Simulation and Computation, 51 (6):3073–3080, 2022. CODEN CSSCDB. ISSN 0361-0918.

Veerman:2022:EVC

[VLvdW22] Jurre R. Veerman, Gwenaël G. R. Leday, and Mark A. van de Wiel. Estimation of variance components, heritability and the ridge penalty in high-dimensional generalized linear models.

Communications in Statistics: Simulation and Computation, 51(1):116–134, 2022. CODEN CSSCDB. ISSN 0361-0918.

vanOest:2021:UCB

[vO21] Rutger van Oest. Unconstrained Cholesky-based parametrization of correlation matrices. Communications in Statistics: Simulation and Computation, 50(11):3607–3613, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Vovan:2021:AGA

[VPT21] Tai Vovan, Dinh Phamtoan, and Dung Tranthituy. Automatic genetic algorithm in clustering for discrete elements. Communications in Statistics: Simulation and Computation, 50(6):1679–1694, 2021. CODEN CSSCDB. ISSN 0361-0918.

Varathan:2021:MAU

[VW21] Nagarajah Varathan and Pushpakanthie Wijekoon. Modified almost unbiased Liu estimator in logistic regression. Communications in Statistics: Simulation and Computation, 50(11): 3530–3546, 2021. CODEN CSSCDB. ISSN 0361-0918.

Vanlengenberg:2021:DGA

[VWZ21] Christopher D. Vanlengenberg, Wenshuang Wang, and Haimeng Zhang. Data generation for axially symmetric processes on the sphere. Communications in Statistics: Simulation and Computation, 50(6):1750–1769, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:SLR

[WCF21] Yuh-Jenn Wu, Li-Hsueh Cheng, and Wei-Quan Fang. Sparse linear regression models of high dimensional covariates with non-Gaussian outliers and Berkson error-in-variable under heteroscedasticity. Communications in Statistics: Simulation and Computation, 50(11):3146–3165, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Westgate:2020:AUM

[Wes20] Philip M. Westgate. Approaches for the utilization of multiple criteria to select a working correlation structure for use within generalized estimating equations. Communications in Statistics: Simulation and Computation, 49(2):305–316, 2020. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:FSA

[WHW21] Huiwen Wang, Tingting Huang, and Shanshan Wang. A flexible spatial autoregressive modelling framework for mixed covariates of multiple data types. Communications in Statistics: Simulation and Computation, 50(11):3498–3515, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Wehrhahn:2021:SSB

[WJB21] Claudia Wehrhahn, Alejandro Jara, and Andrés F. Barrientos. On the small sample behavior of Dirichlet process mixture models for data supported on compact intervals. Communications in Statistics: Simulation and Computation, 50 (3):786–810, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:RVS

[WJZ^{+}22] Yan Wang, Yunlu Jiang, Jiantao Zhang, Zhongran Chen, Baojian Xie, and Chengxiang Zhao. Robust variable selection based on the random quantile LASSO. Communications in Statistics: Simulation and Computation, 51(1):29–39, 2022. CODEN CSSCDB. ISSN 0361-0918.

Watters:2020:CFM

[WL20] Christine G. Watters and Lynn R. LaMotte. A comparison of four methods of inverse prediction. Communications in Statistics: Simulation and Computation, 49(3):709–716, 2020. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:IEC

[WL22] Liang Wang and Huanyu Li. Inference for exponential competing risks data under generalized progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 51(3):1255–1271, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:NAP

[WLC^{+}22] Yan Wang, Xiaoqin Li, Ling Chen, Minghui Wang, and Xuejun Wang. A note on the asymptotic properties of the estimators in a semiparametric regression model. Communications in Statistics: Simulation and Computation, 51(2):358– 373, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:IDC

[WLT22] Liang Wang, Mengyang Li, and Yogesh Mani Tripathi. Inference for dependent competing risks from bivariate Kumaraswamy distribution under generalized progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 51(6):3100–3123, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2022:JML

[WLY22] Leyang Wang, Zhiqiang Li, and Fengbin Yu. Jackknife method for the location of gross errors in weighted total least squares. Communications in Statistics: Simulation and Computation, 51(4):1946–1966, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:MID

[WMJ21] Pengyue Wu, Jing Ma, and Xiongfei Jiang. A model of innovation diffusion based on policy incentives. Communications in Statistics: Simulation and Computation, 50(9):2544–2560, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wijekularathna:2022:PAS

[WMS22] Danush K. Wijekularathna, Ananda B. W. Manage, and Stephen M. Scariano. Power analysis of several normality tests: a Monte Carlo simulation study. Communications in Statistics: Simulation and Computation, 51(3):757–773, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wiegand:2021:GMR

[WN21a] Martin Wiegand and Saralees Nadarajah. General moments of roundoff error. Communications in Statistics: Simulation and Computation, 50(10):2885–2905, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Wiegand:2021:MME

[WN21b] Martin Wiegand and Saralees Nadarajah. MEPDF: Multivariate empirical density functions. Communications in Statistics: Simulation and Computation, 50(2):367–375, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wan:2022:BDA

[WP22] Shuwen Wan and Kai Peng. A Bayesian discriminant analysis method under semiparametric density ratio models. Communications in Statistics: Simulation and Computation, 51(4): 1759–1766, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2020:RPM

[WS20] Leyang Wang and Jianqiang Sun. Regression prediction method that is based on the partial errors-in-variables model. Communications in Statistics: Simulation and Computation, 49(12):3380–3395, 2020. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:DJF

[WS21a] Yanhua Wu and Yufeng Shi. Detection of jumps in financial time series. Communications in Statistics: Simulation and Computation, 50(2):313–322, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wywial:2021:VSM

[WS21b] Janusz L. Wywiał and Grzegorz Sitek. On variance of sample matrix eigenvalue. Communications in Statistics: Simulation and Computation, 50(7):1943–1954, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wu:2021:OSM

[Wu21] Shu-Fei Wu. One stage multiple comparisons with the control for exponential mean lifetimes based on doubly censored samples under heteroscedasticity. Communications in Statistics: Simulation and Computation, 50(5):1473–1483, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2022:NSS

[WWTM22] Cong Wang, Tonghui Wang, David Trafimow, and Hunter A. Myüz. Necessary sample sizes for specified closeness and confidence of matched data under the skew normal setting. Communications in Statistics: Simulation and Computation, 51 (5):2083–2094, 2022. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:ETI

[WWYX21] Xiaohong Wang, Dehui Wang, Kai Yang, and Da Xu. Estimation and testing for the integer-valued threshold autoregressive models based on negative binomial thinning. Communications in Statistics: Simulation and Computation, 50 (6):1622–1644, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:IHD

[WX21] Zhaoliang Wang and Liugen Xue. Inference for high dimensional linear models with error-in-variables. Communications in Statistics: Simulation and Computation, 50(1):164–179, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:JRM

[WY21] Leyang Wang and Fengbin Yu. Jackknife resample method for precision estimation of weighted total least squares. Communications in Statistics: Simulation and Computation, 50 (5):1272–1289, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wan:2021:DSV

[WZ21a] Qiang Wan and Mei Zhu. Detecting a shift in variance using economically designed VSI control chart with combined attribute-variable inspection. Communications in Statistics: Simulation and Computation, 50(11):3483–3497, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:2021:SEF

[WZ21b] Yu Wang and Haixiang Zhang. Some estimation and forecasting procedures in Poisson–Lindley INAR(1) process. Communications in Statistics: Simulation and Computation, 50(1): 49–62, 2021. CODEN CSSCDB. ISSN 0361-0918.

Wang:2021:FIG

[WZY^{+}21] Xiao Wang, Chenchen Zou, Li Yi, Juan Wang, and Xinmin Li. Fiducial inference for gamma distributions: two-sample problems. Communications in Statistics: Simulation and Computation, 50(3):811–821, 2021. CODEN CSSCDB. ISSN 0361-0918.

Xue:2021:OUI

[XH21] Yishu Xue and Guanyu Hu. Online updating of information based model selection in the big data setting. Communications in Statistics: Simulation and Computation, 50(11): 3516–3529, 2021. CODEN CSSCDB. ISSN 0361-0918.

Xing:2020:RBT

[XW20] Yixun Xing and Wayne A. Woodward. R/S-bootstrapping test for fractional integration. Communications in Statistics: Simulation and Computation, 49(6):1467–1473, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Xing:2022:MIE

[XWL22] Yanchun Xing, Ma Wenqing, and Chunhui Liang. A methodology for improving efficiency estimation based on conditional mix-GEE models in longitudinal studies. Communications in Statistics: Simulation and Computation, 51(1):254–265, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xie:2022:LPR

[XXH22] Lin Xie, Hongmin Xiao, and Yan He. The limit property of a risk model based on entrance processes. Communications in Statistics: Simulation and Computation, 51(3):955–972, 2022. CODEN CSSCDB. ISSN 0361-0918.

Xu:2021:RRM

[XZJ^{+}21] Qifa Xu, Xingxuan Zhuo, Cuixia Jiang, Fang Sun, and Xue Huang. Reverse restricted MIDAS model with application to US interest rate forecasts. Communications in Statistics: Simulation and Computation, 50(2):462–482, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yavarizadeh:2021:NMR

[YA21] Bahareh Yavarizadeh and Syed Ejaz Ahmed. The new mixed ridge estimator in a linear mixed model with measurement error under stochastic linear mixed restrictions. Communications in Statistics: Simulation and Computation, 50(8):2185– 2196, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yada:2022:BAD

[Yad22] Shinjo Yada. A Bayesian adaptive design for addressing correlated late-onset outcomes in phase I/II randomized trials of drug combinations in oncology. Communications in Statistics: Simulation and Computation, 51(6):2963–2977, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yashchin:2022:GAM

[Yas22] Emmanuel Yashchin. Gradient analysis of Markov-type control schemes and its applications. Communications in Statistics: Simulation and Computation, 51(4):1814–1836, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yadav:2022:BES

[YBC22] Abhimanyu Singh Yadav, Hassan S. Bakouch, and Christophe Chesneau. Bayesian estimation of the survival characteristics for Hjorth distribution under progressive type-II censoring. Communications in Statistics: Simulation and Computation, 51(3):882–900, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yildirim:2022:MDB

[YC22] Emre Yildirim and Mehmet Ali Cengiz. Modeling dependency between industry production and energy market via stochastic copula approach. Communications in Statistics: Simulation and Computation, 51(4):2006–2019, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Yang:2021:EFD

[YDHW21] Lianqiang Yang, Mengzhen Ding, Yongmiao Hong, and Xuejun Wang. Estimating functions and derivatives via adaptive penalized splines. Communications in Statistics: Simulation and Computation, 50(7):2054–2071, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yan:2020:GFI

[YDLX20] Liang Yan, Xiaofang Dong, Xuhua Liu, and Xingzhong Xu. Generalized fiducial inference in the multiple regression model with measurement errors. Communications in Statistics: Simulation and Computation, 49(9):2239–2255, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Yada:2021:AGP

[YH21a] Shinjo Yada and Chikuma Hamada. Application of gamma process to two-agent combinations with delayed toxicity. Communications in Statistics: Simulation and Computation, 50(1):153–163, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yoon:2021:TIQ

[YH21b] Jae Eun Yoon and Sun Young Hwang. On the threshold innovation in quasi-likelihood for conditionally heteroscedastic time series. Communications in Statistics: Simulation and Computation, 50(7):2042–2053, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yazdi:2021:PIM

[YHGA21] Ahmad Ahmadi Yazdi, Ali Zeinal Hamadani, Mohammad Hosein Karimi Gavareshki, and Amirhossein Amiri. Phase II monitoring of multivariate profiles with estimated parameters and optimal phase I subgroups. Communications in Statistics: Simulation and Computation, 50(10):2858–2884, 2021. CODEN CSSCDB. ISSN 0361-0918.

Yerlikaya-Ozkurt:2020:NCM

[YÖT20] Fatma Yerlikaya-Özkurt and Pakize Taylan. New computational methods for classification problems in the existence of outliers based on conic quadratic optimization. Communications in Statistics: Simulation and Computation, 49(3):753– 770, 2020. CODEN CSSCDB. ISSN 0361-0918.

Yavarizadeh:2022:REL

[YRAB22] Bahareh Yavarizadeh, Abdolrahman Rasekh, Syed Ejaz Ahmed, and Babak Babadi. Ridge estimation in linear mixed measurement error models with stochastic linear mixed restrictions. Communications in Statistics: Simulation and Computation, 51(6):3037–3053, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yau:2020:NHT

[YV20] Tat S. Yau and Julia Volaufova. A note on hypothesis testing in the envelope model. Communications in Statistics: Simulation and Computation, 49(3):606–628, 2020. CODEN CSSCDB. ISSN 0361-0918.

Yin:2020:LAB

[YW20] Shanshan Yin and Lichun Wang. Linear approximate Bayes estimator for variance components in random effects model*. Communications in Statistics: Simulation and Computation, 49(5):1352–1366, 2020. CODEN CSSCDB. ISSN 0361-0918.

Yin:2020:NEM

[YWLD20] Junhui Yin, Liucang Wu, Hanchi Lu, and Lin Dai. New estimation in mixture of experts models using the Pearson type VII distribution. Communications in Statistics: Simulation and Computation, 49(2):472–483, 2020. CODEN CSSCDB. ISSN 0361-0918.

Ye:2022:NEB

[YWT22] Gen Ye, Songjian Wang, and Niansheng Tang. Nonparametric empirical Bayesian method for noncontractual setting of customer-base analysis. Communications in Statistics: Simulation and Computation, 51(3):1234–1254, 2022. CODEN CSSCDB. ISSN 0361-0918.

Yuan:2021:TPP

[YZ21] Mingao Yuan and Yue Zhang. Test for the parametric part in partial functional linear regression based on B-spline. Communications in Statistics: Simulation and Computation, 50 (1):1–15, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:SAC

[ZA21] Qing Zhang and Hongshik Ahn. Subgroup analysis of censored data on cancer treatment. Communications in Statistics: Simulation and Computation, 50(12):4041–4058, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2020:BSI

[ZB20] Ningning Zhao and Zhidong Bai. Bayesian statistical inference based on rounded data. Communications in Statistics: Simulation and Computation, 49(1):135–146, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2020:EPL

[ZBZS20] Xiaojun Zhu, Narayanaswamy Balakrishnan, Yiliang Zhou, and Hon-Yiu So. Exact predictive likelihood inference for Laplace distribution based on a time-constrained experiment. Communications in Statistics: Simulation and Computation, 49(3):647–668, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2022:MLM

[ZC22] Qingle Zheng and Yuezhang Che. Multicategory large margin classification with unequal costs. Communications in Statistics: Simulation and Computation, 51(4):1801–1813, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:HTO

[ZD22] Haodong Zhang and Jieli Ding. Hypothesis testing in outcome-dependent sampling design under generalized linear models. Communications in Statistics: Simulation and Computation, 51(4):1721–1745, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:PIA

[ZFG22] Jun Zhang, Sanying Feng, and Yujie Gai. Partial index additive models with additive distortion measurement errors. Communications in Statistics: Simulation and Computation, 51(5):2191–2216, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:LRM

[ZG21] Jun Zhang and Yujie Gai. Linear regression models with general distortion measurement errors. Communications in Statistics: Simulation and Computation, 50(11):3383–3396, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:CCB

[ZG22a] Jun Zhang and Yujie Gai. Correlation coefficient-based measure for checking symmetry or asymmetry of a continuous variable with additive distortion. Communications in Statistics: Simulation and Computation, 51(5):2158–2190, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:NRM

[ZG22b] Jun Zhang and Yujie Gai. Nonlinear regression models with profile nonlinear least squares estimation. Communications in Statistics: Simulation and Computation, 51(5):2140–2157, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:EFM

[ZGC21] Caiya Zhang, Xiaolu Gu, and Yingyu Chen. Estimation for frailty measurement error Cox models based on profile likelihood and Bayes methods. Communications in Statistics: Simulation and Computation, 50(4):1025–1038, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Zhang:2021:DMH

[ZGL21] Jun Zhang, Yujie Gai, and Bingqing Lin. Detection of marginal heteroscedasticity for partial linear single-index models. Communications in Statistics: Simulation and Computation, 50(3):724–743, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:TSA

[ZGL22] Jun Zhang, Yujie Gai, and Feng Li. Testing symmetry for additive distortion measurement errors data. Communications in Statistics: Simulation and Computation, 51(3):1046–1065, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:VIV

[Zha22] Jiamin Zhang. Variational inference for varying-coefficient model. Communications in Statistics: Simulation and Computation, 51(2):670–685, 2022. CODEN CSSCDB. ISSN 0361-

Zheng:2021:IRL

[Zhe21] Songfeng Zheng. Iteratively reweighted least square for asymmetric L_{2}-loss support vector regression. Communications in Statistics: Simulation and Computation, 50(7):2151–2167, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:FPM

[ZHW21] Tao Zhang, Haifeng Huo, and Yanling Wan. Functional polynomial multiple-index model. Communications in Statistics: Simulation and Computation, 50(12):3958–3971, 2021. CO-DEN CSSCDB. ISSN 0361-0918.

Zeybek:2020:EGN

[ZK20a] Melis Zeybek and Onur Köksoy. The effects of gamma noise on quality improvement. Communications in Statistics: Simulation and Computation, 49(7):1783–1797, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zmyslony:2020:SHT

[ZK20b] Roman Zmyślony and Arkadiusz Kozioł. Simultaneous hypotheses testing in models with blocked compound symmetric covariance structure. Communications in Statistics: Simulation and Computation, 49(3):817–823, 2020. CODEN CSS-CDB. ISSN 0361-0918.

Zhu:2020:NMR

[ZL20] Lixia Zhu and Kung-Jong Lui. Notes on misspecifying the random effects distribution regarding analysis under the AB/ BA crossover trial in dichotomous data — a Monte Carlo evaluation. Communications in Statistics: Simulation and Computation, 49(2):419–435, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:AMS

[ZL22] Yuyang Zhang and Bo Lu. Accounting for matching structure in post-matching analysis of observational studies. Communications in Statistics: Simulation and Computation, 51(6): 3081–3099, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2020:BOA

[ZLSC20] Yicheng Zhou, Zhenzhou Lu, Yan Shi, and Kai Cheng. Bayesian optimum accelerated life test plans based on quantile regression. Communications in Statistics: Simulation and Computation, 49(9):2402–2418, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2022:NCP

[ZNKN22] Kaimeng Zhang, Chi Tim Ng, Yong Man Kwon, and Myung Hwan Na. New concepts of principal component analysis based on maximum separation of clusters. Communications in Statistics: Simulation and Computation, 51(5):2429– 2439, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:EED

[ZNLW20] Jun Zhang, Cuizhen Niu, Tao Lu, and Zhenghong Wei. Estimation of the error distribution function for partial linear single-index models. Communications in Statistics: Simulation and Computation, 49(1):29–44, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:SSB

[ZPW21] Xuemao Zhang, Sudhir Paul, and You-Gan Wang. Small sample bias correction or bias reduction? Communications in Statistics: Simulation and Computation, 50(4):1165–1177, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:SMI

[ZT20] Shuxia Zhang and Boping Tian. Semiparametric method for identifying multiple change-points in financial market. Communications in Statistics: Simulation and Computation, 49 (6):1429–1444, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:SRP

[ZWY20] Jie Zhang, Dehui Wang, and Kai Yang. A study of RCINAR(1) process with generalized negative binomial marginals. Communications in Statistics: Simulation and Computation, 49(6):1487–1510, 2020. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2021:NDC

[ZWY21] Weijia Zhang, Xikui Wang, and Po Yang. A new design of the continual reassessment method. Communications in Statistics: Simulation and Computation, 50(7):2015–2024, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:EPL

[ZXF22] Yang Zhao, Liugen Xue, and Sanying Feng. Estimation for a partially linear single-index varying-coefficient model. Communications in Statistics: Simulation and Computation, 51 (4):1685–1703, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zeng:2021:RBM

[ZZ21] Guoping Zeng and Emily Zeng. On the relationship between multicollinearity and separation in logistic regression. Communications in Statistics: Simulation and Computation, 50 (7):1989–1997, 2021. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2022:BBS

[ZZ22] Jian-Qiang Zhao and Yan-Yong Zhao. Bootstrap bandwidth selection in time-varying coefficient models with jumps. Communications in Statistics: Simulation and Computation, 51 (6):3124–3137, 2022. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2020:RBK

[ZZFG20] Jun Zhang, Jing Zhang, Zhenghui Feng, and Xu Guo. Residuals based Kolmogorov–Smirnov and Cramér–von Mises tests for varying coefficient models. Communications in Statistics: Simulation and Computation, 49(7):1798–1814, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Zhao:2020:NOE

[ZZWH20] Peixin Zhao, Xiaoshuang Zhou, Xiuli Wang, and Xingshou Huang. A new orthogonality empirical likelihood for varying coefficient partially linear instrumental variable models with longitudinal data. Communications in Statistics: Simulation and Computation, 49(12):3328–3344, 2020. CODEN CSSCDB. ISSN 0361-0918.